### v3

# Chevalley formulae for the motivic Chern classes of Schubert cells and for the stable envelopes

Abstract.
We prove a Chevalley formula to multiply the motivic Chern classes of Schubert cells in a generalized flag manifold $G/P$ by the class of any line bundle $\mathcal{L}_{\lambda}$. Our formula is given in terms of the $\lambda$-chains of Lenart and Postnikov. Its proof relies on a change of basis formula in the affine Hecke algebra due to Ram, and on the Hecke algebra action on torus-equivariant K-theory of the complete flag manifold $G/B$ via left Demazure–Lusztig operators. We revisit some wall-crossing formulae for the stable envelopes in $T^{*}(G/B)$. We use our Chevalley formula, and the equivalence between motivic Chern classes of Schubert cells and K-theoretic stable envelopes in $T^{*}(G/B)$, to give formulae for the change of polarization, and for the change of slope for stable envelopes. We prove several additional applications, including Serre, star, and Dynkin, dualities of the Chevalley coefficients, new formulae for the Whittaker functions, and for the Hall–Littlewood polynomials. We also discuss positivity properties of Chevalley coefficients, and properties of the coefficients arising from multiplication by minuscule weights.
Key words and phrases:
motivic Chern classes of Schubert cells, stable envelopes, $\lambda$-chains, Hecke algebra, Demazure–Lusztig operators
2020 Mathematics Subject Classification:
Primary 14M15, 14C17; Secondary 14N15, 17B10, 33D80.
L. C. Mihalcea was supported in part by the NSF grant DMS-2152294 and a Simons Collaboration Grant. H. Naruse was supported in part by JSPS KAKENHI Grant Number 16H03921.
1. Introduction
Let $G$ be a complex, semisimple, Lie group and $T\subset B\subset P\subset G$ be a parabolic subgroup containing a Borel subgroup and the (standard) maximal torus. Let $W$ be the Weyl group determined by $(G,T)$. In the study of cohomology and K-theory rings of (generalized) flag manifolds $G/P$, the Chevalley formula expresses the multiplication of a Schubert class by the class of line bundle, or, equivalently, a Schubert divisor in $G/P$. If one works equivariantly, this formula determines completely the multiplication in the equivariant K ring. In this paper we prove a Chevalley formula for the coefficients $C_{u,\lambda}^{w}(y)\in K_{T}(pt)[y]$ arising in the multiplication
(1) $MC_{y}(X(w)^{\circ})\cdot\mathcal{L}_{\lambda}=\sum C_{u,\lambda}^{w}(y)MC_{y}(X(u)^{\circ})$
in the equivariant K-theory ring $K_{T}(G/B)[y]$; see (27) and Theorem 5.2 below. Here $MC_{y}(X(w)^{\circ})\in K_{T}(G/B)[y]$ is the motivic Chern class of a Schubert cell $X(w)^{\circ}\subset G/B$, and $\mathcal{L}_{\lambda}=G\times^{B}{\mathbb{C}}_{\lambda}$ is the line bundle on $G/B$ associated to the one dimensional $B$-module of weight $\lambda$.
The motivic Chern classes $MC_{y}(X(w)^{\circ})\in K_{T}(G/B)[y]$ have been defined by Brasselet, Schürmann, and Yokura [BSY10] more generally for elements $[Y\to X]$ in the Grothendieck group $G_{0}(var/X)$ of varieties over $X$. They are the unique classes which are functorial with respect to proper morphisms $f:X_{1}\to X_{2}$, and which satisfy the normalization condition $MC[id_{X}:X\to X]=\lambda_{y}(T^{*}_{X})$ for $X$ smooth, where $\lambda_{y}(T^{*}_{X})=\sum y^{i}[\wedge^{i}T^{*}_{X}]$ is the Hirzebruch $\lambda_{y}$ class; see §4 below. They may be thought of as the K-theoretic generalizations of Chern–Schwartz–MacPherson classes defined by MacPherson [Mac74].
The motivic Chern classes of Schubert cells generalize well studied classes from Schubert calculus. If $y=0$, the motivic class $MC_{y}(X(w)^{\circ})$ is equal to the class of the ideal sheaf $[{\mathcal{O}}_{X(w)}(-\partial X(w))]$ of the boundary $\partial X(w)=X(w)\setminus X(w)^{\circ}$, where $X(w)=\overline{X(w)^{\circ}}$ is the Schubert variety. If $y=-1$, then $MC_{y}(X(w)^{\circ})$ is equal to the class of the unique $T$-fixed point in $X(w)^{\circ}$; see [AMSS22]. The Poincaré duals of the classes $MC_{y}(X(w)^{\circ})$, the Segre motivic classes [AMSS19, MSA22], specialize when $y=0$ to the Grothendieck classes of the structure sheaves of the opposite Schubert varieties. Our Chevalley formula (1), and its analogous formula for the Segre motivic classes, specialize to known Chevalley formulae for K-theoretic Schubert classes and ideal sheaves from [GR04, LP07].
The formula for the coefficients $C_{u,\lambda}^{w}(y)\in K_{T}(pt)[y]$ from (1) follows from a formula of Ram [Ram06] in the affine Hecke algebra $\mathbb{H}$, calculating transition coefficients between two bases $\{T_{w}X^{\lambda}\}$ and $\{X^{\lambda}T_{w}\}$ of the affine Hecke algebra:
(2) $T_{w}X^{-\lambda}=\sum_{\mu\in X^{*}(T),u\in W}(-q)^{\ell(w)-\ell(u)}c_{u,\mu}^{w,\lambda}X^{-\mu}T_{u}\/.$
Here $T_{w}$ is an element in the standard basis of $\mathbb{H}$, $X^{-\lambda}$ is an affine element in $\mathbb{H}$, and $X^{*}(T)$ denotes the weight lattice of $T$. Ram’s formula is stated in terms of a combinatorial model utilizing alcove walks, and it is convenient for our purposes to rewrite it utilizing in terms of $\lambda$-chains, a model introduced and studied by Lenart and Postnikov [LP07, LP08] in relation to equivariant K theory of flag manifolds. We refer to Theorem 3.9 and Theorem 3.10 for the precise statements in the Hecke algebra in terms of $\lambda$-chains, and to §5.1 for the formulae involving motivic Chern classes. We also note that (affine) Hecke algebras have long been used to obtain Chevalley formulas in various contexts, for example in [PR99, LP07]. We state next our main result.
Assume $\lambda$ is an integral weight and fix a reduced $\lambda$-chain $(\beta_{1},\beta_{2},\ldots,\beta_{l})$. The chain corresponds to an alcove walk from the fundamental alcove $\operatorname{A_{\circ}}$ to $\operatorname{A_{\circ}}-\lambda$, with separating hyperplanes $H_{-\beta_{j},d_{j}}$. Denote by $s_{\beta}$ the reflection determined by the root $\beta$. We refer the reader to §3 below for full definitions. The following is our main result, cf. Theorem 5.5.
Theorem 1.1.
The following Chevalley formula holds in $K_{T}(G/B)[y]$:
${\mathcal{L}}_{-\lambda}\otimes MC_{y}(X(w)^{\circ})=\sum_{\mu\in X^{*}(T),u\in W}C_{u,-\lambda}^{w}MC_{y}(X(u)^{\circ}),$
where the Chevalley coefficients are given by
(3) $C_{u,-\lambda}^{w}=\sum_{J\subset\{1,2,\ldots,l\}}(-1)^{n(J)}(1+y)^{|J|}(-y)^{\frac{\ell(w)-\ell(u)-|J|}{2}}e^{-w\tilde{r}_{J_{>}}(\lambda)},$
and the sum is over subsets $J=\{j_{1}<\ldots<j_{t}\}\subset\{1,2,\ldots,l\}$ such that $u<us_{\beta_{j_{1}}}<us_{\beta_{j_{1}}}s_{\beta_{j_{2}}}<\ldots<us_{\beta_{j_{1}}}s_{\beta_{j_{2}}}\cdot\ldots\cdot s_{\beta_{j_{t}}}=w$; the Weyl group element $\tilde{r}_{J_{>}}$ is defined in (19). For the multiplication ${\mathcal{L}}_{\lambda}\otimes MC_{y}(X(w)^{\circ})$, the Chevalley coefficients are given by
(4) $C_{u,\lambda}^{w}=\sum_{J\subset\{1,2,\ldots,l\}}(-1)^{n(J)}(-1-y)^{|J|}(-y)^{\frac{\ell(w)-\ell(u)-|J|}{2}}e^{-w\hat{r}_{J_{<}}(-\lambda)},$
where the sum is over subsets $J=\{j_{1}<\ldots<j_{t}\}\subset\{1,2,\ldots,l\}$ such that $u<us_{\beta_{j_{t}}}<us_{\beta_{j_{t}}}s_{\beta_{j_{t-1}}}<\ldots<us_{\beta_{j_{t}}}\cdot\ldots\cdot s_{\beta_{j_{1}}}=w$, and with $\hat{r}_{J_{<}}$ defined in (16).
The connection between the Hecke algebra coefficients from (2) and the Chevalley coefficients above is given by
(5) $C_{u,-\lambda}^{w}=\sum_{\mu\in X^{*}(T)}y^{\ell(w)-\ell(u)}e^{-\mu}c_{u,\mu}^{w,\lambda}|_{q=-y}.$
The coefficients $c_{u,\mu}^{w,\lambda}$ are in general Laurent polynomials in $y$, while $C_{u,-\lambda}^{w}$ are polynomials in $K_{T}(pt)[y]$. In fact, the power $y^{\ell(w)-\ell(u)}$ from the formula (5) is absorbed into $c_{u,\mu}^{w,\lambda}$ so it becomes polynomial in $y$.
[…]
We now give a rough idea on the proof of Theorem 1.1. The key connection between the Chevalley formula in the Hecke algebra to motivic Chern classes, proved in [MNS22b], and ultimately based on results from [AMSS19], is that the motivic Chern classes are recursively obtained by certain left Demazure–Lusztig operators $\mathcal{T}_{w}^{L}$ acting on $K_{T}(G/B)[y]$:
$MC_{y}(X(w)^{\circ})=\mathcal{T}_{w}^{L}[\mathcal{O}_{1.B}]\/.$
These operators commute with elements in $K_{G}(G/B)[y]$ (i.e., the Weyl-group invariants of $K_{T}(G/B)$), and an argument based on equivariant localization shows that
$\begin{split}MC_{y}(X(w)^{\circ})\cdot\mathcal{L}_{\lambda}=\mathcal{T}_{w}^{L}[\mathcal{O}_{1.B}]\cdot\mathcal{L}_{\lambda}=\mathcal{T}_{w}^{L}(\mathcal{L}_{\lambda}\cdot[\mathcal{O}_{1.B}])=\mathcal{T}_{w}^{L}(e^{\lambda}\cdot[\mathcal{O}_{1.B}])\/.\end{split}$
Therefore, the knowledge of the expansion from (2) implies the Chevalley formula in the geometric case. This argument may be generalized to any homogeneous bundle, see Remark 5.3 below. In cohomology (i.e., for the Chern–Schwartz–MacPherson classes), and for $G=\mathrm{SL}_{n}$, this argument is implicitly utilized in the paper [FGX22] to obtain a Murnaghan–Nakayama formula.
We briefly survey next other results from this note. Having established a formula to calculate the Chevalley coefficients, in §6 we utilize several dualities with geometric origin (the Serre duality, the star duality, and the Dynkin automorphism duality) to obtain several symmetries of the coefficients $C_{v,\lambda}^{w}(y)$. See e.g. 6.5. Combining these dualities shows that the polynomials $C_{v,\lambda}^{w}(y)$ are palindromic.
A remarkable property of the motivic Chern classes of Schubert cells, proved in [AMSS19, FRW21], is that they are equivalent to the K-theoretic version of Maulik and Okounkov’s stable envelopes, see [MO19, AO21]. The stable envelopes are elements in the $T\times{\mathbb{C}}^{*}$-equivariant K-theory of the cotangent bundle, $K_{T\times{\mathbb{C}}^{*}}(T^{*}(G/B))$. In this context, the formal variable $y$ may be identified to the (inverse) of the character given by the ${\mathbb{C}}^{*}$ fibre dilation on the cotangent bundle. If $\iota:G/B\hookrightarrow T^{*}(G/B)$ is the inclusion of the zero section, then $\iota^{*}(\operatorname{stab}(w))$ is a multiple of the motivic Chern class of the (opposite) Schubert cell for $w$, where $\operatorname{stab}$ is a stable envelope, appropriately normalized.
The stable envelopes depend on three parameters: a chamber, a polarization, and a slope, and the precise normalizations are essential for this paper. A variation in the chamber results in conjugating by the Borel subgroup [AMSS19], and it is encoded in the left Weyl group action [MNS22b] and certain $R$-matrix operators [RTV15, RTV19]. Varying the polarization, or the slope, results in the multiplication of $\operatorname{stab}(w)$ by a line bundle $\mathcal{L}_{\lambda}$ pulled back from $G/B$; cf. [AMSS19, Oko17], see also §7 below. In particular, the coefficients $C_{v,\lambda}^{w}(y)$ from (1) give ‘wall-crossing’ formulae, recording the change of stable envelopes when its defining parameters are varied. While these wall crossing formulae have been worked out in [Oko17, SZZ20, SZZ21] (see also [KW23]), in §7 we revisit some of these from the point of view of Theorem 1.1. In particular, we utilize the Chevalley formula to give an explicit combinatorial rule relating the stable envelope for the fundamental alcove $\operatorname{A_{\circ}}$ to the one corresponding to any translation $\operatorname{A_{\circ}}+\lambda$; see Theorem 7.8.
In addition to our application mentioned above to wall crossing formulae for stable envelopes, in §8 we utilize known relations between motivic Chern classes of Schubert cells, Whittaker functions, and Hall–Littlewood polynomials, to obtain new formulae for the latter.
Finally, in an Appendix we obtain an analogue of the Chevalley formula (1) for the homological analogue of the motivic Chern classes, the Chern–Schwartz–MacPherson classes. While this formula may be obtained by a specialization argument as in [AMSS22], the degenerate affine Hecke algebra is much simpler in this case, and a direct proof of the Chevalley formula may be obtained rather quickly.
Acknowledgments. HN and CS thank Arun Ram and Andrei Okounkov for answering their questions and providing references; LM and CS thank Bogdan Ion for a related collaboration on Hall–Littlewood polynomials, and Paolo Aluffi and Jörg Schürmann for previous collaborations on motivic Chern classes and CSM classes; LM thanks Andrzej Weber for stimulating discussions. This project started in Spring 2021 while LM was in residence at the Institute for Computational and Experimental Research in Mathematics in Providence, RI, during the Combinatorial Algebraic Geometry program; LM wishes to thank the institute for excellent working conditions.
Notation
We fix the notation utilized throughout the paper. Let $G$ be a simply connected complex Lie group with Borel subgroup $B$ and maximal torus $T\subset B$. Denote by $\mathfrak{g}=\mathrm{Lie}(G)$ and by $\mathfrak{h}=\mathrm{Lie}(T)$ be corresponding Lie algebras. Let $R^{+}\subset\mathfrak{h}^{*}:=\mathfrak{h}^{*}_{\mathbb{Q}}$ denote the positive roots, which by convention are the roots in $B$, and by $\Sigma=\{\alpha_{i}:i\in I\}$ the set of simple roots. The set of all roots is $R:=R^{+}\sqcup-R^{+}$. We use $\alpha>0$ (resp. $\alpha<0$) to denote $\alpha\in R^{+}$ (resp. $\alpha\in-R^{+}$). For any root $\alpha\in R$, let $\alpha^{\vee}\subset\mathfrak{h}$ denote the corresponding coroot. Denote by $\langle\cdot,\cdot\rangle:\mathfrak{h}^{*}\times\mathfrak{h}\to{\mathbb{Q}}$ the evaluation pairing, and let $X^{*}(T)\subset\mathfrak{h}^{*}$ be the weight lattice. For any weight $\lambda\in X^{*}(T)$, let ${\mathcal{L}}_{\lambda}:=G\times^{B}{\mathbb{C}}_{\lambda}$ be the line bundle on $G/B$ associated to $\lambda$. The Weyl group is $W=N_{G}(T)/T$ and it is generated by simple reflections $s_{i}=s_{\alpha_{i}}$ ($i\in I$). It is equipped with a length function $\ell:W\to{\mathbb{Z}}_{\geq 0}$ defined as the length of a minimal expression of $w$ in terms of the simple reflections; we denote by $w_{0}$ the longest element. The Bruhat order on $W$ is a partial order determined by the covering relations $u\leq us_{\alpha}$ where $\alpha\in R$ and $\ell(us_{\alpha})=\ell(u)+1$.
For any $w\in W$, let $X(w)^{\circ}:=BwB/B\subset G/B$ and $Y(w)^{\circ}:=B^{-}wB/B\subset G/B$ be Schubert cells, where $B^{-}$ is the opposite Borel subgroup. Let $X(w):=\overline{X(w)^{\circ}}$ and $Y(w):=\overline{Y(w)^{\circ}}$ be the Schubert varieties, respectively. Let $P(\supseteq B)$ be a parabolic subgroup with simple roots $\Sigma_{P}\subset\Sigma$. Let $R_{P}^{+}$ denote the positive roots spanned by $\Sigma_{P}$. Let $W_{P}$ be the Weyl group generated by the simple reflections $s_{\alpha}$, $\alpha\in\Sigma_{P}$. Let $W^{P}\simeq W/W_{P}$ denote the set of minimal length representatives. For any $w\in W^{P}$, let $X(wW_{P})^{\circ}:=BwP/P\subset G/P$ (resp. $Y(wW_{P})^{\circ}:=B^{-}wP/P\subset G/P$) denote the Schubert cell with closure $X(wW_{P})$ (resp. $Y(wW_{P})$). Let $X^{*}(T)_{P}:=\{\lambda\in X^{*}(T)\mid\langle\lambda,\gamma^{\vee}\rangle=0\text{ for all }\gamma\in R^{+}_{P}\}$ be the set of integral weights which vanish on $(R^{+}_{P})^{\vee}$. For any $\lambda\in X^{*}(T)_{P}$, we still use ${\mathcal{L}}_{\lambda}$ to denote the line bundle $G\times^{P}{\mathbb{C}}_{\lambda}\in\operatorname{Pic}(G/P)$, which has fiber over $1.P$ the $T$-module of weight $\lambda$.
2. Affine Hecke algebra via alcove walk algebra
In this section, we introduce the alcove walk algebra, and a formula of Ram [Ram06] describing a change of bases matrix for the affine Hecke algebra.
2.1. Affine Hecke algebra
The affine Hecke algebra ${\mathbb{H}}$ is a free ${\mathbb{Z}}[q,q^{-1}]$ module with basis $\{T_{w}X^{\lambda}|w\in W,\lambda\in X^{*}(T)\}$, such that
  • For any $\lambda,\mu\in X^{*}(T)$, $X^{\lambda}X^{\mu}=X^{\lambda+\mu}$.
  • For any simple root $\alpha$, $(T_{s_{\alpha}}+1)(T_{s_{\alpha}}-q)=0$.
  • For any $w,y\in W$, such that $\ell(wy)=\ell(w)+\ell(y)$, $T_{w}T_{y}=T_{wy}$
  • For any simple root $\alpha$ and $\lambda\in X^{*}(T)$,
    $T_{s_{\alpha}}X^{\lambda}-X^{s_{\alpha}\lambda}T_{s_{\alpha}}=(1-q)\frac{X^{s_{\alpha}\lambda}-X^{\lambda}}{1-X^{-\alpha}}.$
For our geometric application we will need two other bases of the affine Hecke algebra ${\mathbb{H}}$: $\{T_{w^{-1}}^{-1}X^{\lambda}\mid w\in W,\lambda\in X^{*}(T)\}$ and $\{X^{\lambda}T_{w^{-1}}^{-1}\mid w\in W,\lambda\in X^{*}(T)\}$. Define the transition matrix coefficients $c_{u,\mu}^{w,\lambda}\in{\mathbb{Z}}[q,q^{-1}]$ by
(6) $T_{w^{-1}}^{-1}X^{\lambda}=\sum_{\mu\in X^{*}(T),u\in W}c_{u,\mu}^{w,\lambda}X^{\mu}T_{u^{-1}}^{-1}.$
The main result of this section is a formula for $c_{u,\mu}^{w,\lambda}$ obtained by Ram [Ram06], see Theorem 2.4 below.
For the later application to the motivic Chern classes, we also introduce the Iwahori–Matsumoto ${\mathbb{Z}}[q,q^{-1}]$-algebra involution $\Theta$ on ${\mathbb{H}}$ defined by
$\Theta(T_{s_{\alpha}})=-qT_{s_{\alpha}}^{-1},\textit{\quad and \quad}\Theta(X^{\lambda})=X^{-\lambda},$
where $s_{\alpha}$ is a simple reflection; see [EM97, Section 5.1]. Hence, $\Theta(T_{w^{-1}}^{-1})=(-q)^{-\ell(w)}T_{w}$. Applying $\Theta$ to Equation (6), we obtain:
(7) $T_{w}X^{-\lambda}=\sum_{\mu\in X^{*}(T),u\in W}(-q)^{\ell(w)-\ell(u)}c_{u,\mu}^{w,\lambda}X^{-\mu}T_{u}$
2.2. Alcove walk algebra
In this section, we review Ram’s definition of the alcove walk algebra, and state his formula for the matrix coefficients $c_{u,\mu}^{w,\lambda}$. We refer the reader to [Ram06] for a more detailed account of the alcove walk algebras.
2.2.1. Alcoves
Let $\mathfrak{t}^{*}_{\mathbb{R}}$ be the dual of the Lie algebra of the maximal torus $T$. For any root $\alpha$ and $j\in{\mathbb{Z}}$, define
$H_{\alpha,j}:=\{\lambda\in\mathfrak{t}^{*}_{\mathbb{R}}\mid\langle\lambda,\alpha^{\vee}\rangle=j\}.$
Notice that $H_{\alpha,j}=H_{-\alpha,-j}$. The connected components of $\mathfrak{t}^{*}_{\mathbb{R}}\setminus\cup_{\alpha>0,j\in{\mathbb{Z}}}H_{\alpha,j}$ are called alcoves. The codimension 1 faces of any alcove are called the walls of that alcove. The fundamental alcove $\operatorname{A_{\circ}}$ is defined by:
$\operatorname{A_{\circ}}=\{\lambda\in\mathfrak{t}^{*}_{\mathbb{R}}\mid 0<\langle\lambda,\alpha^{\vee}\rangle<1,\textit{ for any positive root }\alpha\}.$
If $\alpha_{1},\alpha_{2},\ldots,\alpha_{r}$ denote the simple roots, and $\theta^{\vee}$ denotes the highest coroot, then the walls of the fundamental alcove $\operatorname{A_{\circ}}$ are $H_{\theta,1}$ and $H_{\alpha_{i},0}$ ($1\leq i\leq r$). We label these walls of $\operatorname{A_{\circ}}$ by $0,1,\ldots,r$ respectively.
The affine Weyl group for the dual root system is defined by $W_{\operatorname{aff}}:=Q\rtimes W$, where $Q$ is the root lattice. Then $W_{\operatorname{aff}}$ acts simply transitively on the set of alcoves, and this action is determined by the reflections across the hyperplanes $h=H_{\alpha,j}$, given by
(8) $s_{\alpha,j}(\mu)=\hat{r}_{h}(\mu):=s_{\alpha}(\mu)+j\alpha\text{ for }\mu\in X^{*}(T),$
The affine Weyl group is a Coxeter group generated by the reflections $s_{0}:=s_{\theta,1}$ and $s_{i}$ ($1\leq i\leq r)$ along the walls of $\operatorname{A_{\circ}}$. In fact, $\operatorname{A_{\circ}}$ is a fundamental domain for the action of $W_{\operatorname{aff}}$ on the set of alcoves, in the sense that any element in $\mathfrak{t}^{*}_{\mathbb{R}}\setminus\cup_{\alpha>0,j\in{\mathbb{Z}}}H_{\alpha,j}$ is sent to exactly one element in $\operatorname{A_{\circ}}$. See [Ram06, Hum90] for more details.
The extended affine Weyl group for the dual root system is $W^{\operatorname{ext}}_{\operatorname{aff}}=X^{*}(T)\rtimes W$, where $X^{*}(T)$ is the weight lattice. For any $\lambda\in X^{*}(T)$, let $t_{\lambda}$ denote the corresponding element in $W^{\operatorname{ext}}_{\operatorname{aff}}$. There is a length function $\ell$ on $W^{\operatorname{ext}}_{\operatorname{aff}}$ defined by the following formula (see [Mac96, Equation (2.8)]):
$\ell(t_{\mu}w)=\sum_{\alpha\in R^{+}}|\langle\mu,w(\alpha^{\vee})\rangle+\chi(w(\alpha))|\quad\textrm{ where }\quad\chi(\alpha)=\begin{cases}0&\alpha\in R^{+}\\ 1&\alpha\in R^{-}\/.\end{cases}$
Let $\Omega\subset W^{\operatorname{ext}}_{\operatorname{aff}}$ be the subgroup of length zero elements in $W^{\operatorname{ext}}_{\operatorname{aff}}$. Then $W_{\operatorname{aff}}^{\operatorname{ext}}\simeq W_{\operatorname{aff}}\rtimes\Omega$, see [Mac96, Equation (2.10)], and $\ell(wg)=\ell(w)$ for any $w\in W_{\operatorname{aff}}$ and $g\in\Omega$. The elements in $\Omega$ preserve the fundamental alcove $\operatorname{A_{\circ}}$ and act as automorphisms.
Utilizing a bijection between $W_{\operatorname{aff}}$ and the alcoves in $\mathfrak{t}^{*}_{\mathbb{R}}$, one can define a bijection between $W_{\operatorname{aff}}^{\operatorname{ext}}\simeq W_{\operatorname{aff}}\rtimes\Omega$ and the alcoves in $\Omega\times\mathfrak{t}^{*}_{\mathbb{R}}$ ($|\Omega|$ copies of $\mathfrak{t}^{*}_{\mathbb{R}}$, each tiled by alcoves). We label the walls of every alcove in $\Omega\times\mathfrak{t}^{*}_{\mathbb{R}}$ in an $W^{\operatorname{ext}}_{\operatorname{aff}}$-equivariant way. This means that for each $w\in W^{\operatorname{ext}}_{\operatorname{aff}}$ the walls of $w\operatorname{A_{\circ}}$ are $wH_{\alpha_{i},0}$ ($1\leq i\leq n$) and $wH_{\theta,1}$, and they are labeled by $i$ and 00, respectively. In particular, if two adjacent alcoves $A_{1}$ and $A_{2}$ are separated by a wall labeled by $i$ (in both $A_{1}$ and $A_{2}$), and $A_{1}=w\operatorname{A_{\circ}}$ for some $w\in W^{\operatorname{ext}}_{\operatorname{aff}}$, then $A_{2}=ws_{i}\operatorname{A_{\circ}}$. Equivalently, in terms of the wall crossings, if $w=gs_{i_{1}}s_{i_{2}}\cdot\ldots\cdot s_{i_{\ell}}\in W^{\operatorname{ext}}_{\operatorname{aff}}$, with $g\in\Omega$ and $0\leq i_{j}\leq r$, then the alcove $w\operatorname{A_{\circ}}$ in $\Omega\times\mathfrak{t}^{*}_{\mathbb{R}}$ is the alcove obtained from rotating the fundamental alcove $\operatorname{A_{\circ}}$ according to the automorphism $g$, then reflect along the walls labelled (in order) by $i_{1},\ldots,i_{\ell}$; see also Lemma 3.3 below.
Example 2.1.
[…]
2.2.2. Alcove walk algebra
In this section we recall a realization of the Hecke algebra in terms of alcove walks; see [Ram06]. For each positive root $\alpha$ and hyperplane $H_{\alpha,j}$, set the positive side of it to be $\{\lambda\in\mathfrak{t}^{*}_{\mathbb{R}}\mid\langle\lambda,\alpha^{\vee}\rangle>j\}$.
Definition 2.2.
The alcove walk algebra is generated over ${\mathbb{Z}}[q,q^{-1}]$ by elements $g\in\Omega$, and for $0\leq i\leq r$, the elements $c_{i}^{+}$ (positive $i$-crossing), $c_{i}^{-}$ (negative $i$-crossing), $f_{i}^{+}$ (positive $i$-fold) and $f_{i}^{-}$ (negative $i$-fold), subject to the following relations, sometimes called straightening laws:
$c_{i}^{+}=c_{i}^{-}+f_{i}^{+},\quad c_{i}^{-}=c_{i}^{+}+f_{i}^{-}\/,\quad\textrm{ and }\quad gc_{i}^{\pm}=c_{g(i)}^{\pm}g\/,\quad gf_{i}^{\pm}=f_{g(i)}^{\pm}g\/.$
In terms of pictures, these generators can be drawn as follows:
--$+$$i$$c_{i}^{+}$-$+$$i$$c_{i}^{-}$-$+$$i$$f_{i}^{+}$-$+$$i$$f_{i}^{-}$
$(0\leq i\leq r)$
Here, $c_{i}^{+}$ represents a crossing of a wall labeled by $i$ from its negative to its positive side, and similarly for the other generators. The product is given as concatenation. An alcove walk is a word in the generators such that,
  • the tail of the first step is in the fundamental alcove $\operatorname{A_{\circ}}$;
  • at every step, either we change the sheet according to an element in $\Omega$ (thus rotating the alcove according to this elements), or the head of each arrow is in the same alcove as the tail of the next arrow.
An alcove walk $p$ is called nonfolded if there is no $f_{i}^{\pm}$ in its word. The length of an alcove walk is the number of letters $c_{i}^{\pm},f_{i}^{\pm}$ in an alcove walk. (In particular, rotation with respect to an element of $\Omega$ does not contribute to the length.) For a minimal alcove walk between two alcoves, one can show that the walk is non-folded, thus its length is the number of $c_{i}^{\pm}$ in the walk [Ram06]. From this it follows that if $w\in W^{\operatorname{ext}}_{\operatorname{aff}}$, then $\ell(w)$=length of a minimal length walk from $\operatorname{A_{\circ}}$ to $w\operatorname{A_{\circ}}$.
Pick a square root $q^{\frac{1}{2}}$ of $q$. The following is proved by A. Ram.
Proposition 2.3.
[Ram06, §3.2]
  1. (a)
    The affine Hecke algebra ${\mathbb{H}}$ is isomorphic as a ${\mathbb{Z}}[q^{\frac{1}{2}},q^{-\frac{1}{2}}]$-algebra to the quotient of the alcove walk algebra by the relations
    (9) $c_{i}^{+}=(c_{i}^{-})^{-1},\quad f_{i}^{+}=(q^{\frac{1}{2}}-q^{-\frac{1}{2}}),\quad f_{i}^{-}=-(q^{\frac{1}{2}}-q^{-\frac{1}{2}})$
    and
    (10) $p=p^{\prime}\text{ if }p\text{ and }p^{\prime}\text{ are nonfolded alcove walks with }end(p)=end(p^{\prime}),$
    where $end(p)$ means the final alcove of $p$.
  2. (b)
    Under the previous isomorphism, for any $w\in W$ and $\lambda\in X^{*}(T)$  [^1] The $q^{2}$ and $T_{s_{i}}$ in [Ram06, Proposition 3.2(b)] are our $q$ and $q^{-\frac{1}{2}}T_{i}$, respectively.:
    • a minimal length alcove walk from $\operatorname{A_{\circ}}$ to $w\operatorname{A_{\circ}}$ is sent to $q^{\frac{\ell(w)}{2}}T^{-1}_{w^{-1}}$, and,
    • a minimal length alcove walk from $\operatorname{A_{\circ}}$ to $t_{\lambda}\operatorname{A_{\circ}}$ is sent to $X^{\lambda}$.
For an alcove walk $p$, define the functions weight $wt(p)\in X^{*}(T)$, and final direction $\varphi(p)\in W$ of $p$ by the condition that $p$ ends in the alcove $t_{wt(p)}\varphi(p)\operatorname{A_{\circ}}$. Let
$f^{-}(p)$ $=\textit{number of negative folds of }p,$
$f^{+}(p)$ $=\textit{number of positive folds of }p,\textit{ and}$
$f(p)$ $=f^{+}(p)+f^{-}(p).$
Now we can state the formula for the matrix coefficients $c_{u,\mu}^{w,\lambda}$ defined in Equation (6), see also Equation (7).
Theorem 2.4.
[Ram06, Theorem 3.3] Let $\lambda\in X^{*}(T)$ and $w\in W$. Fix a minimal length walk $p_{w}=c_{i_{1}}^{-}c_{i_{2}}^{-}\cdots c_{i_{r}}^{-}$ from $\operatorname{A_{\circ}}$ to $w\operatorname{A_{\circ}}$ and a minimal length walk $p_{\lambda}=c_{j_{1}}^{\epsilon_{1}}c_{j_{2}}^{\epsilon_{2}}\cdots c_{j_{s}}^{\epsilon_{s}}g$ from $\operatorname{A_{\circ}}$ to $t_{\lambda}\operatorname{A_{\circ}}$, where $g\in W^{ext}_{\operatorname{aff}}$ is defined by the condition $gW_{\operatorname{aff}}=t_{\lambda}W_{\operatorname{aff}}$ 222 We need to add this extra $g\in\Omega$ since $t_{\lambda}\operatorname{A_{\circ}}$ may not on the same sheet as $\operatorname{A_{\circ}}$, see Example 2.5. The stated conditions determine $g$ uniquely., and $\epsilon_{i}=\pm$ for each $i$. Then
(11) $T^{-1}_{w^{-1}}X^{\lambda}=\sum_{p}(-1)^{f^{-}(p)}(q^{\frac{1}{2}}-q^{-\frac{1}{2}})^{f(p)}q^{\frac{\ell(\varphi(p))-\ell(w)}{2}}X^{wt(p)}T_{\varphi(p)^{-1}}^{-1},$
where the sum is over all alcove walks $p$ of the form
(12) $p=c_{i_{1}}^{-}c_{i_{2}}^{-}\cdots c_{i_{r}}^{-}p_{j_{1}}p_{j_{2}}\cdots p_{j_{s}}g\text{ such that }p_{j_{k}}\in\{c_{j_{k}}^{\pm},f_{j_{k}}^{\epsilon_{k}}\}\/.$
Therefore the matrix coefficients $c_{u,\mu}^{w,\lambda}$ in Equation 6 are given by:
[…]
Example 2.5.
Let $G=\operatorname{SL}(2,{\mathbb{C}})$. We use the same notation as in Example 2.1. We check the above theorem for $w=s_{1}$ and $\lambda=\omega$. From the alcove picture in Example 2.1, $T_{s_{1}}^{-1}$ is represented by the minimal length walk $q^{-1/2}c_{1}^{-1}$, while $X^{\omega}$ is represented by the walk $gc_{1}^{+}=c_{0}^{+}g$. Thus, the sum in the right hand side of the Theorem is over the alcove walks $c_{1}^{-}c_{0}^{-}g$ and $c_{1}^{-}f_{0}^{+}g$, which end at the alcoves $t_{-\omega}s_{1}\operatorname{A_{\circ}}$ and $t_{-\omega}\operatorname{A_{\circ}}$, respectively. (Note that $c_{1}^{-}c_{0}^{+}g$, which represents $T_{s_{1}}^{-1}X^{\omega}$, is not an alcove walk.) Therefore, the identity in the theorem is
$T_{s_{1}}^{-1}X^{\omega}=X^{-\omega}T_{s_{1}}^{-1}-q^{-1}(1-q)X^{-\omega}.$
On the other hand, it is easy to check the above equation using the definition of the affine Hecke algebra in Section 2.1.
Remark 2.6.
In Theorem 2.4, one may relax the hypotheses about the alcove walks $p_{w}$ and $p_{\lambda}$ to be of non-minimal length. This follows from analyzing the proof of Ram’s result in [Ram06]. We do not use this, but it is consistent with our use of non-reduced $\lambda$-chains in §3 below.
Consider an ordered collection of hyperplanes =$\mathcal{H}=\{H_{\beta_{1},k_{1}},\ldots,H_{\beta_{s},k_{s}}\}$ and set $h_{i}:=H_{\beta_{i},k_{i}}$. Associated to this sequence we define the elements
$\hat{r}_{\mathcal{H}}=s_{\beta_{1},k_{1}}\cdot\ldots\cdot s_{\beta_{s},k_{s}}\in W_{\operatorname{aff}}\/;\quad r_{\mathcal{H}}={s}_{\beta_{1}}\cdot\ldots\cdot s_{\beta_{s}}\in W\/.$
(Note that these depend on the order of $\mathcal{H}$.) We also define an $\mathcal{H}$-restricted version of the Bruhat order on $W$ by
(14) $w\stackrel{{\scriptstyle\mathcal{H}}}{{\Longrightarrow}}u\iff w>ws_{\beta_{1}}>ws_{\beta_{1}}s_{\beta_{2}}>\cdots>ws_{\beta_{1}}s_{\beta_{2}}\cdot\ldots\cdot s_{\beta_{s}}=u.$
The following is a mild generalization of [Len11, Proposition 2.5].
Lemma 2.7.
Let $w\in W$ and $\lambda\in X^{*}(T)$. Fix:
  • an alcove walk $p_{w}$ from $\operatorname{A_{\circ}}$ to $w\operatorname{A_{\circ}}$;
  • an alcove walk $p_{\lambda}=c_{j_{1}}^{\epsilon_{1}}c_{j_{2}}^{\epsilon_{2}}\cdots c_{j_{s}}^{\epsilon_{s}}g$ from $\operatorname{A_{\circ}}$ to $t_{\lambda}\operatorname{A_{\circ}}$.
Let $h_{i}=H_{\beta_{i},k_{i}}$, for $1\leq i\leq s$ be the sequence of hyperplanes defined by the walls of alcoves crossed by $p_{\lambda}$, with $\beta_{i}\in R^{\epsilon_{i}}$ and $k_{i}\in{\mathbb{Z}}$. Then there is a bijection between the following two sets:
  1. (1)
    The set of alcove walks of the form $\bar{p}=p_{w}p_{j_{1}}\ldots p_{j_{s}}$ such that $p_{j_{k}}\in\{c_{j_{k}}^{\pm},f_{j_{k}}^{\epsilon_{k}}\}$;
  2. (2)
    The set of subsets =$\mathcal{M}=\{h_{m_{1}},\ldots,h_{m_{t}}\}\subset\mathcal{H}=\{h_{1},\ldots,h_{s}\}$ with $m_{1}<m_{2}<\ldots<m_{t}$, s.t. $w\stackrel{{\scriptstyle\mathcal{M}}}{{\Longrightarrow}}wr_{\mathcal{M}}$.
Under this bijection, the indices $m_{i}$ correspond to the positions of foldings $f_{m_{i}}^{\epsilon_{m_{i}}}$. Furthermore,
$\varphi(\bar{p}g)=wr_{\mathcal{M}}\textrm{ and }\mathrm{wt}(\bar{p}g)=w\hat{r}_{\mathcal{M}}(\lambda)\/.$
Remark 2.8.
This statement is a slight generalization of Lenart’s result. We do not require $\lambda$ to be dominant, and therefore we need to allow $\beta_{i}\in R^{\epsilon_{i}}$ to be a negative root.
Proof.
The proof follows the same outline as that of [Len11, Proposition 2.5], so we will be brief. To start, consider the unique unfolded alcove walk $p_{0}=p_{w}p_{j_{1}}\ldots p_{j_{s}}$ such that $p_{j_{k}}\in\{c_{j_{k}}^{\pm}\}$. Any other alcove walk $\bar{p}$ as in the statement is of the form
$\bar{p}=\phi_{m_{t}}\cdots\phi_{m_{1}}(p_{0})\/,$
where $\phi_{m_{j}}$ is the folding operation at position $m_{j}$ (cf.  [Len11]) and the $m_{i}$’s with $m_{i-1}<m_{i}$ are the folding positions of $\bar{p}$. This alcove walk has the property that if $k\notin\{m_{1},\ldots,m_{t}\}$ then $p_{j_{k}}\in\{c_{j_{k}}^{\pm}\}$ and $p_{j_{m_{i}}}=f_{j_{m_{i}}}^{\epsilon^{\prime}_{m_{i}}}$ ($1\leq i\leq t$). In addition $\varphi(\bar{p}g)=wr_{\mathcal{M}}$ and $\mathrm{wt}(\bar{p}g)=w\hat{r}_{\mathcal{M}}(\lambda)$. Let $\mathcal{M}_{i}=\{h_{m_{1}},\ldots,h_{m_{i}}\}\subset\mathcal{H}=\{h_{1},\ldots,h_{s}\}$ for $1\leq i\leq t$ with the convention that $\mathcal{M}_{0}=\emptyset$ and $r_{\emptyset}=id$. A key point is that $wr_{\mathcal{M}_{i-1}}>wr_{\mathcal{M}_{i}}$ if and only if the folding orientations satisfy $\epsilon_{m_{i}}=\epsilon^{\prime}_{m_{i}}$. (The proof uses the condition that if $\beta>0$, then $ws_{\beta}>w$ if and only if $w(\beta)>0$.) All this put together implies that $\{h_{m_{1}},\ldots,h_{m_{i}}\}\leftrightarrow\phi_{m_{i}}\cdots\phi_{m_{1}}(p_{0})$ gives the bijection in the statement. ∎
For a subset $\mathcal{M}\subset\mathcal{H}$ as in Lemma 2.7, set $p(\mathcal{M})$ to be the alcove walk $\bar{p}g$ associated to $\mathcal{M}$, and set $f^{+}(\mathcal{M})=f^{+}(p(\mathcal{M}))$.
The following is a reformulation of Ram’s result from Theorem 2.4 in terms of paths in the Bruhat order.
Corollary 2.9.
Let $u,w\in W$ and $\lambda,\mu\in X^{*}(T)$. Assume the same hypotheses and notation as in Lemma 2.7. Then
$c_{u,\mu}^{w,\lambda}=\sum_{\mathcal{M}}(-1)^{f^{+}(\mathcal{M})}(1-q)^{|\mathcal{M}|}q^{\frac{\ell(u)-\ell(w)-|\mathcal{M}|}{2}},$
where the sum is over subsets $\mathcal{M}\subset\mathcal{H}$ which satisfy $w\stackrel{{\scriptstyle\mathcal{M}}}{{\Longrightarrow}}u=wr_{\mathcal{M}}$ and $\mu=w\hat{r}_{\mathcal{M}}(\lambda)$.
Remark 2.10.
A priori, the coefficients $c_{u,\mu}^{w,\lambda}$ appearing in the walk algebra are in ${\mathbb{Z}}[q^{\pm\frac{1}{2}}]$. However, after matching these with the initial definition of Hecke algebras, it turns out that $c_{u,\mu}^{w,\lambda}\in{\mathbb{Z}}[q^{\pm 1}]$. This can also be seen directly in the formula above, by observing that $\ell(u)-\ell(w)-|\mathcal{M}|$ is even. (This uses that a reflection has odd length, and the cancellation property of non-reduced expressions.)
3. A $\lambda$-chain formula for the transition coefficients $c_{u,\mu}^{w,\lambda}$
In this section, we reformulate the alcove walk formula 2.9 in terms of the notion of $\lambda$-chains introduced in [LP07]. This notion was utilized to obtain a $K$-theory Chevalley formula for the structure sheaves of Schubert varieties. The main results of this section are Theorem 3.9 and Theorem 3.10. Specializing $y\mapsto 0$, these recover [LP07, Theorem 6.1]. Throughout this section, we only consider alcoves on the identity sheet $\mathfrak{t}_{{\mathbb{R}}}^{*}$. Recall that $\operatorname{A_{\circ}}+\lambda:=\{x+\lambda\mid x\in\operatorname{A_{\circ}}\}$ is the alcove on $\mathfrak{t}_{\mathbb{R}}^{*}$. (If $\lambda$ is not a root, the alcove $t_{\lambda}(\operatorname{A_{\circ}})$ is not the alcove $\operatorname{A_{\circ}}+\lambda$.)
3.1. Alcove paths and $\lambda$-chains
For any two alcoves $A$ and $B$, which are separated by a common wall lying on a hyperplane $H_{\beta,k}$, write $A\stackrel{{\scriptstyle\beta}}{{\longrightarrow}}B$ if the root $\beta$ points from $A$ to $B$.
Definition 3.1.
[LP07, Definition 5.2] An alcove path is a sequence of alcoves $(A_{0},A_{1},\ldots,A_{l})$ such that $A_{j-1}$ and $A_{j}$ are adjacent, $1\leq j\leq l$. We denote this by
$A_{0}\stackrel{{\scriptstyle-\beta_{1}}}{{\longrightarrow}}A_{1}\stackrel{{\scriptstyle-\beta_{2}}}{{\longrightarrow}}\cdots\stackrel{{\scriptstyle-\beta_{l}}}{{\longrightarrow}}A_{l}.$
When the length $l$ is minimal among all alcove paths from $A_{0}$ to $A_{l}$, it is called a reduced alcove path.
Remark 3.2.
By definition, there is a one-to-one correspondence between alcove paths
(15) $\operatorname{A_{\circ}}=A_{0}\stackrel{{\scriptstyle-\beta_{1}}}{{\longrightarrow}}A_{1}\stackrel{{\scriptstyle-\beta_{2}}}{{\longrightarrow}}\cdots\stackrel{{\scriptstyle-\beta_{l}}}{{\longrightarrow}}A_{l}=\operatorname{A_{\circ}}-\lambda$
from $\operatorname{A_{\circ}}$ to $A_{l}=v_{-\lambda}(\operatorname{A_{\circ}})$ and alcove walks from $A_{0}$ to $A_{l}$ of the form $c^{\epsilon_{1}}_{i_{1}}c^{\epsilon_{2}}_{i_{2}}\cdots c^{\epsilon_{l}}_{i_{l}}$ , where $-\beta_{j}\in R^{\epsilon_{j}}$ for $1\leq j\leq l$.
Recall that $s_{0}=s_{\theta,1}$ is the affine reflection along the hyperplane $H_{\theta,1}$ with $\theta^{\vee}$ the highest coroot. Define $\alpha_{0}=-\theta$, $\bar{s}_{0}=s_{\theta}$ and $\bar{s_{i}}=s_{i}$ for $1\leq i\leq r$.
Lemma 3.3.
[LP07, Lemma 5.3] For any $v\in W_{\operatorname{aff}}$, there is a one-to-one correspondence between decompositions of $v$ in the simple reflections $s_{i}\in W_{\operatorname{aff}}$ ($0\leq i\leq r$) and alcove paths from $\operatorname{A_{\circ}}$ to $v(\operatorname{A_{\circ}})$ as follows.
For any decomposition $v=s_{i_{1}}s_{i_{2}}\cdots s_{i_{l}}$, define
$\beta_{j}:=\bar{s}_{i_{1}}\bar{s}_{i_{2}}\cdots\bar{s}_{i_{j-1}}(\alpha_{i_{j}}),\quad 1\leq j\leq l.$
Then
$\operatorname{A_{\circ}}=A_{0}\stackrel{{\scriptstyle-\beta_{1}}}{{\longrightarrow}}A_{1}\stackrel{{\scriptstyle-\beta_{2}}}{{\longrightarrow}}\cdots\stackrel{{\scriptstyle-\beta_{l}}}{{\longrightarrow}}A_{l}=v(\operatorname{A_{\circ}})$
is an alcove path from $\operatorname{A_{\circ}}$ to $v(\operatorname{A_{\circ}})$.
Moreover, the affine reflection along the $j$-th hyperplane separating $A_{j-1}$ and $A_{j}$ is $s_{i_{1}}s_{i_{2}}\cdots s_{i_{j-1}}s_{i_{j}}s_{i_{j-1}}\cdots s_{i_{2}}s_{i_{1}}$; in particular, the separating wall is labeled by $i_{j}$. Under this correspondence, a reduced alcove path corresponds to a reduced decomposition for $v$.
Let $\lambda$ be an integral weight and let $v_{-\lambda}\in W_{\operatorname{aff}}$ be the affine Weyl group element which satisfies $v_{-\lambda}(\operatorname{A_{\circ}})=\operatorname{A_{\circ}}-\lambda$, i.e., $t_{-\lambda}=v_{-\lambda}g$ with $g\in\Omega$. Choose a (possibly non-reduced) decomposition $v_{-\lambda}=s_{i_{1}}\cdots s_{i_{l}}$ and let $\beta_{j}$ be defined by
$\beta_{j}:=\bar{s}_{i_{1}}\bar{s}_{i_{2}}\cdots\bar{s}_{i_{j-1}}(\alpha_{i_{j}}),\quad 1\leq j\leq l\/,$
with the convention that $\alpha_{0}=-\theta$ (see Lemma 3.3). Then
$\operatorname{A_{\circ}}=A_{0}\stackrel{{\scriptstyle-\beta_{1}}}{{\longrightarrow}}A_{1}\stackrel{{\scriptstyle-\beta_{2}}}{{\longrightarrow}}\cdots\stackrel{{\scriptstyle-\beta_{l}}}{{\longrightarrow}}\operatorname{A_{\circ}}-\lambda$
is an alcove path from $\operatorname{A_{\circ}}$ to $\operatorname{A_{\circ}}-\lambda$.
Definition 3.4.
[LP07, Definition 5.4] The sequence of roots $(\beta_{1},\ldots,\beta_{l})$ is a $\lambda$-chain of roots associated to the decomposition of $v_{-\lambda}$. A $\lambda$-chain is called reduced if the decomposition $v_{-\lambda}=s_{i_{1}}\cdots s_{i_{l}}$ is reduced.
Let $H_{-\beta_{j},d_{j}}$ be the hyperplane separating the alcoves $A_{j}$ and $A_{j+1}$. The sequence of integers $d_{j}$ are determined by the sequence of roots $\beta_{j}$, but we occasionally keep the information of $d$’s in the notation, and we refer to the sequence of pairs $(\beta_{1},d_{1}),(\beta_{2},d_{2}),\ldots,(\beta_{l},d_{l})$ as a $\lambda$-chain  [^3] If $\lambda$ is dominant, this definition was extended to the Kac–Moody situation in [LP08].. Following [LP07, Prop. 6.8] we recall a combinatorial construction of a $\lambda$-chain for an integral weight $\lambda$.
Fix a linear order on the index of Dynkin nodes (for example $1<2<\cdots<r$ in $I=\{1,2,\ldots,r\}$). The set $\mathcal{R}_{\lambda}\subset W_{\mathrm{aff}}$ of affine reflections $s_{\alpha,k}$ for the hyperplanes $H_{\alpha,k}$ separating the fundamental alcove $\operatorname{A_{\circ}}$ and $\operatorname{A_{\circ}}-\lambda$ is given by
$\mathcal{R}_{\lambda}=\bigcup_{\alpha\in R^{+}}\begin{cases}\{s_{\alpha,k}:0\geq k>-\langle\lambda,\alpha^{\vee}\rangle\}&\textrm{ if }\langle\lambda,\alpha^{\vee}\rangle>0\\ \{s_{\alpha,k}:0<k\leq-\langle\lambda,\alpha^{\vee}\rangle\}&\textrm{ if }\langle\lambda,\alpha^{\vee}\rangle<0\\ \emptyset&\textrm{ otherwise }\end{cases}$
One defines a ‘height function’
$h:\mathcal{R}_{\lambda}\to\mathbb{R}^{r+1}\/;\quad h(s_{\alpha,k})=\frac{1}{\langle\lambda,\alpha^{\vee}\rangle}(-k,\langle\varpi_{1},\alpha^{\vee}\rangle,\cdots,\langle\varpi_{r},\alpha^{\vee}\rangle)\/.$
It turns out that this function is injective. Now order the images of $h$ in lexicographic order, so we obtain $h(s_{\alpha_{1},k_{1}})<\ldots<h(s_{\alpha_{l},k_{l}})$. Define another function $b:\mathcal{R}_{\lambda}\to R^{+}\cup R^{-}$ by
$b(s_{\alpha,k})=\begin{cases}\alpha&\textrm{ if }k\leq 0,\alpha\in R^{+}\/;\\ -\alpha&\textrm{ if }k>0,\alpha\in R^{+}\/.\end{cases}$
Then $b(s_{\alpha_{1},k_{1}}),\ldots,b(s_{\alpha_{l},k_{l}})$ is a (reduced) $\lambda$-chain of roots.
Remark 3.5.
A particularly nice situation occurs for a minuscule fundamental weight $\varpi_{i}$, i.e. when $\langle\varpi_{i},\alpha^{\vee}\rangle\in\{0,1\}$ for any positive root $\alpha$. In this case all $k_{i}=0$ and $v_{-\varpi_{i}}=(w^{P_{i}})^{-1}\in W$, with $w^{P_{i}}$ being the longest minimal length representative for cosets of $W/W_{P_{i}}$, where $W_{P_{i}}=\operatorname{Stab}_{W}(\varpi_{i})$; equivalently, the Schubert variety $X(w^{P_{i}}W_{P_{i}})=G/P_{i}$. A reduced decomposition and the associated $\varpi_{i}$-chain may be read from the associated Young poset of $G/P_{i}$ see e.g. [Pro99, Ste01] and also [BCMP18, §3.1] and Example 3.6 below. It may also be obtained as a reverse linear extension of the heap $H(w^{P_{i}})$, and this gives a one-to-one correspondence between reduced $\varpi_{i}$-chains and reverse linear extensions of the heap $H(w^{P_{i}})$. We refer the readers to [MNS22a, NO19] for the heap perspective.
Example 3.6.
Consider $G=\mathrm{SL}_{5}$ and the fundamental weight $\varpi_{2}$. The stabilizer is the maximal parabolic $P_{2}$ so that $G/P_{2}$ is the Grassmannian $\mathrm{Gr}(2,5)$. The inversion set and a reduced decomposition of $v_{-\varpi_{2}}$ may be read from the Young diagrams below, with the notation $(i-j)=\alpha_{i}+\ldots+\alpha_{j-1}$.
[…]
Then $v_{-\varpi_{2}}=s_{2}s_{3}s_{4}s_{1}s_{2}s_{3}$ and a $\varpi_{2}$-chain of roots is given by $\{\alpha_{2},\alpha_{2}+\alpha_{3},\alpha_{2}+\alpha_{3}+\alpha_{4},\alpha_{1}+\alpha_{2},\alpha_{1}+\alpha_{2}+\alpha_{3},\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}\}$.
Example 3.7.
Let $G=\operatorname{SL}_{3}$, with the Weyl group $W=S_{3}$, and consider $\lambda=2\varpi_{1}-2\varpi_{2}$. An example of alcove path from $\operatorname{A_{\circ}}$ to $\operatorname{A_{\circ}}-\lambda$ is
$\operatorname{A_{\circ}}=A_{0}\stackrel{{\scriptstyle-\beta_{1}}}{{\longrightarrow}}A_{1}\stackrel{{\scriptstyle-\beta_{2}}}{{\longrightarrow}}A_{2}\stackrel{{\scriptstyle-\beta_{3}}}{{\longrightarrow}}A_{3}\stackrel{{\scriptstyle-\beta_{4}}}{{\longrightarrow}}A_{4}\stackrel{{\scriptstyle-\beta_{5}}}{{\longrightarrow}}A_{5}\stackrel{{\scriptstyle-\beta_{6}}}{{\longrightarrow}}A_{6}=\operatorname{A_{\circ}}-\lambda\;(A_{i}=r_{i}A_{i-1},1\leq i\leq 6),$
which is the red path below, and it corresponds to the decomposition $v_{-\lambda}=s_{0}s_{1}s_{0}s_{1}s_{2}s_{1}$. The corresponding alcove walk is $\bar{p}=c_{0}^{+}c_{1}^{+}c_{0}^{-}c_{1}^{-}c_{2}^{-}c_{1}^{+}$, and the corresponding $\lambda$-chain of roots is $(\beta_{1},1),(\beta_{2},1),(\beta_{3},0),(\beta_{4},-1),(\beta_{5},1),(\beta_{6},2)$, as calculated below. Here we included the $d$’s in the notation.
[…]
$\begin{array}[]{lll}i_{1}=0,&\beta_{1}=\alpha_{0}=-(\alpha_{1}+\alpha_{2})\\ i_{2}=1,&\beta_{2}=\bar{s_{0}}(\alpha_{1})=-\alpha_{2}\\ i_{3}=0,&\beta_{3}=\bar{s_{0}}\bar{s_{1}}(\alpha_{0})=\alpha_{1}\\ i_{4}=1,&\beta_{4}=\bar{s_{0}}\bar{s_{1}}\bar{s_{0}}(\alpha_{1})=\alpha_{1}+\alpha_{2}\\ i_{5}=2,&\beta_{5}=\bar{s_{0}}\bar{s_{1}}\bar{s_{0}}\bar{s_{1}}(\alpha_{2})=\alpha_{1}\\ i_{6}=1,&\beta_{6}=\bar{s_{0}}\bar{s_{1}}\bar{s_{0}}\bar{s_{1}}\bar{s_{2}}(\alpha_{1})=-\alpha_{2}\\ \end{array}$
$s_{0}=\hat{r}_{H_{-\beta_{1},1}}$
$s_{0}s_{1}s_{0}=\hat{r}_{H_{-\beta_{2},1}}$
$(s_{0}s_{1})s_{0}(s_{1}s_{0})=s_{1}=\hat{r}_{H_{-\beta_{3},0}}$
$(s_{0}s_{1}s_{0})s_{1}(s_{0}s_{1}s_{0})=s_{0}=\hat{r}_{H_{-\beta_{4},-1}}$
$(s_{0}s_{1}s_{0}s_{1})s_{2}(s_{1}s_{0}s_{1}s_{0})=\hat{r}_{H_{-\beta_{5},1}}$
$(s_{0}s_{1}s_{0}s_{1}s_{2})s_{1}(s_{2}s_{1}s_{0}s_{1}s_{0})=\hat{r}_{H_{-\beta_{6},2}}$
We can choose another $\lambda$-chain for $\lambda=2\varpi_{1}-2\varpi_{2}$ by using the reduced decomposition $v_{-\lambda}=s_{1}s_{0}s_{2}s_{1}$. The corresponding reduced $\lambda$-chain is $(\alpha_{1},0),(-\alpha_{2},1),(\alpha_{1},1),(-\alpha_{2},2)$.
Lemma 3.8.
[LP07, Remark 5.5] Let $L=(\beta_{1},\ldots,\beta_{l})$ be a $\lambda$-chain. Then $\bar{L}:=(-\beta_{l},\ldots,-\beta_{1})$ is a $(-\lambda)$-chain. If $H_{-\beta_{j},d_{j}}$ is the $j$-th hyperplane of the alcove path from $\operatorname{A_{\circ}}$ to $\operatorname{A_{\circ}}-\lambda$ determined by $L$, then the $j$-th hyperplane of alcove path from $\operatorname{A_{\circ}}$ to $\operatorname{A_{\circ}}+\lambda$ determined by $\bar{L}$ is $H_{\beta_{l+1-j},\langle\lambda,\beta_{l+1-j}^{\vee}\rangle-d_{l+1-j}}$. If $L$ is reduced, then $\bar{L}$ is also reduced.
3.2. $\lambda$-chain formulae
Next we state the main theorem of this section. For any root hyperplane $h=H_{\beta,k}$, let $r_{h}$ denote the reflection along the hyperplane $H_{\beta,0}$, and $\hat{r}_{h}$ be the reflection along $h$.
Assume $\lambda$ is an integral weight and fix a reduced $\lambda$-chain $(\beta_{1},\beta_{2},\ldots,\beta_{l})$, which corresponds to an alcove walk from $\operatorname{A_{\circ}}$ to $\operatorname{A_{\circ}}-\lambda$, with separating hyperplanes $H_{-\beta_{j},d_{j}}=:h_{j}$.
For a subset $J=\{j_{1}<j_{2}<\cdots<j_{t}\}\subset\{1,2,\ldots,l\}$, define the following relation
$u\stackrel{{\scriptstyle J_{>}}}{{\longrightarrow}}w\stackrel{{\scriptstyle def.}}{{\iff}}u<ur_{h_{j_{t}}}<ur_{h_{j_{t}}}r_{h_{j_{t-1}}}<\cdots<ur_{h_{j_{t}}}\cdots r_{h_{j_{1}}}=w,$
and let
(16) $\hat{r}_{J_{<}}:=\hat{r}_{h_{j_{1}}}\cdots\hat{r}_{h_{j_{t}}}.$
Hence, $w\stackrel{{\scriptstyle J}}{{\Longrightarrow}}u$ from Equation 14 is equivalent to $u\stackrel{{\scriptstyle J_{>}}}{{\longrightarrow}}w$. Therefore, 2.9 can be restated as follows.
Theorem 3.9.
In the above setting,
(17) $c_{u,\mu}^{w,-\lambda}=\sum(-1)^{n(J)}(1-q)^{|J|}q^{\frac{\ell(u)-\ell(w)-|J|}{2}},$
where the sum is over subsets $J\subset\{1,2,\ldots,l\}$ such that $u\stackrel{{\scriptstyle J_{>}}}{{\longrightarrow}}w$ and $w\hat{r}_{J_{<}}(-\lambda)=\mu$, and where $n(J):=\#\{j\in J\mid\beta_{j}<0\}$.
Proof.
In the summation in 2.9, we let $J\subset\{1,2,\ldots,l\}$ be the indices of the hyperplanes in ${\mathcal{M}}$. Then by Remark 3.2, $f^{+}({\mathcal{M}})=n(J)$. This finishes the proof. ∎
Using the transformation between a $\lambda$-chain and a $(-\lambda)$-chain, we can get rid of the negative sign in front of $\lambda$ in Theorem 3.9 as follows.
First of all, $(-\beta_{l},-\beta_{l-1},\ldots,-\beta_{1})$ is a $(-\lambda)$-chain, which corresponds to an alcove walk from $\operatorname{A_{\circ}}$ to $\operatorname{A_{\circ}}+\lambda$, with the $j$-th separating hyperplane being
(18) $h^{\prime}_{j}:=H_{\beta_{l+1-j},\langle\lambda,\beta^{\vee}_{l+1-j}\rangle-d_{l+1-j}}\/.$
Then $r_{h^{\prime}_{j}}=r_{h_{l+1-j}}$. Let $\tilde{r}_{h_{l+1-j}}:=\hat{r}_{h^{\prime}_{j}}$ be the reflection along $h^{\prime}_{j}$. Define
$u\stackrel{{\scriptstyle J_{<}}}{{\longrightarrow}}w\stackrel{{\scriptstyle def.}}{{\iff}}u<ur_{h_{j_{1}}}<ur_{h_{j_{1}}}r_{h_{j_{2}}}<\cdots<ur_{h_{j_{1}}}\cdots r_{h_{j_{t}}}=w\/,$
and
(19) $\tilde{r}_{J_{>}}:=\tilde{r}_{h_{j_{t}}}\cdots\tilde{r}_{h_{j_{1}}}.$
Theorem 3.10.
The following holds,
(20) $c_{u,\mu}^{w,\lambda}=\sum(-1)^{n(J)}(q-1)^{|J|}q^{\frac{\ell(u)-\ell(w)-|J|}{2}},$
where the sum is over subsets $J\subset\{1,2,\ldots,l\}$ such that $u\stackrel{{\scriptstyle J_{<}}}{{\longrightarrow}}w$ and $w\tilde{r}_{J_{>}}(\lambda)=\mu$, where $n(J):=\#\{j\in J\mid\beta_{j}<0\}$.
Proof.
Applying Theorem 3.9 to the $(-\lambda)$-chain $(-\beta_{l},-\beta_{l-1},\ldots,-\beta_{1})$, we get
$c_{u,\mu}^{w,\lambda}=\sum(-1)^{\#\{j_{i}\mid-\beta_{l+1-j_{i}}<0\}}(1-q)^{t}q^{\frac{\ell(u)-\ell(w)-t}{2}},$
where the summation is over subsets $J^{\prime}:=\{1\leq j_{1}<j_{2}<\dots<j_{t}\leq l\}$, such that
(21) $u<ur_{h^{\prime}_{j_{t}}}<\cdots<ur_{h^{\prime}_{j_{t}}}\cdots r_{h^{\prime}_{j_{1}}}=w$
and
(22) $w\hat{r}_{h^{\prime}_{j_{1}}}\cdots\hat{r}_{h^{\prime}_{j_{t}}}(\lambda)=\mu.$
Let
$J:=\{1\leq l+1-j_{t}<l+1-j_{t-1}<\cdots<l+1-j_{1}\leq l\}.$
Then
$(-1)^{\#\{j_{i}\mid-\beta_{l+1-j_{i}}<0\}}(1-q)^{t}=(-1)^{n(J)}(q-1)^{t},$
where $n(J)$ is defined in Theorem 3.9. Since $r_{h^{\prime}_{j}}=r_{h_{l+1-j}}$, the condition Equation 21 is equivalent to $u\stackrel{{\scriptstyle J_{<}}}{{\longrightarrow}}w$. On the other hand, the condition Equation 22 is equivalent to $w\tilde{r}_{J_{>}}(\lambda)=\mu$ as $\tilde{r}_{h_{l+1-j}}=\hat{r}_{h^{\prime}_{j}}$. This finishes the proof. ∎
For further use, we also record the following technical result, which will allow to rewrite the elements $w\hat{r}_{J_{<}}(-\lambda)$ from Theorem 3.9 and $w\tilde{r}_{J_{<}}$ from Theorem 3.10.
Proposition 3.11.
Consider a $\lambda$-chain $\beta_{1},\ldots,\beta_{l}$. For any subsequence $J=\{j_{1}<j_{2}<\ldots<j_{t}\}\subset\{1,2,\ldots,l\}$, we have
$-\hat{r}_{J_{<}}(-\lambda)=r_{J}\;\tilde{r}_{J_{>}}(\lambda),\text{ where }{r}_{J}={r}_{h_{j_{1}}}{r}_{h_{j_{2}}}\cdots{r}_{h_{j_{t}}}.$
Proof.
Let $m_{j}=\langle\lambda,\beta_{j}^{\vee}\rangle$ ($1\leq j\leq l$). By induction on $|J|$, it is easy to show the following three equalities:
$\begin{array}[]{lcl}\hat{r}_{h_{j_{1}}}\hat{r}_{h_{j_{2}}}\cdots\hat{r}_{h_{j_{t}}}(-\lambda)&=&-\lambda+(m_{j_{1}}-d_{j_{1}})\beta_{j_{1}}+(m_{j_{2}}-d_{j_{2}})r_{h_{j_{1}}}\beta_{j_{2}}+\cdots\\ &&+(m_{j_{t}}-d_{j_{t}})r_{j_{1}}r_{h_{j_{2}}}\cdots r_{h_{j_{t-1}}}\beta_{j_{t}},\\ \tilde{r}_{h_{j_{t}}}\tilde{r}_{h_{j_{t-1}}}\cdots\tilde{r}_{h_{j_{1}}}(\lambda)&=&\lambda-d_{j_{t}}\beta_{j_{t}}-d_{j_{t-1}}r_{h_{j_{t}}}\beta_{j_{t-1}}-\cdots-d_{j_{1}}r_{h_{j_{t}}}r_{h_{j_{t-1}}}\cdots r_{h_{j_{2}}}\beta_{j_{1}},\\ r_{h_{j_{1}}}r_{h_{j_{2}}}\cdots r_{h_{j_{t}}}(\lambda)&=&\lambda-m_{j_{1}}\beta_{j_{1}}-m_{j_{2}}r_{h_{j_{1}}}\beta_{j_{2}}-\cdots-m_{j_{t}}r_{h_{j_{1}}}r_{h_{j_{2}}}\cdots r_{h_{j_{t-1}}}\beta_{j_{t}}.\end{array}$
As $r_{J}=r_{h_{j_{1}}}r_{h_{j_{2}}}\cdots r_{h_{j_{t}}}$ is a linear transformation, we get
$r_{J}\tilde{r}_{J_{>}}(\lambda)=r_{J}(\lambda)+d_{j_{t}}r_{h_{j_{1}}}\cdots r_{h_{j_{t-1}}}\beta_{j_{t}}+d_{j_{t-1}}r_{h_{j_{1}}}\cdots r_{h_{j_{t-2}}}\beta_{j_{t-1}}+\cdots+d_{j_{1}}\beta_{j_{1}}=-\hat{r}_{J_{<}}(-\lambda).$
Remark 3.12.
With the notation as above, the condition that $u\stackrel{{\scriptstyle J_{>}}}{{\longrightarrow}}w$ in Theorem 3.9 implies that $ur_{J}^{-1}=w$, thus by 3.11,
$w\hat{r}_{J_{<}}(-\lambda)=-u\tilde{r}_{J_{>}}(\lambda)\/.$
Similarly, in Theorem 3.10, we have that $w=ur_{J}$, thus
$w\tilde{r}_{J_{>}}(\lambda)=-u\hat{r}_{J_{<}}(-\lambda)\/.$
This leads to alternative formulae in the aforementioned theorems 3.9 and 3.10.
In the appendix §B below we included a fully worked out example illustrating calculations of some coefficients $c_{u,\mu}^{w,\lambda}$ utilizing Theorem 3.9 and Theorem 3.10.
4. Motivic Chern classes of Schubert cells
In this section, we recall basic properties of the motivic Chern classes of the Schubert cells in the (partial) flag varieties.
4.1. Definition of the Motivic Chern classes
Let $X$ be a quasi-projective complex algebraic variety, and let $G_{0}(var/X)$ be the (relative) Grothendieck group of varieties over $X$. It consists of isomorphism classes of morphisms $[f:Z\to X]$ modulo the scissor relations. Brasselet, Schürmann and Yokura [BSY10] defined the motivic Chern transformation $MC_{y}:G_{0}(var/X)\to K(X)[y]$ with values in the K-theory group of coherent sheaves in $X$ to which one adjoins a formal variable $y$. The transformation $MC_{y}$ is a group homomorphism, it is functorial with respect to proper push-forwards, and if $X$ is smooth it satisfies the normalization condition
$MC_{y}[id_{X}:X\to X]=\sum[\wedge^{j}T^{*}_{X}]y^{j}\/.$
Here $[\wedge^{j}T^{*}_{X}]$ is the K-theory class of the bundle of degree $j$ differential forms on $X$. As explained in [BSY10], the motivic Chern class is related by a Hirzebruch–Riemann–Roch type statement to the Chern–Schwartz–MacPherson (CSM) class in the homology of $X$; see Section A.2.
There is also an equivariant version of the motivic Chern class transformation, which uses equivariant varieties and morphisms, and has values in the suitable equivariant K-theory group. Its definition was given in [AMSS19, FRW21], following closely the approach of [BSY10].
Assume $X$ is smooth and there is a torus $T$ acting on $X$. Let $K_{T}(X)$ denote the equivariant K-theory of $X$, see [CG09]. If $X$ is a point, $K_{T}(pt)=K^{0}(\operatorname{Rep}(T))={\mathbb{Z}}[T]$. For any ${\mathcal{F}}\in K_{T}(X)$, let
$\chi_{T}(X,{\mathcal{F}}):=\sum_{i}(-1)^{i}H^{i}(X,{\mathcal{F}})\in K_{T}(pt).$
Let $\langle$-,$\rangle$ - , - ⟩ be the non-degenerate pairing on $K_{T}(X)$ defined by
$\langle{\mathcal{F}},{\mathcal{G}}\rangle=\chi_{T}(X,{\mathcal{F}}\otimes{\mathcal{G}})\in K_{T}(pt),$
where ${\mathcal{F}},{\mathcal{G}}\in K_{T}(X)$. For a vector bundle $E$, the $\lambda_{y}$-class of $E$ is the class
$\lambda_{y}(E):=\sum_{k}[\wedge^{k}E]y^{k}\in K_{T}(X)[y]\/.$
The $\lambda_{y}$-class is multiplicative, i.e. for any short exact sequence $0\to E_{1}\to E\to E_{2}\to 0$ of equivariant vector bundles there is an equality $\lambda_{y}(E)=\lambda_{y}(E_{1})\lambda_{y}(E_{2})$ as elements in $K_{T}(X)[y]$.
Recall the (relative) motivic Grothendieck group ${G}_{0}^{T}(var/X)$ of varieties over $X$ is the free abelian group generated by the isomorphism classes $[f:Z\to X]$ where $Z$ is a quasi-projective $T$-variety and $f:Z\to X$ is a $T$-equivariant morphism modulo the usual additivity relations
$[f:Z\to X]=[f:U\to X]+[f:Z\setminus U\to X]$
for $U\subset Z$ an open $T$-invariant subvariety. For any equivariant morphism $f:X\to Y$ of quasi-projective $T$-varieties there are well defined push-forwards $f_{!}:G_{0}^{T}(var/X)\to G_{0}^{T}(var/Y)$ given by composition. The equivariant motivic Chern class is defined by the following theorem.
Theorem 4.1.
[AMSS19, FRW21] Let $X$ be a quasi-projective, non-singular, complex algebraic variety with an action of the torus $T$. There exists a unique natural transformation $MC_{y}:G_{0}^{T}(var/X)\to K_{T}(X)[y]$ satisfying the following properties:
  1. (1)
    It is functorial with respect to $T$-equivariant proper morphisms of non-singular, quasi-projective varieties.
  2. (2)
    It satisfies the normalization condition
    $MC_{y}[id_{X}:X\to X]=\lambda_{y}(T^{*}_{X})=\sum y^{i}[\wedge^{i}T^{*}_{X}]_{T}\in K_{T}(X)[y]\/.$
The transformation $MC_{y}$ satisfies a Verdier–Riemann–Roch (VRR) formula: for any smooth, $T$-equivariant morphism $\pi:X\to Y$ of quasi-projective and non-singular algebraic varieties, and any $[f:Z\to Y]\in G_{0}^{T}(var/Y)$, the following holds:
$\lambda_{y}(T^{*}_{\pi})\cap\pi^{*}MC_{y}[f:Z\to Y]=MC_{y}[\pi^{*}f:Z\times_{Y}X\to X]\/.$
Define the Grothendieck–Serre dual operator ${\mathcal{D}}:K_{T}(X)\to K_{T}(X)$ as follows. For any $[{\mathcal{F}}]\in K_{T}(X)$, define
(23) ${\mathcal{D}}[{\mathcal{F}}]:=[RHom({\mathcal{F}},\omega^{\bullet}_{X})]:=[\omega_{X}^{\bullet}\otimes{\mathcal{F}}^{\vee}]\in K_{T}(X)\/,$
where $\omega^{\bullet}_{X}\simeq\omega_{X}[\dim X]$ is the (equivariant) dualizing complex of $X$ (the canonical bundle $\omega_{X}$ shifted by dimension). The class $[{\mathcal{F}}^{\vee}]$ is obtained by taking an equivariant resolution of ${\mathcal{F}}$ by vector bundles, and then taking duals. Extend the operators ${\mathcal{D}}$ and $(-)^{\vee}$ to $K_{T}(X)[y^{\pm 1}]$ by sending $y\mapsto y^{-1}$. We also let $(-)^{\vee}$ denote the operator on $K_{T}(pt)[y^{\pm 1}]$, which sends $e^{\lambda}$ to $e^{-\lambda}$ and $y$ to $y^{-1}$.
Definition 4.2.
Assume $\Omega\hookrightarrow X$ is a $T$-stable subvariety.
  1. (1)
    The motivic Chern (MC) class of $\Omega$ is
    $MC_{y}(\Omega):=MC_{y}[\Omega\hookrightarrow X]\in K_{T}(X)[y].$
  2. (2)
    If $\Omega$ is pure dimensional, the Segre motivic class (SMC) $SMC_{y}(\Omega)\in K_{T}(X)[[y]]$ is the class
    $SMC_{y}(\Omega):=(-y)^{\dim\Omega}\frac{{\mathcal{D}}(MC_{y}(\Omega))}{\lambda_{y}(T^{*}_{X})}\in K_{T}(X)[[y]]\/.$
4.2. Complete flag variety case
Both $\{MC_{y}(X(w)^{\circ})\mid w\in W\}$ and $\{SMC_{y}(Y(w)^{\circ})\mid w\in W\}$ are bases for the localized equivariant K-theory
$K_{T}(G/B)[[y]]_{\operatorname{loc}}:=K_{T}(G/B)[[y]]\otimes_{K_{T}(pt)}\operatorname{Frac}K_{T}(pt)\/.$
These classes can be calculated recursively using the Demazure–Lusztig operators as follows.
The left multiplication action of $G$ on $G/B$ induces a left Weyl group action on $K_{T}(G/B)$. For any $w\in W$, let $w^{L}$ denote this action.
Definition 4.3.
[MNS22b, Section 5.3] For any simple reflection $s_{i}$, we define the following left Demazure–Lusztig operator on $K_{T}(G/B)_{\operatorname{loc}}$:
(24) ${\mathcal{T}}^{L}_{i}:=\frac{1+ye^{-\alpha_{i}}}{1-e^{-\alpha_{i}}}s_{i}^{L}-\frac{1+y}{1-e^{-\alpha_{i}}}.$
These operators enjoy the following properties.
Lemma 4.4.
[MNS22b] The operators ${\mathcal{T}}^{L}_{i}$ satisfy the braid relation and the quadratic relation
$({\mathcal{T}}^{L}_{i}+1)({\mathcal{T}}^{L}_{i}+y)=0.$
Moreover, they commute with the operators of tensoring by elements in $K_{G}(G/B)$.
Moreover, it is easy to verify the following lemma.
Lemma 4.5.
The following map $\Psi$ defines an action of the affine Hecke algebra ${\mathbb{H}}$ (with $q=-y$) on $K_{T}(G/B)_{\operatorname{loc}}$,
$T_{i}\mapsto{\mathcal{T}}^{L}_{i},\textit{\quad and \quad}X^{\lambda}\mapsto e^{\lambda},$
where $e^{\lambda}\in K_{T}(\operatorname{pt})$ acts on $K_{T}(G/B)_{\operatorname{loc}}$, by multiplication.
We have the following theorem.
Theorem 4.6.
  1. (1)
    [MNS22b, Theorem 7.6] For $w\in W$ and a simple root $\alpha_{i}$,
    $\mathcal{T}^{L}_{i}(MC_{y}(X(w)^{\circ}))=\begin{cases}MC_{y}(X(s_{i}w)^{\circ})&\textrm{ if }s_{i}w>w;\\ -(y+1)MC_{y}(X(w)^{\circ})-yMC_{y}(X(s_{i}w)^{\circ})&\textrm{ if }s_{i}w<w\/.\end{cases}$
    In particular,
    $MC_{y}(X(w)^{\circ})={\mathcal{T}}^{L}_{w}([{\mathcal{O}}_{X(id)}]).$
  2. (2)
    [MNS22b, Theorem 7.1] For any $w,u\in W$:
    $\langle MC_{y}(X(w)^{\circ}),SMC_{y}(Y(u)^{\circ})\rangle=\delta_{u,w}\/.$
Remark 4.7.
  1. (1)
    By [AMSS22, Thm. 5.1 and Cor. 5.3],
    $MC_{0}(X(w)^{\circ})=[{\mathcal{O}}_{X(w)}(-\partial X(w))],\textit{\quad and \quad}SMC_{0}(Y(w)^{\circ})=[{\mathcal{O}}_{Y(w)}],$
    where $\partial X(w):=X(w)\setminus X(w)^{\circ}$. Thus, the duality in the second part of the theorem reduces to the classical fact
    $\langle[{\mathcal{O}}_{X(w)}(-\partial X(w))],[{\mathcal{O}}_{Y(u)}]\rangle=\delta_{u,w}.$
  2. (2)
    By definition, for any $w\in W$,
    $MC_{y}(Y(w))=\sum_{u\geq w}MC_{y}(Y(u)^{\circ}).$
    Besides, by the linearity of the Grothendieck–Serre dual operator ${\mathcal{D}}$,
    $SMC_{y}(Y(w))=\sum_{u\geq w}(-y)^{\ell(u)-\ell(w)}SMC_{y}(Y(u)^{\circ}).$
    Therefore,
    $SMC_{0}(Y(w))=[{\mathcal{O}}_{Y(w)}].$
4.3. Partial flag variety case
For any $w\in W$, let $\ell(wW_{P})$ denote the length of the minimal length representative in $wW_{P}$. Let $\pi:G/B\to G/P$ be the natural projection. It is proved in [AMSS19, Remeark 5.7] that
(25) $\pi_{*}MC_{y}(X(w)^{\circ})=(-y)^{\ell(w)-\ell(wW_{P})}MC_{y}(X(wW_{P})^{\circ})\/.$
In particular, if $w$ is a minimal length representative, then $\pi_{*}MC_{y}(X(w)^{\circ})=MC_{y}(X(wW_{P})^{\circ})$; this also follows directly from the functoriality property of the motivic classes.
Proposition 4.8.
  1. (1)
    [MNS22b, Proposition 6.3] Let $\Omega\subset G/P$ be a $T$-stable subvariety of $G/P$ and $\pi:G/B\to G/P$ be the projection. Then
    $\pi^{*}SMC_{y}(\Omega)=SMC_{y}(\pi^{-1}\Omega)\in K_{T}(G/B)[[y]]\/.$
  2. (2)
    [MNS22b, Theorem 7.2] Let $u,w\in W^{P}$. The Segre motivic classes are dual to the motivic Chern classes for any $G/P$, i.e.
    $\langle MC_{y}(X(wW_{P})^{\circ}),SMC_{y}(Y(uW_{P})^{\circ})\rangle=\delta_{u,w}\/.$
Remark 4.9.
By the first property in the proposition, for any $w\in W^{P}$,
$\pi^{*}(SMC_{y}(Y(wW_{P})))=SMC_{y}(Y(w)).$
Letting $y=0$, we get
$\pi^{*}(SMC_{0}(Y(wW_{P})))=SMC_{0}(Y(w))=[{\mathcal{O}}_{Y(w)}],$
where the second equality follows from Remark 4.7. By the definition of the SMC class,
(26) $SMC_{y}(Y(wW_{P}))=\sum_{u\in W^{P},u\geq w}(-y)^{\ell(u)-\ell(w)}SMC_{y}(Y(uW_{P})^{\circ}).$
Hence,
$\pi^{*}(SMC_{0}(Y(wW_{P})^{\circ}))=\pi^{*}(SMC_{0}(Y(wW_{P})))=[{\mathcal{O}}_{Y(w)}].$
Since $\pi^{*}([{\mathcal{O}}_{Y(wW_{P})}])=[{\mathcal{O}}_{Y(w)}]$ and $\pi^{*}$ is injective, we get
$SMC_{0}(Y(wW_{P})^{\circ})=SMC_{0}(Y(wW_{P}))=[{\mathcal{O}}_{Y(wW_{P})}].$
Combining with the second part of the proposition, we get
$MC_{0}(X(wW_{P})^{\circ})=[{\mathcal{O}}_{X(wW_{P})}(-\partial X(wW_{P}))].$
This can also be proved using Remark 4.7(1), Equation (25), and the fact that the pushforward of an ideal sheaf is an ideal sheaf, see [Bri02].
5. Chevalley formulae for the motivic Chern classes
In this section, we obtain several Chevalley formulae for the motivic Chern classes, in terms of alcove walks, $\lambda$-chains, and certain operators. The main technique is to reinterpret formulae from Hecke algebras such as Theorem 3.9 in terms of multiplications of motivic Chern classes by line bundles. Our main results are Theorem 5.5 and Theorem 5.12. In §5.3 we discuss several positivity properties and conjectures of the Chevalley coefficients. Finally, in §5.5 we discuss parabolic Chevalley formulae.
5.1. Chevalley coefficients
Consider a torus weight $\lambda\in X^{*}(T)$ and $u,w\in W$. The Chevalley coefficient $C_{u,\lambda}^{w}$ is defined by the following formula:
(27) ${\mathcal{L}}_{\lambda}\otimes MC_{y}(X(w)^{\circ})=\sum_{u\leq w}C_{u,\lambda}^{w}MC_{y}(X(u)^{\circ}).$
Note that for any simple reflection $s_{i}$ there is a short exact sequence of equivariant sheaves
$0\to\mathcal{L}_{\varpi_{i}}\otimes{\mathbb{C}}_{-w_{0}(\varpi_{i})}\to{\mathcal{O}}_{G/B}\to{\mathcal{O}}_{X(w_{0}s_{i})}\to 0$
with $\varpi_{i}$ the fundamental weight, see e.g., [BM15, §8]. Therefore,
$[{\mathcal{O}}_{X(w_{0}s_{i})}]=1-e^{-w_{0}(\omega_{i})}{\mathcal{L}}_{\varpi_{i}}\in K_{T}(G/B)\/,$
and the coefficients from (27) for $\lambda=\varpi_{i}$ also recover the multiplication of $[{\mathcal{O}}_{X(w_{0}s_{i})}]$ with the MC classes of the Schubert cells.
The coefficients $C_{u,\lambda}^{w}$ also arise from Chevalley formulae involving Segre motivic classes:
Lemma 5.1.
The following equation holds:
(28) ${\mathcal{L}}_{\lambda}\otimes SMC_{y}(Y(u)^{\circ})=\sum_{w\geq u}C_{u,\lambda}^{w}SMC_{y}(Y(w)^{\circ}).$
Proof.
This follows from the (Poincaré) duality in Theorem 4.6(2), as the Chevalley coefficients in Equation 27 and Equation 28 are given by
$C_{u,\lambda}^{w}=\langle{\mathcal{L}}_{\lambda}\otimes MC_{y}(X(w)^{\circ}),SMC_{y}(Y(u)^{\circ})\rangle\/.$
We now relate the Chevalley coefficients above to the coefficients $c_{u,\mu}^{w,\lambda}$ from Equation 7 in the Hecke algebra.
Theorem 5.2.
Let $\lambda$ be any weight in $X^{*}(T)$. The following Chevalley formula holds in $K_{T}(G/B)[y]$:
${\mathcal{L}}_{-\lambda}\otimes MC_{y}(X(w)^{\circ})=\sum_{\mu\in X^{*}(T),u\in W}y^{\ell(w)-\ell(u)}e^{-\mu}c_{u,\mu}^{w,\lambda}|_{q=-y}MC_{y}(X(u)^{\circ}).$
In particular, the following equation holds for the Chevalley coefficients in (27):
(29) $C_{u,-\lambda}^{w}=\sum_{\mu\in X^{*}(T)}y^{\ell(w)-\ell(u)}e^{-\mu}c_{u,\mu}^{w,\lambda}|_{q=-y}.$
From Equation 27, the Chevalley coefficients $C_{u,\lambda}^{v}$ belong to a localization $K_{T}(pt)[y]$ which allows division by $1+e^{\alpha}y$ for any root $\alpha$. However, the expansion from Equation 7 implies that the coefficients $(-q)^{\ell(w)-\ell(v)}c_{u,\mu}^{w,\lambda}$ are polynomials in ${\mathbb{Z}}[q]$. Then it follows that $C_{u,\lambda}^{v}$ are in fact polynomials in $K_{T}(pt)[y]$. This will be seen explicitly in Theorem 5.4 below.
Proof of Theorem 5.2.
Applying the map $\Psi$ in Lemma 4.5 to Equation 7, we get
(30) ${\mathcal{T}}^{L}_{w}e^{-\lambda}=\sum_{\mu\in X^{*}(T),u\in W}y^{\ell(w)-\ell(u)}e^{-\mu}c_{u,\mu}^{w,\lambda}|_{q=-y}{\mathcal{T}}^{L}_{u}\in\operatorname{End}_{{\mathbb{C}}}K_{T}(G/B)[y].$
Then the theorem follows by applying both sides to $[{\mathcal{O}}_{X(id)}]$, and utilizing that
${\mathcal{T}}^{L}_{w}e^{-\lambda}([{\mathcal{O}}_{X(id)}])=$ ${\mathcal{T}}^{L}_{w}({\mathcal{L}}_{-\lambda}\otimes[{\mathcal{O}}_{X(id)}])$
$=$ ${\mathcal{L}}_{-\lambda}\otimes MC_{y}(X(w)^{\circ}).$
Here the first equality follows from ${\mathcal{L}}_{-\lambda}\otimes[{\mathcal{O}}_{X(id)}]=e^{-\lambda}[{\mathcal{O}}_{X(id)}]$, while the second one follows from Lemma 4.4 and Theorem 4.6. ∎
Remark 5.3.
The above argument can be generalized to the case when the line bundle ${\mathcal{L}}_{\lambda}$ is replaced by any homogeneous bundle $\mathcal{V}=G\times^{B}V\to G/B$ associated to a $B$-representation of $V$. If the character of $V$ is $ch(V)=\sum_{\lambda}a_{\lambda}e^{\lambda}$, then a localization argument shows that the class of $\mathcal{V}$ in $K_{T}(G/B)$ is equal to
$[\mathcal{V}]=\sum a_{\lambda}\mathcal{L}_{\lambda}\/.$
It follows that for any $w\in W$,
$\mathcal{V}\otimes MC(X(w)^{\circ})=\sum_{\lambda}a_{\lambda}\mathcal{L}_{\lambda}\otimes MC(X(w)^{\circ})=\sum_{u}\sum_{\lambda}a_{\lambda}C_{u,\lambda}^{w}MC(X(u)^{\circ})\/.$
We illustrate this for $G/B=\mathrm{Fl}(n)$, the complete flag manifold. This is equipped with the tautological sequence $\mathcal{F}_{0}=0\subset\mathcal{F}_{1}\subset\mathcal{F}_{2}\subset\ldots\subset\mathcal{F}_{n}={\mathbb{C}}^{n}$. For $11\leq i\leq n-$ 1 define $X_{i}=\mathcal{F}_{i}/\mathcal{F}_{i-1}$ regarded in $K_{T}(G/B)$. Then
$\wedge^{j}\mathcal{F}_{i}=e_{j}(X_{1},\ldots,X_{i})\/,\quad Sym^{j}\mathcal{F}_{i}=h_{j}(X_{1},\ldots,X_{i})$
where $e_{j}$ and $h_{j}$ denote the elementary symmetric function, respectively the complete homogeneous symmetric function. Note that if $\varpi_{i}$ denotes the $i$th fundamental weight, then $X_{i}=\mathcal{L}_{\varpi_{i}-\varpi_{i-1}}$ for $11\leq i\leq n-$ 1, with the convention that $\varpi_{0}=0$. Theorem 5.2 gives a formula for the multiplication by monomials $X_{1}^{a_{1}}\cdot\ldots\cdot X_{n-1}^{a_{n-1}}$, which in turn gives formulae to multiply by $e_{j}(X_{1},\ldots,X_{i})$ and $h_{j}(X_{1},\ldots,X_{i})$.
In the next section, we give explicit formulae for the Chevalley coefficients $C_{\lambda,w}^{u}$ based on the formulae for the Hecke algebra coefficients $c_{u,\mu}^{w,\lambda}$.
5.2. Chevalley formulae via alcove walks and the $\lambda$-chains
Let us recall the setting of 2.9. For $\lambda\in X^{*}(T)$, choose a minimal length alcove walk $p_{v_{\lambda}}=c_{j_{1}}^{\epsilon_{1}}c_{j_{2}}^{\epsilon_{2}}\cdots c_{j_{s}}^{\epsilon_{s}}$ from $\operatorname{A_{\circ}}$ to $\operatorname{A_{\circ}}+\lambda$, and let =$\mathcal{H}=\{h_{1},h_{2},\ldots,h_{s}\}$ be the ordered sequence of hyperplanes defined by the walls of alcoves crossed by $p_{\lambda}$. Define $\beta_{i}\in R^{\epsilon_{i}}$ and $k_{i}\in\mathbb{Z}$ ($1\leq i\leq s$) by the condition $h_{i}=H_{\beta_{i},k_{i}}$. Combining 2.9 and Theorem 5.2, we get the following formula.
Theorem 5.4 (Alcove walk formula for the Chevalley coefficients).
$C_{u,-\lambda}^{w}=\sum_{\mathcal{M}}(-1)^{f^{+}(\mathcal{M})}(-1-y)^{|\mathcal{M}|}(-y)^{\frac{\ell(w)-\ell(u)-|\mathcal{M}|}{2}}{e^{-w\hat{r}_{\mathcal{M}}(\lambda)}}$
where the sum is over ordered subsets $\mathcal{M}\subset\mathcal{H}$ such that $w\stackrel{{\scriptstyle\mathcal{M}}}{{\Longrightarrow}}u=wr_{\mathcal{M}}$.
We now recall the setup of Theorem 3.9. Assume $\lambda$ is an integral weight and fix a reduced $\lambda$-chain $(\beta_{1},\beta_{2},\ldots,\beta_{l})$, which corresponds to an alcove walk from $\operatorname{A_{\circ}}$ to $\operatorname{A_{\circ}}-\lambda$, with separating hyperplanes $h_{j}:=H_{-\beta_{j},d_{j}}$.
Combining Theorem 3.9 , Theorem 3.10, and Theorem 5.2, we get the following formulae.
Theorem 5.5 ($\lambda$-chain formula for the Chevalley coefficients).
The following hold:
(31) $C_{u,-\lambda}^{w}=\sum_{J\subset\{1,2,\ldots,l\}}(-1)^{n(J)}(1+y)^{|J|}(-y)^{\frac{\ell(w)-\ell(u)-|J|}{2}}e^{-w\tilde{r}_{J_{>}}(\lambda)},$
where the sum is over subsets $J=\{j_{1}<\ldots<j_{t}\}\subset\{1,2,\ldots,l\}$ such that $u<ur_{h_{j_{1}}}<ur_{h_{j_{1}}}r_{h_{j_{2}}}<\ldots<ur_{h_{j_{1}}}r_{h_{j_{2}}}\cdot\ldots\cdot r_{h_{j_{t}}}=w$ and $\tilde{r}_{J_{>}}$ is defined in (19). Furthermore,
(32) $C_{u,\lambda}^{w}=\sum_{J\subset\{1,2,\ldots,l\}}(-1)^{n(J)}(-1-y)^{|J|}(-y)^{\frac{\ell(w)-\ell(u)-|J|}{2}}e^{-w\hat{r}_{J_{<}}(-\lambda)},$
where the sum is over subsets $J=\{j_{1}<\ldots<j_{t}\}\subset\{1,2,\ldots,l\}$ such that $u<ur_{h_{j_{t}}}<ur_{h_{j_{t}}}r_{h_{j_{t-1}}}<\ldots<ur_{h_{j_{t}}}\cdot\ldots\cdot r_{h_{j_{1}}}=w$ and $\hat{r}_{J_{<}}$ is defined in (16).
Remark 5.6.
(1) It follows from Remark 3.12 that $-w\tilde{r}_{J_{>}}(\lambda)=u\hat{r}_{J_{<}}(-\lambda)$ in Equation 31, and that $-w\hat{r}_{J_{<}}(-\lambda)=u\tilde{r}_{J_{>}}(\lambda)$ in Equation 32, giving alternative ways to calculate these.
(2) By Remark 4.7 (1), the MC and SMC classes specialize at $y=0$ to the ideal sheaves and the structure sheaves, respectively. Under this specialization, and using that $-w\tilde{r}_{J_{>}}(\lambda)=u\hat{r}_{J_{<}}(-\lambda)$, Equation 31 reduces to the equivariant $K$-theory Chevalley formula of Lenart–Postnikov [LP07, Theorem 6.1].
One can also consider the more general situation of Kac–Moody flag varieties defined e.g. in [Kum02]. These are ind-varieties, and one can define the motivic Chern classes of the finite-dimensional Schubert cells using the ind-structure. There are analogues of the (left and right) Demazure operators, and the notion of $\lambda$-chains extends to this setting, by results of Lenart and Postnikov [LP08]. Note that for an infinite Weyl group $W$, a $\lambda$-chain may be an infinite sequence, i.e. $l=\infty$. But, for given $u\leq w$, $t=|J|$ is finite, and the number of $J$ which satisfies the condition in Theorem 5.5 is also finite. Based on these similarities to the finite case, we expect the following conjecture to hold.
Conjecture 1.
For a Kac–Moody Weyl group $W$ and a dominant integral weight $\lambda$, the analogues of the equations Equation 31 and Equation 32 hold.
5.3. Miscelaneous
In this section we discuss positivity properties of the coefficients from the Chevalley formula, and some special properties of the multiplication by line bundles given by minuscule weights. We start with the following consequence of Theorem 5.5.
Proposition 5.7.
Let $\lambda$ be a dominant weight and set $q=-y$. Then:
(a) $C_{u,\lambda}^{w}$ may be written as a combination of terms of the form $e^{\mu}q^{a}(q-1)^{b}$ with non-negative integer coefficients.
(b) $C_{u,-\lambda}^{w}$ may be written as a combination of terms of the form $(-1)^{b}e^{\mu}q^{a}(q-1)^{b}$ with non-negative integer coefficients.
Furthermore, in both situations $b$ has the same parity as $\ell(w)-\ell(u)$.
Proof.
Both statements follow from Theorem 5.5, using that for a reduced $\lambda$-chain with $\lambda$ dominant we have $n(J)=0$, and that $\ell(w)-\ell(u)-|J|$ is an even integer. ∎
Example 5.8.
Consider type $A_{2}$, with $\lambda=2\varpi_{1}+\varpi_{2}$, and with the $\lambda$-chain from §B. Take $w=s_{2}s_{1}$. Then from Theorem 5.5 we get (with $q=-y$):
$\begin{split}{\mathcal{L}}_{\lambda}\otimes MC_{-q}(X(s_{2}s_{1})^{\circ})&=e^{\varpi_{1}-3\varpi_{2}}MC_{-q}(X(s_{2}s_{1})^{\circ})\\ &+(q-1)(e^{-\varpi_{2}}+e^{-\varpi_{1}+\varpi_{2}}+e^{-2\varpi_{1}+3\varpi_{2}})MC_{-q}(X(s_{1})^{\circ})\\ &+(q-1)(e^{2\varpi_{1}-2\varpi_{2}}+e^{3\varpi_{1}-\varpi_{2}})MC_{-q}(X(s_{2})^{\circ})\\ &+(q-1)^{2}(e^{2\varpi_{1}+\varpi_{2}}+e^{2\varpi_{2}}+e^{\varpi_{1}})MC_{-q}(X(id)^{\circ}).\end{split}$
$\begin{split}{\mathcal{L}}_{-\lambda}\otimes MC_{-q}(X(s_{2}s_{1})^{\circ})&=e^{-\varpi_{1}+3\varpi_{2}}MC_{-q}(X(s_{2}s_{1})^{\circ})\\ &-(q-1)(e^{\varpi_{1}-\varpi_{2}}+e^{\varpi_{2}}+e^{-\varpi_{1}+3\varpi_{2}})MC_{-q}(X(s_{1})^{\circ})\\ &-(q-1)(e^{-2\varpi_{1}+2\varpi_{2}}+e^{-\varpi_{1}+3\varpi_{2}})MC_{-q}(X(s_{2})^{\circ})\\ &+(q-1)^{2}(e^{-\varpi_{1}}+e^{\varpi_{1}-\varpi_{2}}+e^{-2\varpi_{1}+2\varpi_{2}}+e^{\varpi_{2}}+e^{-\varpi_{1}+3\varpi_{2}})MC_{-q}(X(id)^{\circ})\/.\end{split}$
We now investigate the special multiplication by ±$\mathcal{L}_{\pm\varpi_{i}}$ in the case of minuscule fundamental weights. In this case the coefficients have a particularly pleasing factorization.
Lemma 5.9.
If $\lambda=\varpi_{i}$ is a minuscule weight, then for any $u,w\in W$,
$C_{u,\lambda}^{w}=e^{u(\lambda)}P_{u,\lambda}^{w}(y)\quad\textrm{ and }\quad C_{u,-\lambda}^{w}=(-1)^{\ell(w)-\ell(u)}e^{-w(\lambda)}P_{w_{0}w,\lambda}^{w_{0}u}(y)\/,$
where $P_{u,\lambda}^{w}(y)\in{\mathbb{Z}}[y]$. Furthermore, these polynomials satisfy
$P_{u,\lambda}^{w}(y^{-1})\cdot y^{\ell(w)-\ell(u)}=P_{u,\lambda}^{w}(y)\/,$
i.e., they are palindromic.
Proof.
Since $\lambda=\varpi_{i}$ is minuscule, for any reduced $\lambda$-chain $(\beta_{1},\beta_{2},\ldots,\beta_{l})$, the separating hyperplanes must be of the form $h_{j}:=H_{-\beta_{j},0}$, thus $\hat{r}_{h_{j}}=r_{h_{j_{1}}}$. Therefore the first equality follows from Equation 32, since $-w\hat{r}_{J_{<}}(-\lambda)=-wr_{J_{<}}(-\lambda)=u(\lambda)$. The second equality follows this and from the ‘Star duality’ in 6.2 below. The palindromic property follows from 6.5(a) below. ∎
Remark 5.10.
[…]
$MC(X(w)^{\circ})\cdot\mathcal{L}_{\lambda}=\sum C_{u,\lambda}^{w}(y)MC(X(u)^{\circ})\/$
are polynomials with non-negative coefficients. Since then, we found counterexamples to this conjecture in Lie types $D_{6},E_{6}$ and $A_{7}$.
The next example shows that even in the cases when a coefficient $C_{u,\lambda}^{w}(y)$ happens to have positive coefficients, cancellations may still occur in the formula (32) which calculates it.
Example 5.11.
Consider type $A_{3}$, $\lambda=\varpi_{2}$, $w=s_{1}s_{2}s_{3}s_{1}s_{2}s_{1}$ and $u=s_{3}s_{1}$. An $\varpi_{2}$-chain is $\beta_{1}=\alpha_{2},\beta_{2}=\alpha_{2}+\alpha_{3},\beta_{3}=\alpha_{1}+\alpha_{2},\beta_{4}=\alpha_{1}+\alpha_{2}+\alpha_{3}$; see Example 3.6. We have two paths from $u$ to $w$ in (32):
  • $J_{1}=\{2,3\}$ which gives $u<us_{\beta_{3}}<us_{\beta_{3}}s_{\beta_{2}}=w$, where $\ell(us_{\beta_{3}})+3=\ell(w)$;
  • $J_{2}=\{1,2,3,4\}$ which gives $u<us_{\beta_{4}}<us_{\beta_{4}}s_{\beta_{3}}<us_{\beta_{4}}s_{\beta_{3}}s_{\beta_{2}}<us_{\beta_{4}}s_{\beta_{3}}s_{\beta_{2}}s_{\beta_{1}}=w$.
The path $J_{1}$ gives coefficient $(-1-y)\times(-1-y)(-y)e^{u(\lambda)}=-y(y+1)^{2}e^{u(\lambda)}$, and the path $J_{2}$ gives coefficient $(-1-y)^{4}e^{u(\lambda)}=(y+1)^{4}e^{u(\lambda)}$. Therefore
$C_{u,\lambda}^{w}=e^{u(\lambda)}\left((y+1)^{4}-y(y+1)^{2}\right)=e^{u(\lambda)}(y^{2}+y+1)(y+1)^{2}\/.$
5.4. Operator formula
In this section, we reformulate the $\lambda$-chain Chevalley formula from Theorem 5.5 via operators generalizing to motivic Chern classes similar ones from [LP07].
Let $h:=(\rho,\theta^{\vee})+1$ be the Coxeter number, where $\rho:=\sum_{i=1}^{r}\varpi_{i}$ and $\theta^{\vee}$ is the highest coroot. Let $\tilde{R}(T):={\mathbb{Z}}[e^{\pm\varpi_{1}/h},\ldots,e^{\pm\varpi_{r}/h}]$, and let
$\tilde{K}_{T}(G/B):=K_{T}(G/B)[y]\otimes_{K_{T}(\operatorname{pt})[y]}\operatorname{Frac}(\tilde{R}(T)[y]).$
Then $\tilde{K}_{T}(G/B)$ has a basis over $\operatorname{Frac}(\tilde{R}(T)[y])$ given by the motivic Chern classes of the Schubert cells. Define $\operatorname{Frac}(\tilde{R}(T)[y])$-linear operators $B_{\beta}$ $(\beta\in R^{+})$ and $E^{\mu}$ ($\mu\in X^{*}(T)$) on $\tilde{K}_{T}(G/B)$ by
$\begin{array}[]{lll}B_{\beta}(MC_{y}(X(w)^{\circ})):=\begin{cases}(-1-y)(-y)^{\frac{\ell(w)-\ell(ws_{\beta})-1}{2}}MC_{y}(X(ws_{\beta})^{\circ})&\text{ if }ws_{\beta}<w\\[5.69046pt] 0&\text{ otherwise}\end{cases},\\ \\ E^{\mu}(MC_{y}(X(w)^{\circ})):=e^{\frac{w(\mu)}{h}}MC_{y}(X(w)^{\circ}).\end{array}$
If $\beta\in R^{-}$, define $B_{\beta}:=-B_{-\beta}$. Then
$E^{\mu}E^{\mu^{\prime}}=E^{\mu+\mu^{\prime}},\textit{\quad and \quad}B_{\beta}E^{s_{\beta}\mu}=E^{\mu}B_{\beta}.$
Given a $\lambda$-chain $(\beta_{1},\ldots,\beta_{l})$, define
$R^{[\lambda]}:=R_{\beta_{l}}R_{\beta_{l-1}}\cdots R_{\beta_{1}},\;\text{ where}\;R_{\beta}:=E^{\rho}(E^{\beta}+B_{\beta})E^{-\rho}=E^{\beta}+E^{(\rho,\beta^{\vee})\beta}B_{\beta}\/.$
Then we have the following operator formula.
Theorem 5.12 (Operator Chevalley formula).
For any integral weight $\lambda\in X^{*}(T)$,
(33) ${\mathcal{L}}_{\lambda}\otimes MC_{y}(X(w)^{\circ})=R^{[\lambda]}\left(MC_{y}(X(w)^{\circ})\right).$
Remark 5.13.
  1. (1)
    The theorem implies that the definition of $R^{[\lambda]}$ does not depend on the choice of the $\lambda$-chain, which is equivalent to the Yang–Baxter equations ([LP07, Definition 9.1]) satisfied by the operators $R_{\beta}$.
  2. (2)
    The formula analogous to (33) involving SMC classes is obtained by replacing the operator $R^{|\lambda|}$ by an operator defined via the adjoint operators of $B_{\beta}$ and $E^{\mu}$.
  3. (3)
    Recall from Remark 4.7(1) that $MC_{0}(X(w)^{\circ})=[{\mathcal{O}}_{X(w)}(-\partial X(w))]$. Specializing $y=0$ in the Theorem, we get a dual version of Lenart–Postnikov’s formula [LP07, Theorem 13.1].
For the proof, we need to recall the following result from [LP07]. Recall $(\beta_{1},\ldots,\beta_{l})$ is a $\lambda$-chain. The hyperplane $h_{j}:=H_{-\beta_{j},d_{j}}$ separates the alcoves $A_{j-1}$ and $A_{j}$ in the corresponding alcove path, and $\hat{r}_{h_{j}}$ is the reflection along $h_{j}$.
Lemma 5.14.
[LP07, Proof of Prop. 14.5] For any $1\leq j_{1}<j_{2}<\cdots<j_{t}\leq l$,
$-\rho+\beta_{1}+\cdots+\beta_{j_{1}-1}+s_{\beta_{j_{1}}}(\beta_{j_{1}+1}+\cdots\beta_{j_{2}-1})+\cdots+s_{\beta_{j_{1}}}\cdots s_{\beta_{j_{t}}}(\beta_{j_{t}+1}+\cdots\beta_{l}+\rho)$
$=$ $-h\hat{r}_{j_{1}}\cdots\hat{r}_{j_{t}}(-\lambda).$
Proof of Theorem 5.12.
By definition,
$R^{[\lambda]}=$ $R_{\beta_{l}}R_{\beta_{l-1}}\cdots R_{\beta_{1}}$
$=$ $E^{\rho}(E^{\beta_{l}}+B_{\beta_{l}})\cdots(E^{\beta_{2}}+B_{\beta_{2}})(E^{\beta_{1}}+B_{\beta_{1}})E^{-\rho}$
$=$ $\sum_{J}R^{[\lambda]}_{J}.$
Here $J=\{j_{1}<j_{2}<\cdots<j_{t}\}$ is a subset of $\{1,2,\dots,l\}$, and $R^{[\lambda]}_{J}$ is the term that contains $B_{\beta_{j}}$ if $j\in J$, and $E^{\beta_{j}}$, otherwise. Moving all the $B$-operators to the left, we get
$R^{[\lambda]}_{J}=B_{\beta_{j_{t}}}\cdots B_{\beta_{j_{1}}}E^{-h\hat{r}_{j_{1}}\cdots\hat{r}_{j_{t}}(-\lambda)}=B_{\beta_{j_{t}}}\cdots B_{\beta_{j_{1}}}E^{-h\hat{r}_{J<}(-\lambda)}.$
Here we have used the relation $B_{\beta}E^{s_{\beta}\mu}=E^{\mu}B_{\beta}$ and Lemma 5.14. Then the Theorem follows from this and Theorem 5.5. ∎
5.5. Parabolic case
In this section, we extend the above Chevalley formula to the partial flag variety case $G/P$.
Theorem 5.15 (Chevalley formula for the $G/P$ case).
Let $w\in W^{P}$ and $\lambda\in X^{*}(T)_{P}$. Then we have
${\mathcal{L}}_{\lambda}\otimes MC_{y}(X(wW_{P})^{\circ})=\sum_{u\in W^{P}}\left(\sum_{v\in uW_{P}}(-y)^{\ell(v)-\ell(u)}C^{w}_{v,\lambda}\right)MC_{y}(X(uW_{P})^{\circ}),$
and
${\mathcal{L}}_{\lambda}\otimes SMC_{y}(Y(uW_{P})^{\circ})=\sum_{w\in W^{P}}\left(\sum_{v\in uW_{P}}(-y)^{\ell(v)-\ell(u)}C^{w}_{v,\lambda}\right)SMC_{y}(Y(wW_{P})^{\circ}),$
where $C^{v}_{u,\lambda}$ are the Chevalley coefficients for full flag $G/B$ given explicitly Theorem 5.5.
Proof.
The first equality follows from equations (25) and (27), while the second equality follows from the first one and the duality in 4.8(2). ∎
Remark 5.16.
In a paper in preparation we use this theorem to give a combinatorial Chevalley formula for minuscule flag varieties and a $K$-theoretic generalization of Nakada’s colored hook formula. We also recently learned about work by Neil Fan, Peter Guo, Changjian Su, and Rui Xiong who proved Pieri formulae for the motivic Chern classes on Grassmannians.
6. Dualities of Chevalley coefficients
Recall the expansion from Equation 27:
${\mathcal{L}}_{\lambda}\otimes MC_{y}(X(w)^{\circ})=\sum_{u\leq w}C_{u,\lambda}^{w}MC_{y}(X(u)^{\circ}).$
In this section we state and prove several duality properties of the Chevalley coefficients $C_{u,\lambda}^{w}$. All these dualities have a geometric origin (Serre duality, star duality, Dynkin automorphisms) and we name the Chevalley dualities correspondingly.
Recall the Grothendieck-Serre operator $\mathcal{D}$ from (23), and the duality functor $(-)^{\vee}$ on $K_{T}(G/B)[y^{\pm 1}]$ and $K_{T}(\operatorname{pt})[y^{\pm 1}]$, which sends a vector bundle to its dual and $y^{i}$ to $y^{-i}$. This induces the following property of the Chevalley coefficients in Equation 27.
Proposition 6.1 (Serre duality).
$C^{w}_{u,\lambda}=(-y)^{\ell(w)-\ell(u)}w_{0}(C^{w_{0}u}_{w_{0}w,-\lambda})^{\vee}.$
Proof.
Since
$SMC_{y}(Y(w)^{\circ})=(-y)^{\dim Y(w)}\frac{{\mathcal{D}}(MC_{y}(Y(w)^{\circ}))}{\lambda_{y}(T^{*}(G/B))},$
Equation 28 can be written as
$\mathcal{L}_{\lambda}\otimes{\mathcal{D}}(MC_{y}(Y(u)^{\circ}))=\sum_{w\in W}C^{w}_{u,\lambda}(-y)^{\ell(u)-\ell(w)}{\mathcal{D}}(MC_{y}(Y(w)^{\circ})).$
Taking the dual ${\mathcal{D}}$ on both sides, we have
$\mathcal{L}_{-\lambda}\otimes MC_{y}(Y(u)^{\circ})=\sum_{w\in W}(C^{w}_{u,\lambda})^{\vee}(-y)^{\ell(w)-\ell(u)}MC_{y}(Y(w)^{\circ}).$
Finally, applying the left $w_{0}$ action to the above identity, we get
$\mathcal{L}_{-\lambda}\otimes MC_{y}(X(w_{0}u)^{\circ})=\sum_{w\in W}w_{0}(C^{w}_{u,\lambda})^{\vee}(-y)^{\ell(w)-\ell(u)}MC_{y}(X(w_{0}w)^{\circ}).$
This finishes the proof of the theorem by comparing the above equation with Equation 27. ∎
The star duality *** acts on $K_{T}(G/B)[y^{\pm 1}]$ and $K_{T}(pt)[y^{\pm 1}]$ by sending a vector bundle to its dual, and leaving $y^{i}$ unchanged. Consider the composition
$\iota:=w_{0}*:K_{T}(\operatorname{pt})[y^{\pm 1}]\to K_{T}(\operatorname{pt})[y^{\pm 1}]\/.$
Then we have the following result by combining Theorem 9.1(a) and Remark 4.7 of [AMSS22].
(34) ${\mathbb{C}}_{-\rho}\otimes{\mathcal{L}}_{-\rho}\otimes MC_{y}(X(w)^{\circ})=(-1)^{\dim G/B-\ell(w)}\prod_{\alpha>0}(1+ye^{-\alpha})*\bigg{(}SMC_{y}(X(w)^{\circ})\bigg{)}.$
Proposition 6.2 (Star duality).
$C_{u,\lambda}^{w}=(-1)^{\ell(w)-\ell(u)}\iota(C_{w_{0}w,-\lambda}^{w_{0}u}).$
Proof.
Applying the star duality functor *** to Equation 27, then use Equation 34 to get
${\mathcal{L}}_{-\lambda}\otimes SMC_{y}(X(w)^{\circ})=\sum_{u}(-1)^{\ell(w)-\ell(u)}*(C_{u,\lambda}^{w})SMC_{y}(X(u)^{\circ}).$
Applying the left $w_{0}$ action to the above equation and comparing with Equation 28, we get the result. ∎
The map sending $\alpha_{i}\mapsto-w_{0}(\alpha_{i})$ for every simple root $\alpha_{i}$ induces an automorphism on the Dynkin diagram, hence also on the flag variety $G/B$. This automorphism maps $X(w)^{\circ}$ to $X(w_{0}ww_{0})^{\circ}$, and it induces a ring automorphism $\phi$ on $K_{T}(G/B)$, which sends ${\mathcal{L}}_{\lambda}$ to ${\mathcal{L}}_{-w_{0}\lambda}$, and twists the base ring $K_{T}(\operatorname{pt})$ by the map $\iota$ above.
Proposition 6.3 (Dynkin duality).
$C_{u,\lambda}^{w}=\iota(C_{w_{0}uw_{0},-w_{0}\lambda}^{w_{0}ww_{0}}).$
Proof.
Applying the Dynkin automorphism $\phi$ to Equation 27, we obtain
$\mathcal{L}_{-w_{0}\lambda}\otimes MC_{y}(X(w_{0}ww_{0})^{\circ})=\sum_{u\in W}\iota(C^{w}_{u,\lambda})MC_{y}(X(w_{0}uw_{0})^{\circ}).$
Then the claim follows from the definition of the coefficients $C_{u,\lambda}^{w}$. ∎
Combining the above dualities we obtain the following.
Proposition 6.4.
  1. (1)
    (Serre duality + star duality)
    $(C^{w}_{u,\lambda})|_{y\to y^{-1}}\times y^{\ell(w)-\ell(u)}=C^{w}_{u,\lambda}.$
  2. (2)
    (Star duality + Dynkin duality)
    $C^{w}_{u,\lambda}=(-1)^{\ell(w)-\ell(u)}C^{uw_{0}}_{ww_{0},w_{0}\lambda}.$
Proof.
The first part follows directly from 6.1 and 6.2, and the second part follows from 6.2 and 6.3. This finishes the proof. ∎
Using the $\lambda$-chain formula in Theorem 5.5, we can also give a direct combinatorial proof for the various duality identities in this section. The proofs are similar to [LP07, Theorem 8.6] and [LP07, Theorem 8.7].
Proposition 6.5.
For any integral weight $\lambda$ and $w,u\in W$,
  • (a)
    $C^{w}_{u,\lambda}\in K_{T}(pt)[y]$ and $(C^{w}_{u,\lambda})|_{y\to y^{-1}}\times y^{\ell(w)-\ell(u)}=C^{w}_{u,\lambda}$ i.e., $C^{w}_{u,\lambda}$ is palindromic regarded as a polynomial in $y$; (cf. 6.4 (1));
  • (b)
    $C^{w}_{u,\lambda}=(-1)^{\ell(w)-\ell(u)}C^{uw_{0}}_{ww_{0},w_{0}\lambda};$  (cf. 6.4 (2));
  • (c)
    $C^{w}_{u,\lambda}=(-1)^{\ell(w)-\ell(u)}\iota\left(C^{{w_{o}}u}_{{w_{o}}w,-\lambda}\right);$  (cf. 6.2).
  • (d)
    $C^{w}_{u,\lambda}=\iota\left(C^{w_{0}ww_{0}}_{w_{0}uw_{0},-w_{0}\lambda}\right)$;  (cf. 6.3 ).
Proof.
Part (a) is straightforward from Equation 32; part (d) follows from (b) and (c). Parts (b) and (c) follow from known properties of $\lambda$-chain, i.e., if $(\beta_{1},\ldots,\beta_{l})$ is a $\lambda$-chain, then $(w_{0}\beta_{l},\ldots,w_{0}\beta_{1})$ is a $(w_{0}\lambda)$-chain and $(-\beta_{l},\ldots,-\beta_{1})$ is a $(-\lambda)$-chain. ∎
Remark 6.6.
The Serre duality extends to any $G/P$. For $w,u\in W^{P}$and $\lambda\in X^{*}(T)_{P}$ recall the expansion:
${\mathcal{L}}_{\lambda}\otimes MC_{y}(X(wW_{P})^{\circ})=\sum_{u\in W^{P}}C^{w,P}_{u,\lambda}MC_{y}(X(uW_{P})^{\circ}).$
Then Theorem 5.15 can be rewritten as $C^{w,P}_{u,\lambda}=\displaystyle\sum_{v\in uW_{P}}(-y)^{\ell(v)-\ell(u)}C^{w}_{v,\lambda}$. By the same argument as in the proof of 6.1, we have the Serre duality on parabolic Chevalley coefficients;
(35) $C^{w,P}_{u,\lambda}=(-y)^{\ell(w)-\ell(u)}w_{0}(C^{\overline{w_{0}u},P}_{\overline{w_{0}w},-\lambda})^{\vee},$
where $\overline{w_{0}w}$ and $\overline{w_{0}u}\in W^{P}$ are minimal coset representatives of $w_{0}wW_{P}$ and $w_{0}uW_{P}$.
7. K-theoretic stable envelopes for $T^{*}(G/B)$
In this section we apply the Chevalley formula for motivic classes to calculate the transformation of stable envelopes in $T^{*}(G/B)$ under the change of arbitrary alcoves. For alcoves adjacent to the fundamental alcove, a formula for this transformation was obtained in [SZZ20, Theorem 5.4], see also [KW23].
7.1. Definition of the stable envelopes
The stable envelopes were defined by Maulik and Okounkov in their seminal work on quantum cohomology of Nakajima quiver varieties [MO19]. Later, this was generalized by Okounkov and his collaborators to $K$-theory and elliptic cohomology [Oko17, AO21]. We recall next the definition of the stable envelopes for $T^{*}(G/B)$.
The torus $T$ acts by left multiplication on $G/B$. Hence, it induces a natural action on $T^{*}(G/B)$. There is also a natural dilation ${\mathbb{C}}^{*}$ action on the cotangent fibers by a character of $q^{-1}$. Throughout this section, we use $q^{1/2}$ to denote the standard representation of ${\mathbb{C}}^{*}$, so that $K_{T\times{\mathbb{C}}^{*}}(\operatorname{pt})=K_{T}(\operatorname{pt})[q^{\pm 1/2}]$.
The definition of the stable envelopes depends on three parameters:
  • a chamber ${\mathcal{C}}$ in the Lie algebra of the maximal torus $T$.
  • a polarization $T^{\frac{1}{2}}\in K_{T\times{\mathbb{C}}^{*}}(T^{*}(G/B))$ of the tangent bundle $T(T^{*}(G/B))$, i.e., a solution of the equation
    $T^{\frac{1}{2}}+q^{-1}(T^{\frac{1}{2}})^{\vee}=T(T^{*}(G/B))$
    in the ring $K_{T\times{\mathbb{C}}^{*}}(T^{*}(G/B))_{\operatorname{loc}}$.
  • an alcove $A$ in $\mathfrak{t}^{*}_{\mathbb{R}}$, which is also called a slope in [Oko17].
Given a polarization $T^{\frac{1}{2}}$, there is an opposite polarization defined by $T^{\frac{1}{2}}_{\operatorname{opp}}:=q^{-1}(T^{\frac{1}{2}})^{\vee}$. A typical example of a polarization is $T(G/B)$ (pulled back from $G/B$ to $T^{*}(G/B)$). Its opposite is equal to $T^{*}(G/B)$. Because we will work with fibers of polarizations over fixed points, in what follows we will assume that the polarization is given by a subbundle of $T(T^{*}(G/B))$, or possibly a virtual vector subbundle, i.e., a formal difference of such subbundles.
The torus fixed point set $(T^{*}(G/B))^{T}=(G/B)^{T}$ is in one-to-one correspondence with the Weyl group $W$. For every $w\in W$, recall that $e_{w}$ denotes the corresponding fixed point. For a chosen Weyl chamber $\mathfrak{C}$ in $\operatorname{Lie}T$, pick any cocharacter $\sigma\in\mathfrak{C}$. The attracting set of the fixed point $e_{w}$, also called the Białynicki-Birula cell in the literature, is defined by
$\operatorname{Attr}_{\mathfrak{C}}(w)=\left\{x\in T^{*}(G/B)\mid\lim\limits_{z\rightarrow 0}\sigma(z)\cdot x=e_{w}\right\}.$
By analyzing the (signs of the roots in the) weight space decomposition of $T_{w}(T^{*}(G/B))$, one may show that $\operatorname{Attr}_{\mathfrak{C}}(w)$ is the conormal bundle of the attracting variety of $w$ in $G/B$, i.e., the conormal bundle of the Schubert cell stable under the Borel subgroup associated to the chamber $\mathfrak{C}$ [^4] We note that $T^{*}(G/B)$ is not compact, so not all points have well defined limits at 00. For example, if $\mathfrak{C}$ is the positive chamber, then the points in the open set $T^{*}(X(w_{0})^{\circ})\setminus X(w_{0})^{\circ}\subset T^{*}(G/B)$ do not have limits at 00. Define a partial order on the fixed point set $W$ to be the (transitive closure of the) following relation:
$e_{w}\preceq_{\mathfrak{C}}e_{v}\text{\quad if \quad}\overline{\operatorname{Attr}_{\mathfrak{C}}(v)}\cap e_{w}\neq\emptyset.$
Then the order determined by the positive (resp., negative) chamber is the same as the Bruhat order (resp., the opposite Bruhat order).
Any chamber $\mathfrak{C}$ determines a decomposition of the tangent space $N_{w}:=T_{w}(T^{*}(G/B))$ as $N_{w}=N_{w,+}\oplus N_{w,-}$ into $T$-weight spaces which are positive and negative with respect to $\mathfrak{C}$ respectively. For every polarization $T^{\frac{1}{2}}$, denote $N_{w}\cap T^{1/2}|_{w}$ by $N_{w}^{\frac{1}{2}}$. Similarly, we have $N_{w,+}^{\frac{1}{2}}$ and $N_{w,-}^{\frac{1}{2}}$. In particular, $N_{w,-}=N^{\frac{1}{2}}_{w,-}\oplus q^{-1}(N_{w,+}^{\frac{1}{2}})^{\vee}$. Consequently, we have
$N_{w,-}-N_{w}^{\frac{1}{2}}=q^{-1}(N_{w,+}^{\frac{1}{2}})^{\vee}-N_{w,+}^{\frac{1}{2}}$
as virtual vector bundles. The determinant bundle of the virtual bundle $N_{w,-}-N_{w}^{\frac{1}{2}}$ is a complete square and its square root will be denoted by $\left(\frac{\det N_{w,-}}{\det N_{w}^{\frac{1}{2}}}\right)^{\frac{1}{2}}$; cf. [Oko17, §9.1.5]. For instance, if we choose the polarization $T^{1/2}=T(G/B)$, the positive chamber, and $w=\operatorname{id}$ then both $N_{\operatorname{id}}^{\frac{1}{2}}$ and $N_{\operatorname{id},-}$ have weights $-\alpha$, where $\alpha$ varies in the set of positive roots; in this case the virtual bundle $N_{\operatorname{id},-}-N_{\operatorname{id}}^{\frac{1}{2}}=0$.
Let $f:=\sum_{\mu}f_{\mu}e^{\mu}\in K_{T\times{\mathbb{C}}^{*}}(\operatorname{pt})$ be a Laurent polynomial, where $e^{\mu}\in K_{T}(\operatorname{pt})$ and $f_{\mu}\in{\mathbb{Q}}[q^{1/2},q^{-1/2}]$. The Newton polytope of $f$, denoted by $\deg_{T}f$, is
$\deg_{T}f=\mbox{Convex hull }(\{\mu|f_{\mu}\neq 0\})\subseteq X^{*}(T)\otimes_{\mathbb{Z}}{\mathbb{Q}},$
where $X^{*}(T)$ denotes the character lattice of $T$.  The following theorem defines the $K$-theoretic stable envelopes.
Theorem 7.1.
[Oko17, §9.1], [OS22, Thm. 1]. For every chamber $\mathfrak{C}$, a polarization $T^{1/2}$, and an alcove $A$, there exists a unique map of $K_{T\times{\mathbb{C}}^{*}}(\operatorname{pt})$-modules
$\operatorname{stab}_{\mathfrak{C},T^{\frac{1}{2}},A}:K_{T\times{\mathbb{C}}^{*}}((T^{*}(G/B))^{T})\rightarrow K_{T\times{\mathbb{C}}^{*}}(T^{*}(G/B))$
such that for every $w\in W$, the class $\Gamma:=\operatorname{stab}_{\mathfrak{C},T^{\frac{1}{2}},A}(w)$ satisfies:
  1. (1)
    (Support) $\operatorname{Supp}\Gamma\subseteq\cup_{z\preceq_{\mathfrak{C}}w}\overline{\operatorname{Attr}_{\mathfrak{C}}(z)}$;
  2. (2)
    (Normalization) $\Gamma|_{w}=(-1)^{\operatorname{rk}N_{w,+}^{\frac{1}{2}}}\left(\frac{\det N_{w,-}}{\det N_{w}^{\frac{1}{2}}}\right)^{\frac{1}{2}}{\mathcal{O}}_{\operatorname{Attr}_{\mathfrak{C}}(w)}|_{w}$;
  3. (3)
    (Degree) For every $e_{v}\prec_{\mathfrak{C}}e_{w}$,
    $\deg_{T}\Gamma|_{v}\subseteq\deg_{T}\operatorname{stab}_{\mathfrak{C},T^{\frac{1}{2}},A}(v)|_{v}+v\lambda-w\lambda\/,$
    where $\lambda\in(X^{*}(T)\otimes_{{\mathbb{Z}}}{\mathbb{Q}})\cap A$ is any rational weight in the alcove $A$.
Strictly speaking, $\operatorname{stab}_{\mathfrak{C},T^{\frac{1}{2}},A}(w)\in K_{T\times{\mathbb{C}}^{*}}(G/B)$ denotes the image of $1\in K_{T\times{\mathbb{C}}^{*}}(e_{w})$ under the map $\operatorname{stab}_{\mathfrak{C},T^{\frac{1}{2}},A}$. From the definition, it is immediate to see that $\{\operatorname{stab}_{\mathfrak{C},T^{\frac{1}{2}},A}(w)\mid w\in W\}$ forms a basis for the localized equivariant K-theory $K_{T\times{\mathbb{C}}^{*}}(T^{*}(G/B))_{\operatorname{loc}}$, called the stable basis. Explicit combinatorial formulae and recursions for the localizations of stable envelopes in $T^{*}(G/B)$ may be found in [AMSS19, §8.3].
7.2. Changing the polarizations
A natural question is to study the change of the stable envelopes when we vary the above three parameters. To start, the change of chambers is encoded in the left Weyl group action. More precisely, the group $G$ acts on $T^{*}(G/B)$ by left multiplication, which induces a left Weyl group action on $K_{T\times{\mathbb{C}}^{*}}(T^{*}(G/B))$, see [MNS22b]. If we change the chamber $\mathfrak{C}$ to another chamber $w(\mathfrak{C})$ ($w\in W$), we have the following formula (see [AMSS19, Lemma 8.2(a)])
$w\cdot(\operatorname{stab}_{{\mathcal{C}},T^{\frac{1}{2}},A}(u))=\operatorname{stab}_{w({\mathcal{C}}),w(T^{\frac{1}{2}}),A}(wu).$
Next we consider the change of polarizations. In this case the results are stated e.g. in [Oko17] in the more general setting of symplectic resolutions; for the convenience of the reader we include proofs for $T^{*}(G/B)$. We have the following lemma relating two arbitrary polarizations see [Oko17, Section 7.5.8].
Lemma 7.2.
For any two polarizations $T_{1}^{\frac{1}{2}}$ and $T_{2}^{\frac{1}{2}}$, there exists a class ${\mathcal{F}}\in K_{T\times{\mathbb{C}}^{*}}(T^{*}(G/B))$ such that $T_{1}^{\frac{1}{2}}-T_{2}^{\frac{1}{2}}={\mathcal{F}}-q^{-1}{\mathcal{F}}^{\vee}$.
Remark 7.3.
Classes of the form ${\mathcal{F}}-q^{-1}{\mathcal{F}}^{\vee}$ are called balanced classes in loc. cit.
Proof.
It suffices to prove the following fact: any solution of the equation ${\mathcal{G}}+q^{-1}{\mathcal{G}}^{\vee}=0\in K_{T\times{\mathbb{C}}^{*}}(T^{*}(G/B))$ must be of the form ${\mathcal{G}}={\mathcal{F}}-q^{-1}{\mathcal{F}}^{\vee}$ for some ${\mathcal{F}}\in K_{T\times{\mathbb{C}}^{*}}(T^{*}(G/B))$. Since ${\mathcal{G}}\in K_{T\times{\mathbb{C}}^{*}}(T^{*}(G/B))\simeq K_{T}(G/B)[q,q^{-1}]$, we can write ${\mathcal{G}}={\mathcal{F}}+q^{-1}{\mathcal{F}}^{\prime}$ for some ${\mathcal{F}}\in K_{T}(G/B)[q]$ and ${\mathcal{F}}^{\prime}\in q^{-1}K_{T}(G/B)[q^{-1}]$. Thus,
${\mathcal{F}}+q^{-1}{\mathcal{F}}^{\prime}+q^{-1}{\mathcal{F}}^{\vee}+({\mathcal{F}}^{\prime})^{\vee}=0.$
Since ${\mathcal{F}},({\mathcal{F}}^{\prime})^{\vee}\in K_{T}(G/B)[q]$ and $q^{-1}{\mathcal{F}}^{\prime},q^{-1}{\mathcal{F}}^{\vee}\in q^{-1}K_{T}(G/B)[q^{-1}]$, we get
${\mathcal{F}}+({\mathcal{F}}^{\prime})^{\vee}=0,\textit{\quad and \quad}q^{-1}{\mathcal{F}}^{\prime}+q^{-1}{\mathcal{F}}^{\vee}=0.$
Therefore, ${\mathcal{F}}^{\prime}=-{\mathcal{F}}^{\vee}$, and ${\mathcal{G}}={\mathcal{F}}-q^{-1}{\mathcal{F}}^{\vee}$. ∎
For any two polarizations $T_{1}^{\frac{1}{2}}$ and $T_{2}^{\frac{1}{2}}$ let $\mathcal{F}$ be defined by $T_{1}^{\frac{1}{2}}-T_{2}^{\frac{1}{2}}={\mathcal{F}}-q^{-1}{\mathcal{F}}^{\vee}$ as in the above lemma. Define a $T\times{\mathbb{C}}^{*}$-equivariant line bundle ${\mathcal{L}}$ on $T^{*}(G/B)$ by the following formula
${\mathcal{L}}:=\det{\mathcal{F}}.$
Here for a virtual bundle ${\mathcal{F}}={\mathcal{V}}_{1}-{\mathcal{V}}_{2}$, $\det{\mathcal{F}}:=\det{\mathcal{V}}_{1}/\det{\mathcal{V}}_{2}$.
Example 7.4.
Consider the polarization $T_{1}^{\frac{1}{2}}:=T(G/B)$ and the opposite polarization $T_{2}^{\frac{1}{2}}=q^{-1}(T_{1}^{\frac{1}{2}})^{\vee}=T^{*}(G/B)$. Then the element ${\mathcal{F}}$ in Lemma 7.2 can be taken to be $T(G/B)$, therefore ${\mathcal{L}}=\det{\mathcal{F}}={\mathcal{L}}_{-2\rho}$.
The next proposition shows that the change of polarizations results in a multiplication by a line bundle, therefore it is encoded in the Chevalley formula for the stable envelopes.
Proposition 7.5.
[Oko17, Exercise 9.1.12] The following holds:
$\operatorname{stab}_{\mathfrak{C},T_{2}^{\frac{1}{2}},A}(w)=(-1)^{\operatorname{rk}N_{w,+,2}^{\frac{1}{2}}-\operatorname{rk}N_{w,+,1}^{\frac{1}{2}}}q^{\frac{\operatorname{rk}{\mathcal{F}}|_{w}}{2}}{\mathcal{L}}\otimes\operatorname{stab}_{\mathfrak{C},T_{1}^{\frac{1}{2}},A}(w),$
where $N_{w,+,i}^{\frac{1}{2}}:=N_{w,+,i}\cap T^{\frac{1}{2}}_{i}$ for $i=1,2$, and if ${\mathcal{F}}=T_{1}^{1/2}-T_{2}^{1/2}={\mathcal{V}}_{1}-{\mathcal{V}}_{2}$ is the virtual bundle from Lemma 7.2, then $\operatorname{rk}{\mathcal{F}}|_{w}:=\operatorname{rk}{\mathcal{V}}_{1}|_{w}-\operatorname{rk}{\mathcal{V}}_{2}|_{w}.$
Proof.
From the characterization Theorem 7.1, it suffices to show that the right hand side satisfies the defining properties of the stable envelope on the left hand side. The support condition is immediate. For the degree condition, we need to check that for every $e_{v}\prec_{\mathfrak{C}}e_{w}$,
$\deg_{T}({\mathcal{L}}\otimes\operatorname{stab}_{\mathfrak{C},T_{1}^{\frac{1}{2}},A}(w))|_{v}\subseteq\deg_{T}({\mathcal{L}}\otimes\operatorname{stab}_{\mathfrak{C},T_{1}^{\frac{1}{2}},A}(v))|_{v}+v\lambda-w\lambda\/,$
for some $\lambda\in A$. The terms ${\mathcal{L}}|_{v}$ on both sides cancel, and the above inclusion reduces to the degree condition for $\operatorname{stab}_{\mathfrak{C},T_{1}^{\frac{1}{2}},A}(w)$. The normalization condition follows from the next calculation. We utilize our assumption that ${\mathcal{F}}$ is represented by a subbundle of $T(T^{*}(G/B))$, and we note also that the weights of $N_{w}=T_{w}(T^{*}(G/B))$ are distinct.
$(-1)^{\operatorname{rk}N_{w,+,2}^{\frac{1}{2}}-\operatorname{rk}N_{w,+,1}^{\frac{1}{2}}}\frac{\operatorname{stab}_{\mathfrak{C},T_{2}^{\frac{1}{2}},A}(w)|_{w}}{\operatorname{stab}_{\mathfrak{C},T_{1}^{\frac{1}{2}},A}(w)|_{w}}$
$=$ $\left(\frac{\det N_{w,-}}{\det N_{w,2}^{\frac{1}{2}}}\right)^{\frac{1}{2}}\left(\frac{\det N_{w,-}}{\det N_{w,1}^{\frac{1}{2}}}\right)^{-\frac{1}{2}}$
$=$ $\left(\frac{\det N_{w,1}^{\frac{1}{2}}}{\det N_{w,2}^{\frac{1}{2}}}\right)^{\frac{1}{2}}$
$=$ $\left(\det N_{w}\cap({\mathcal{F}}|_{w}-q^{-1}{\mathcal{F}}^{\vee}|_{w})\right)^{\frac{1}{2}}$
$=$ $\left(\frac{\det(N_{w}\cap{\mathcal{F}}|_{w})}{\det(N_{w}\cap q^{-1}({\mathcal{F}}|_{w})^{\vee})}\right)^{\frac{1}{2}}$
$=$ $q^{\frac{\operatorname{rk}{\mathcal{F}}|_{w}}{2}}{\mathcal{L}}|_{w}.$
The reason for the last equality is as follows. We have that $N_{w}\cap{\mathcal{F}}|_{w}={\mathcal{F}}|_{w}$, and suppose $e^{\lambda}$ is a torus weight of it. Since $N_{w}$ is a symplectic vector space, $q^{-1}e^{-\lambda}$ is a weight of $N_{w}$. Hence $q^{-1}e^{-\lambda}$ is also a weight of the intersection $N_{w}\cap q^{-1}({\mathcal{F}}|_{w})^{\vee}$ and in fact $N_{w}\cap q^{-1}({\mathcal{F}}|_{w})^{\vee}=q^{-1}({\mathcal{F}}|_{w})^{\vee}$, giving the equality. The case of ${\mathcal{F}}={\mathcal{V}}_{1}-{\mathcal{V}}_{2}$ being a virtual bundle follows from linearity of the constructions. ∎
7.3. Changing the alcoves
We now turn to what happens under the change of alcoves. This can be answered utilizing the recursive formulae from [SZZ20, SZZ21], see Theorem 7.6 below. Our Chevalley formula provides a non-recursive answer, in terms of $\lambda$-chains, relating the stable envelope for the fundamental alcove $\operatorname{A_{\circ}}$ to the stable envelope for a translate $\operatorname{A_{\circ}}+\lambda$; see Theorem 7.8 below. We recall some of the formulae below, and state our Chevalley based formula in Theorem 7.8.
The alcoves in $\mathfrak{t}^{*}_{\mathbb{R}}$ are of the form $x(\operatorname{A_{\circ}})+\lambda$ for some $x\in W$ and some $\lambda$ in the root lattice. It was proved in [AMSS19, Lemma 8.2],[SZZ20, Rmk. 2.3] that
(36) $\operatorname{stab}_{\mathfrak{C},T^{\frac{1}{2}},A+\lambda}(w)=e^{-w\lambda}{\mathcal{L}}_{\lambda}\otimes\operatorname{stab}_{\mathfrak{C},T^{\frac{1}{2}},A}(w)\/,$
where ${\mathcal{L}}_{\lambda}$ is the pullback of $G\times^{B}{\mathbb{C}}_{\lambda}$ from $G/B$ to $T^{*}(G/B)$. Fix the chamber ${\mathcal{C}}$ to be the anti-dominant Weyl chamber
${\mathcal{C}}:=\{\lambda\in\mathfrak{t}^{*}_{\mathbb{R}}\mid\langle\lambda,\alpha^{\vee}\rangle<0,\textit{ for any positive root }\alpha\}$
and the polarization $T^{\frac{1}{2}}=T^{*}(G/B)$. To simplify the notations, we denote $\operatorname{stab}_{{\mathcal{C}},T^{*}(G/B),A}(w)$ by $\operatorname{stab}_{A}(w)$.
Let $Z=T^{*}(G/B)\times_{\mathcal{N}}T^{*}(G/B)$ be the Steinberg variety where ${\mathcal{N}}\subset\mathfrak{g}$ denotes the nilpotent cone. There is an algebra isomorphism due to Kazhdan–Lusztig and Ginzburg [KL87, CG09]
${\mathbb{H}}\simeq K_{G\times{\mathbb{C}}^{*}}(Z),$
where $K_{G\times{\mathbb{C}}^{*}}(Z)$ has an algebra structure given by convolution. The convolution induces an action of the Hecke algebra ${\mathbb{H}}$ on $K_{T\times{\mathbb{C}}^{*}}(T^{*}(G/B))$ which we recall next. For a simple root $\alpha_{i}$, and the corresponding minimal parabolic subgroup $P_{i}\supset B$, define the operator $T_{i}$ on $K_{T\times{\mathbb{C}}^{*}}(T^{*}(G/B))$ by the following formula:
$T_{i}({\mathcal{F}}):=-{\mathcal{F}}-\pi_{1*}(\pi_{2}^{*}{\mathcal{F}}\otimes\pi_{2}^{*}{\mathcal{L}}_{\alpha_{i}}).$
Here ${\mathcal{F}}\in K_{A\times{\mathbb{C}}^{*}}(T^{*}(G/B))$, $Y_{i}:=G/B\times_{G/P_{i}}G/B\subset G/B\times G/B$, $T^{*}_{Y_{i}}$ is the conormal bundle of $Y_{i}$ inside $G/B\times G/B$, and $\pi_{j}:T^{*}_{Y_{i}}\to T^{*}(G/B)$ $(j=1,2)$ are the two projections. These operators satisfy the quadratic relations and the braid relations in ${\mathbb{H}}$. In particular, $T_{w}$ is well defined for any $w\in W$. Recall that $\operatorname{A_{\circ}}$ denotes the fundamental alcove.
The following have been proved in [SZZ20, SZZ21].
Theorem 7.6.
(a) [SZZ20, Theorem 4.5]
$T_{i}(\operatorname{stab}_{\operatorname{A_{\circ}}}(w))=\begin{cases}(q-1)\operatorname{stab}_{\operatorname{A_{\circ}}}(w)+q^{1/2}\operatorname{stab}_{\operatorname{A_{\circ}}}(ws_{i}),&\textit{\quad if \quad}ws_{i}<w,\\ q^{1/2}\operatorname{stab}_{\operatorname{A_{\circ}}}(ws_{i}),&\textit{\quad if\quad}ws_{i}>w.\end{cases}$
(b) [SZZ21, Theorem 5.4] Let $x\in W$. Then
$\operatorname{stab}_{x(\operatorname{A_{\circ}})}(w)=q^{-\ell(x)/2}T_{x}(\operatorname{stab}_{\operatorname{A_{\circ}}}(wx))\/.$
(c) [SZZ21, Lemma 3.5 and Corollary 5.3] Assume $A_{1}$ and $A_{2}$ are two adjacent alcoves separated by a wall of the form $H_{\alpha,n}$, where $\alpha>0$. Assume $A_{2}$ is on the positive side of $H_{\alpha,n}$, i.e., for any $\mu\in A_{2}$, $(\mu,\alpha^{\vee})>n$. Then
$\operatorname{stab}_{A_{1}}(w)=\begin{cases}\operatorname{stab}_{A_{2}}(w)+e^{-nw\alpha}(q^{1/2}-q^{-1/2})\operatorname{stab}_{A_{2}}(ws_{\alpha})&\textit{\quad if \quad}ws_{\alpha}>w,\\ \operatorname{stab}_{A_{2}}(w)&\textit{\quad if \quad}ws_{\alpha}<w.\end{cases}$
Part (a) of the theorem implies that $T_{x}(\operatorname{stab}_{\operatorname{A_{\circ}}}(wx))$ may be recursively calculated in terms of $\{\operatorname{stab}_{\operatorname{A_{\circ}}}(w)\mid w\in W\}$. Part (b) implies that the same is true for $\operatorname{stab}_{x(\operatorname{A_{\circ}})}(w)$. Finally, part (c) may be used to relate directly the stable bases for any two adjacent alcoves, and therefore recursively relate the stable bases for two arbitrary alcoves.
We focus next on relating (36) to our Chevalley formulae obtained earlier in this paper. Together with (a) and (b) from Theorem 7.6 above, this gives an alternative recursion to (c), calculating the stable envelope for an arbitrary alcove $x\operatorname{A_{\circ}}+\lambda$ starting from the stable envelope for $\operatorname{A_{\circ}}$.
Fix $\lambda$ an integral weight. By Theorem 7.6(a), $\operatorname{stab}_{x(\operatorname{A_{\circ}})}(w)$ may be written as a linear combination of $\{\operatorname{stab}_{\operatorname{A_{\circ}}}(w)\mid w\in W\}$. By (36), to determine $\operatorname{stab}_{x(\operatorname{A_{\circ}})+\lambda}(w)$, it suffices to find a formula for ${\mathcal{L}}_{\lambda}\otimes\operatorname{stab}_{\operatorname{A_{\circ}}}(w)$. This can be achieved by the Chevalley formula for the motivic Chern classes. The key is the following result.
Lemma 7.7.
[AMSS19, Theorem 8.6] Let $\iota:G/B\hookrightarrow T^{*}(G/B)$ denote the inclusion of the zero section. For any $w\in W$,
$\iota^{*}(\operatorname{stab}_{\operatorname{A_{\circ}}}(w))=(-1)^{\dim G/B}q^{\dim G/B-\frac{\ell(w)}{2}}MC_{-q^{-1}}(Y(w)^{\circ})\otimes{\mathcal{L}}_{-2\rho}.$
Recall the operator $(-)^{\vee}$ on $K_{T}(pt)[y^{\pm 1}]$, which sends $e^{\mu}$ to $e^{-\mu}$ and $y$ to $y^{-1}$. We have the following Chevalley formula for the stable bases.
Theorem 7.8.
Let $\lambda\in X^{*}(T)$ be a weight and fix $\beta_{1},\ldots,\beta_{l}$ a reduced $\lambda$-chain corresponding to an alcove walk from $\operatorname{A_{\circ}}$ to $\operatorname{A_{\circ}}-\lambda$. Then
${\mathcal{L}}_{\lambda}\otimes\operatorname{stab}_{\operatorname{A_{\circ}}}(u)=\sum_{w}q^{\frac{\ell(u)-\ell(w)}{2}}(C_{u,-\lambda}^{w})^{\vee}|_{y=-q^{-1}}\operatorname{stab}_{\operatorname{A_{\circ}}}(w),$
where $C_{u,\lambda}^{w}$ are the coefficients defined in Equation 27.
One can write this in terms of $\lambda$-chains as follows (with the notation from Theorem 5.5):
${\mathcal{L}}_{\lambda}\otimes\operatorname{stab}_{\operatorname{A_{\circ}}}(u)=\sum_{J\subset\{1,2,\ldots,l\}}(-1)^{n(J)}(q^{-1/2}-q^{1/2})^{|J|}e^{w\tilde{r}_{J_{>}}(\lambda)}\operatorname{stab}_{\operatorname{A_{\circ}}}(ur_{h_{j_{1}}}r_{h_{j_{2}}}\cdots r_{h_{j_{t}}}),$
where the sum is over subsets $J=\{j_{1}<\ldots<j_{t}\}\subset\{1,2,\ldots,l\}$ such that $u<ur_{h_{j_{1}}}<ur_{h_{j_{1}}}r_{h_{j_{2}}}<\ldots<ur_{h_{j_{1}}}r_{h_{j_{2}}}\cdot\ldots\cdot r_{h_{j_{t}}}$.
Remark 7.9.
Recall we have the following duality between the stable bases (see [OS22, Proposition 1]):
$\langle\operatorname{stab}_{{\mathcal{C}},T^{\frac{1}{2}},A}(u),\operatorname{stab}_{-{\mathcal{C}},T_{\operatorname{opp}}^{\frac{1}{2}},-A}(w)\rangle=\delta_{u,w}.$
Therefore, similar arguments as in the proof of Lemma 5.1 will give a Chevalley formula for the dual stale basis $\operatorname{stab}_{-{\mathcal{C}},T(G/B)),-\operatorname{A_{\circ}}}(u)$.
Proof.
By the definition of the SMC classes from definition 4.2(2) , (28) becomes
${\mathcal{L}}_{\lambda}\otimes{\mathcal{D}}(MC_{y}(Y(u)^{\circ}))=\sum_{w\geq u}(-y)^{\ell(u)-\ell(w)}C_{u,\lambda}^{w}{\mathcal{D}}(MC_{y}(Y(w)^{\circ})).$
Taking $(-)^{\vee}$ on both sides of the equation, we get
${\mathcal{L}}_{-\lambda}\otimes MC_{y}(Y(u)^{\circ})=\sum_{w\geq u}(-y)^{\ell(w)-\ell(u)}(C_{u,\lambda}^{w})^{\vee}MC_{y}(Y(w)^{\circ}).$
Then the first equation of the theorem follows from this and Lemma 7.7. The second equation is a consequence of the first and of Equation 31 in Theorem 5.5. ∎
Example 7.10.
In type $A_{2}$, set $u=s_{2}s_{1}$ and $\lambda=2\varpi_{1}+\varpi_{2}$. A $\lambda$-chain of roots is given by $\beta_{1}=\alpha_{2},\;\beta_{2}=\alpha_{1}+\alpha_{2},\;\beta_{3}=\alpha_{1},\;\beta_{4}=\alpha_{1}+\alpha_{2},\;\beta_{5}=\alpha_{1},\;\beta_{6}=\alpha_{1}+\alpha_{2}$; see Appendix B. From Theorem 7.8, we have
$\begin{array}[]{lll}{\mathcal{L}}_{\lambda}\otimes\operatorname{stab}_{\operatorname{A_{\circ}}}(s_{2}s_{1})&=&e^{\varpi_{1}-3\varpi_{2}}\operatorname{stab}_{\operatorname{A_{\circ}}}(s_{2}s_{1})+(q^{-\frac{1}{2}}-q^{\frac{1}{2}})e^{-\varpi_{1}-2\varpi_{2}}\operatorname{stab}_{\operatorname{A_{\circ}}}(s_{1}s_{2}s_{1}).\end{array}$
Therefore by (36), using that $-u(\lambda)=-\varpi_{1}+3\varpi_{2}$, we have
$\begin{array}[]{lll}\operatorname{stab}_{\operatorname{A_{\circ}}+\lambda}(s_{2}s_{1})&=&\operatorname{stab}_{\operatorname{A_{\circ}}}(s_{2}s_{1})+(q^{-\frac{1}{2}}-q^{\frac{1}{2}})e^{-\alpha_{1}}\operatorname{stab}_{\operatorname{A_{\circ}}}(s_{1}s_{2}s_{1}).\end{array}$
Example 7.11.
Consider $u=s_{2}$, $\lambda=2\varpi_{1}+\varpi_{2}$, $w_{0}u=s_{2}s_{1}$.
We can use Serre duality (6.1) and Example 5.8 to get
$\begin{array}[]{lll}{\mathcal{L}}_{\lambda}\otimes\operatorname{stab}_{\operatorname{A_{\circ}}}(s_{2})&=&e^{3\varpi_{1}-\varpi_{2}}\operatorname{stab}_{\operatorname{A_{\circ}}}(s_{2})\\ &&+(q^{-\frac{1}{2}}-q^{\frac{1}{2}})(e^{\varpi_{1}}+e^{-\varpi_{1}+\varpi_{2}}+e^{-3\varpi_{1}+2\varpi_{2}})\operatorname{stab}_{\operatorname{A_{\circ}}}(s_{1}s_{2})\\ &&+(q^{-\frac{1}{2}}-q^{\frac{1}{2}})(e^{2\varpi_{1}-2\varpi_{2}}+e^{\varpi_{1}-3\varpi_{2}})\operatorname{stab}_{\operatorname{A_{\circ}}}(s_{2}s_{1})\\ &&+(q^{-\frac{1}{2}}-q^{\frac{1}{2}})^{2}(e^{-\varpi_{1}-2\varpi_{2}}+e^{-2\varpi_{1}}+e^{-\varpi_{2}})\operatorname{stab}_{\operatorname{A_{\circ}}}(s_{1}s_{2}s_{1}).\\ \end{array}$
As $-s_{2}(\lambda)=-3\varpi_{1}+\varpi_{2}$, we have
$\begin{array}[]{lll}\operatorname{stab}_{\operatorname{A_{\circ}}+\lambda}(s_{2})&=&\operatorname{stab}_{\operatorname{A_{\circ}}}(s_{2})\\ &&+(q^{-\frac{1}{2}}-q^{\frac{1}{2}})(e^{-\alpha_{1}}+e^{-2\alpha_{1}}+e^{-3\alpha_{1}})\operatorname{stab}_{\operatorname{A_{\circ}}}(s_{1}s_{2})\\ &&+(q^{-\frac{1}{2}}-q^{\frac{1}{2}})(e^{-\alpha_{1}-\alpha_{2}}+e^{-2\alpha_{1}-2\alpha_{2}})\operatorname{stab}_{\operatorname{A_{\circ}}}(s_{2}s_{1})\\ &&+(q^{-\frac{1}{2}}-q^{\frac{1}{2}})^{2}(e^{-3\alpha_{1}-2\alpha_{2}}+e^{-3\alpha_{1}-\alpha_{2}}+e^{-2\alpha_{1}-\alpha_{2}})\operatorname{stab}_{\operatorname{A_{\circ}}}(s_{1}s_{2}s_{1}).\\ \end{array}$
8. Whittaker functions and Hall–Littlewood polynomials
In this section we apply the Chevalley formula to obtain combinatorial expressions for Whittaker functions and Hall–Littlewood polynomials. Variants of the formulae we obtain were already available in the literature, and our approach based on the cohomological calculations adds a geometric perspective to this.
8.1. Whittaker functions
In this section we study Whittaker functions. These appear in $p$-adic representation theory, and in this note we utilize a cohomological construction of these functions from [AMSS19], see also [MSA22]. Recall the definition of the Demazure–Lusztig operators on $K_{T}(pt)[y]$:
$\widetilde{T_{i}}(e^{\lambda})=-e^{\lambda}\frac{1+y}{1-e^{-\alpha_{i}}}+e^{s_{i}\lambda}\frac{1+ye^{\alpha_{i}}}{1-e^{-\alpha_{i}}},$
and
$\widetilde{T_{i}^{\vee}}(e^{\lambda})=-e^{\lambda}\frac{1+y}{1-e^{-\alpha_{i}}}+e^{s_{i}\lambda}\frac{1+ye^{-\alpha_{i}}}{1-e^{-\alpha_{i}}}.$
The operators satisfy the usual quadratic and braid relations in the Hecke algebra, therefore for any $w\in W$ there are operators $\widetilde{T_{w}}$ and $\widetilde{T^{\vee}_{w}}$ acting on $K_{T}(pt)[y]$, and defined using any reduced decomposition of $w$. The following has been proved in [MSA22, Thm. 1.1]:
Proposition 8.1.
Let $\lambda_{y}(id):=\prod_{\alpha>0}(1+ye^{\alpha})$, and denote by
$MC^{\prime}_{y}(X(w)^{\circ}):=\lambda_{y}(id)\frac{MC_{y}(X(w)^{\circ})}{\lambda_{y}(T^{*}_{G/B})}\/.$
Then for any $\lambda\in X^{*}(T)$,
  1. (1)
    $\chi_{T}\left(G/B,{\mathcal{L}}_{\lambda}\otimes MC_{y}(X(w)^{\circ})\right)=\widetilde{T^{\vee}_{w}}(e^{\lambda})$;
  2. (2)
    $\chi_{T}\left(G/B,{\mathcal{L}}_{\lambda}\otimes MC_{y}^{\prime}(X(w)^{\circ})\right)=\widetilde{T_{w}}(e^{\lambda}).$
We note that for any anti-dominant weight $\lambda$ and $w\in W$,
(37) $\chi_{T}\left(G/B,{\mathcal{L}}_{\lambda}\otimes MC_{y}^{\prime}(X(w)^{\circ})\right)=\widetilde{T_{w}}(e^{\lambda})={\mathcal{W}}_{\lambda,w},$
where ${\mathcal{W}}_{\lambda,w}$ is the Iwahori–Whittaker function for the Langlands dual group over a non-archimedean local field; we refer to [MSA22] for more details, including the number theoretic definition of ${\mathcal{W}}_{\lambda,w}$. Here we identified $y$ with $-q^{-1}$, where $q$ is the number of elements in the residue field. As explained in [MSA22], from the fact that
$\sum_{w\in W}MC_{y}^{\prime}(X(w)^{\circ})=1$
by the additivity motivic Chern classes, one recovers the Casselman–Shalika formula for the spherical Whittaker function [CS80]:
(38) $\sum_{w\in W}{\mathcal{W}}_{\lambda,w}=\prod_{\alpha>0}(1+ye^{\alpha})\chi_{T}(G/B,{\mathcal{L}}_{\lambda})=\prod_{\alpha>0}(1+ye^{\alpha})\chi_{w_{0}\lambda}\/.$
Here $\chi_{w_{0}\lambda}$ denotes the character for the irreducible representation of $G$ of highest weight $w_{0}(\lambda)$. We also note that, in type A, an interpretation of the Iwahori–Whittaker function in terms of the partition function of the Iwahori lattice model has been obtained in [BBBG19].
Using the Chevalley coefficients in Equation 27, we obtain the following formula for the Iwahori–Whittaker function ${\mathcal{W}}_{\lambda,w}$. Let $\rho$ denote the half sum of the positive roots.
Theorem 8.2.
For any anti-dominant weight $\lambda$ and $w\in W$,
${\mathcal{W}}_{\lambda,w}=e^{\rho}\sum_{u}(-1)^{\ell(u)}C_{\lambda-\rho,u}^{w}|_{y\mapsto y^{-1}}y^{\ell(w)-\ell(u)}\/.$
Proof.
For any $u\in W$,
(39) $\chi_{T}(G/B,MC_{y}(X(u)^{\circ}))=MC_{y}[X(u)^{\circ}\rightarrow\operatorname{pt}]=MC_{y}[{\mathbb{A}}^{1}\rightarrow\operatorname{pt}]^{\ell(u)}=(-y)^{\ell(u)},$
where the second equality follows from [AMSS19, Theorem 4.2(3)]. Therefore, taking the equivariant Euler characteristics of both sides of Equation 27, we get
$\widetilde{T^{\vee}_{w}}(e^{\lambda})=\chi_{T}\left(G/B,{\mathcal{L}}_{\lambda}\otimes MC_{y}(X(w)^{\circ})\right)=\sum_{u}C_{\lambda,u}^{w}(-y)^{\ell(u)}.$
On the other hand, it is immediate to check the following relation between the two Demazure–Lusztig operators:
(40) $\widetilde{T_{w}}=e^{\rho}\widetilde{T^{\vee}_{w}}|_{y\mapsto y^{-1}}e^{-\rho}y^{\ell(w)}.$
Hence,
$\chi_{T}\left(G/B,{\mathcal{L}}_{\lambda}\otimes MC_{y}^{\prime}(X(w)^{\circ})\right)=\widetilde{T_{w}}(e^{\lambda})=e^{\rho}\sum_{u}(-1)^{\ell(u)}C_{\lambda-\rho,u}^{w}|_{y\mapsto y^{-1}}y^{\ell(w)-\ell(u)}\/.$
The proof ends by applying 8.1. ∎
Remark 8.3.
Notice that $C_{0,u}^{w}=\delta_{u,w}$. The above proof shows
$\chi_{T}\left(G/B,{\mathcal{L}}_{\rho}\otimes MC_{y}^{\prime}(X(w)^{\circ})\right)=(-1)^{\ell(w)}e^{\rho}.$
As a corollary we prove a variant of the Casselman–Shalika formula (38), obtained by Li [Li92], see also [BBBG19, Proposition 9.4]. First define
(41) $R_{\lambda}(y):=\chi_{T}(G/B,\lambda_{y}(T^{*}_{G/B})\otimes{\mathcal{L}}_{\lambda})\in K_{T}(pt)[y].$
Corollary 8.4.
Let $\lambda$ be an anti-dominant integral weight. Then
$\sum_{w}y^{-\ell(w)}{\mathcal{W}}_{\lambda,w}=e^{\rho}R_{\lambda-\rho}(y^{-1})\/.$
Proof.
By the additivity of the motivic Chern classes and Proposition 8.1(1),
(42) $R_{\lambda}(y)=\sum\chi_{T}\left(G/B,{\mathcal{L}}_{\lambda}\otimes MC_{y}(X(w)^{\circ})\right)=\sum_{w}\widetilde{T^{\vee}_{w}}(e^{\lambda}).$
On the other hand, for an anti-dominant weight $\lambda$, we have
$\sum_{w}y^{-\ell(w)}{\mathcal{W}}_{\lambda,w}$ $=\sum_{w}y^{-\ell(w)}\widetilde{T_{w}}(e^{\lambda})$
$=e^{\rho}\sum_{w}y^{-\ell(w)}e^{-\rho}\widetilde{T_{w}}e^{\rho}(e^{\lambda-\rho})$
$=e^{\rho}\sum_{w}\widetilde{T^{\vee}_{w}}|_{y\mapsto y^{-1}}(e^{\lambda-\rho})$
$=e^{\rho}R_{\lambda-\rho}(y^{-1}).$
Here, the first equality follows from 8.1(2), the third one follows from Equation 40, and the last one follows from Equation 42. ∎
8.2. Hall–Littlewood polynomials
In this section, we assume either $\lambda$ or $-\lambda$ to be a dominant integral weight. Set $\Sigma_{\lambda}:=\{\alpha\in\Sigma\mid\langle\lambda,\alpha^{\vee}\rangle=0\}$, and let $R^{+}_{\lambda}$ be the set of positive roots which are linear combinations of the simple roots in $\Sigma_{\lambda}$. Denote by $W_{\lambda}\subset W$ the subgroup generated by the simple reflections $s_{\alpha}$, where $\alpha\in\Sigma_{\lambda}$. Let $W^{\lambda}$ be the set of minimal length representatives for the cosets $W/W_{\lambda}$. Finally, let $P_{\lambda}$ be the parabolic subgroup containing the Borel subgroup $B$ defined by the condition that $W_{P_{\lambda}}=W_{\lambda}$.
Definition 8.5.
  1. (1)
    Define $H_{\lambda}(y):=\chi_{T}(G/P_{\lambda},\lambda_{y}(T^{*}_{G/P_{\lambda}})\otimes{\mathcal{L}}_{\lambda})\in K_{T}(pt)[y]$.
  2. (2)
    (Hall–Littlewood polynomial cf. [Mac98, P.208 (2.2)]) For a dominant weight $\lambda$, define
    $HL_{\lambda}({\mathbf{x}};t):=\sum_{w\in W^{\lambda}}w\left({\mathbf{x}}^{\lambda}\prod_{\alpha\in R^{+}\setminus R^{+}_{\lambda}}\frac{1-t{\mathbf{x}}^{-\alpha}}{1-{\mathbf{x}}^{-\alpha}}\right)$
    where ${\mathbf{x}}^{\lambda}$ denotes $e^{\lambda}$.
Let $\pi_{\lambda}:G/B\to G/P_{\lambda}$ be the natural projection. Then
$\lambda_{y}(G/B)=\pi_{\lambda}^{*}(\lambda_{y}(G/P_{\lambda}))\cdot\lambda_{y}(P_{\lambda}/B)\/.$
By the projection formula, and using that $\pi_{\lambda}^{*}(\mathcal{L}_{\lambda})=\mathcal{L}_{\lambda}$, we have that
$H_{\lambda}(y)=\chi_{T}(G/P_{\lambda},\lambda_{y}(T^{*}_{G/P_{\lambda}})\otimes{\mathcal{L}}_{\lambda})=\frac{\chi_{T}(G/B,\lambda_{y}(T^{*}_{G/B})\otimes{\mathcal{L}}_{\lambda})}{\chi_{T}(P_{\lambda}/B,\lambda_{y}(T^{*}_{P_{\lambda}/B}))}=\frac{R_{\lambda}(y)}{\sum_{w\in W_{\lambda}}(-y)^{\ell(w)}}$
Here the last equality follows from Equation 41, and the fact that
$\chi_{T}(P_{\lambda}/B,\lambda_{y}(T^{*}_{P_{\lambda}/B}))=\sum_{w\in W_{\lambda}}\chi(MC_{y}(X(w)^{\circ}))=\sum_{w\in W_{\lambda}}(-y)^{\ell(w)}\/.$
The relation between $H_{\lambda}$ and the Hall–Littlewood polynomial, summarized next, was obtained in a related (upcoming) collaboration with B. Ion.
Lemma 8.6.
We have the following formulae for $H_{\lambda}(y)$:
(43) $H_{\lambda}(y)=\sum_{w\in W^{P}}\sum_{u\in W}C^{w}_{u,\lambda}(-y)^{\ell(u)},$
and
(44) $H_{\lambda}(y)=\displaystyle\sum_{w\in W^{\lambda}}w\left(e^{\lambda}\prod_{\alpha\in R^{+}\setminus R^{+}_{\lambda}}\frac{1+ye^{\alpha}}{1-e^{\alpha}}\right).$
Proof.
Equation 43 follows because $\lambda_{y}(T^{*}_{G/P})=MC_{y}(G/P)=\displaystyle\sum_{w\in W^{P}}MC_{y}(X(wW_{P})^{\circ})$, Theorem 5.15, and $\chi_{T}(MC_{y}(X(uW_{P})^{\circ}))=(-y)^{\ell(u)}$ for any $u\in W^{P}$. Equation 44 follows from the localization formula (cf. [Nie74], [MSA22, Theorem 2.1 (c)].) To be more specific, the torus fixed points in $G/P_{\lambda}$ are $\{wP_{\lambda}\mid w\in W^{P}\}$, and the torus weights of the tangent space at the fixed point $wP_{\lambda}$ are $\{-w\alpha\mid\alpha\in R^{+}\setminus R^{+}_{\lambda}\}$. ∎
Corollary 8.7.
For a dominant integral $\lambda$, the Hall–Littlewood polynomial $HL_{\lambda}({x};t)$ can be expressed using $H_{-\lambda}(y)$ or $H_{\lambda}(y)$ as follows:
(45) $HL_{\lambda}({\mathbf{x}};t)=H_{-\lambda}(y)|_{e^{\alpha}\mapsto{\mathbf{x}}^{-\alpha},y\mapsto-t},$
and
(46) $HL_{\lambda}({\mathbf{x}};t)=\left(\frac{1}{(-y)^{\dim G/P_{\lambda}}}H_{\lambda}(y)\right)|_{e^{\alpha}\mapsto{\mathbf{x}}^{\alpha},y\mapsto-t^{-1}}.$
We assume next that $\lambda$ is a dominant integral weight. Fix a reduced $(-\lambda)$-chain $\Gamma=(-\beta_{1},-\beta_{2},\cdots,-\beta_{l})$ and the sequence of hyperplanes $H_{\beta_{1},d_{1}},H_{\beta_{2},d_{2}},\ldots,H_{\beta_{l},d_{l}}$. This corresponds to an alcove path from $\operatorname{A_{\circ}}$ to $\operatorname{A_{\circ}}+\lambda$. Since $\lambda$ is dominant, $\beta_{j}>0,d_{j}>0$.
Then we can recover the following known formula for the Hall–Littlewood polynomial, see [Len11, Theorem 2.7].
Proposition 8.8.
[Sch06, Ram06, Len11]
(47) $HL_{\lambda}({\bf x};t)=\sum_{(w,J,u)\in\mathcal{A}(\Gamma)}t^{\frac{\ell(w)+\ell(u)-|J|}{2}}(1-t)^{|J|}{\bf x}^{w\hat{r}_{J_{<}}(\lambda)},$
where (with notation as in §3.2),
$\mathcal{A}(\Gamma)=\{(w,J,u)\mid w\in W^{P},u\in W,J\subset\{1,2,\ldots,l\},u\stackrel{{\scriptstyle J_{>}}}{{\longrightarrow}}w\}\/.$
Proof.
This is a direct consequence of Equation 45, Equation 43, and Equation 32 from Theorem 5.5, applied to the $(-\lambda)$-chain $(-\beta_{1},-\beta_{2},\cdots,-\beta_{l})$. ∎
We also get a new formula for $HL_{\lambda}({\mathbf{x}};t)$ as follows.
Proposition 8.9.
(48) $HL_{\lambda}({\bf x};t)=\sum_{(u,J,w)\in\;\mathcal{A}^{op}(\Gamma)}t^{\frac{2\dim G/P_{\lambda}-\ell(w)-\ell(u)-|J|}{2}}(1-t)^{|J|}{\bf x}^{u\hat{r}_{J_{<}}(\lambda))},$
where $\mathcal{A}^{op}(\Gamma)=\{(u,J,w)\mid w\in W^{P},u\in W,J\subset\{1,2,\ldots,l\},u\stackrel{{\scriptstyle J_{<}}}{{\longrightarrow}}w\}$.
Proof.
This is a direct consequence of Equation 46, Equation 43, and Equation 31 from Theorem 5.5, applied to the $(-\lambda)$-chain together with Remark 5.6 (1). ∎
Remark 8.10.
When $P_{\lambda}=B$, the equations (47) and (48) give the same formula. The correspondence may be seen using the Serre duality in Equation 35. However, the formulae are in general different, as shown by the examples below.
Example 8.11.
(Type $A_{2}$) Let $G=GL_{3}(\mathbb{C})$, $T=(\mathbb{C}^{*})^{3}$, and $x_{i}=e^{\varepsilon_{i}}$, for $i=1,2,3$. Let $\lambda=\varpi_{1}=\varepsilon_{1}$, then $W_{\lambda}=\langle s_{2}\rangle\subset W=\langle s_{1},s_{2}\rangle$, and $W^{\lambda}=\{id,s_{1},s_{2}s_{1}\}$. Fix a reduced $(-\lambda)$-chain is $(-\beta_{1}=-\alpha_{1}-\alpha_{2},-\beta_{2}=-\alpha_{1})$.
8.8 sums over the following seven terms:
$\begin{array}[]{c|c|c|r}w&J&u&\\ \hline\cr\hline\cr id&\emptyset&id&x_{1}\\ \hline\cr s_{1}&\emptyset&s_{1}&tx_{2}\\ &\{2\}&id&(1-t)x_{2}\\ \hline\cr\end{array}$          $\begin{array}[]{c|c|c|r}\hline\cr s_{2}s_{1}&\emptyset&s_{2}s_{1}&t^{2}x_{3}\\ &\{1\}&s_{1}&t(1-t)x_{3}\\ &\{2\}&s_{2}&t(1-t)x_{3}\\ &\{1,2\}&id&(1-t)^{2}x_{3}\\ \hline\cr\end{array}$.
Hence, with ${\mathbf{x}}=(x_{1},x_{2},x_{3})$,
$HL_{\lambda}({\mathbf{x}};t)=(x_{1})+(tx_{2}+(1-t)x_{2})+(t^{2}x_{3}+t(1-t)x_{3}+t(1-t)x_{3}+(1-t)^{2}x_{3})=x_{1}+x_{2}+x_{3}.$
On the other hand, 8.9 sums over the following six terms:
$\begin{array}[]{c|c|c|r}w&J&u&\\ \hline\cr\hline\cr id&\emptyset&id&t^{2}x_{1}\\ \hline\cr s_{1}&\emptyset&s_{1}&tx_{2}\\ &\{2\}&id&t(1-t)x_{1}\\ \hline\cr\end{array}$           $\begin{array}[]{c|c|c|r}\hline\cr s_{2}s_{1}&\emptyset&s_{2}s_{1}&x_{3}\\ &\{1\}&s_{1}&(1-t)x_{2}\\ &\{2\}&s_{2}&(1-t)x_{1}\\ \hline\cr\end{array}$.
Hence,
$HL_{\lambda}({\mathbf{x}};t)=(t^{2}x_{1})+(tx_{2}+t(1-t)x_{1})+(x_{3}+(1-t)x_{2}+(1-t)x_{1}))=x_{1}+x_{2}+x_{3}.$
Example 8.12.
Same setup as in the above example, but with $\lambda=2\varpi_{2}$. Then $W_{\lambda}=\langle s_{1}\rangle$, $W^{\lambda}=\{id,s_{2},s_{1}s_{2}\}$, and $d=\dim G/P_{\lambda}=2$. A $(-\lambda)$-chain is $(-\beta_{1}=-(\alpha_{1}+\alpha_{2}),-\beta_{2}=-\alpha_{2},-\beta_{3}=-(\alpha_{1}+\alpha_{2}),-\beta_{4}=-\alpha_{2})$ , and $d_{1}=1,d_{2}=1,d_{3}=2,d_{4}=2$.
8.8 sums over the following twelve terms:
$\begin{array}[]{c|c|c|r}w&J&u&\\ \hline\cr\hline\cr id&\emptyset&id&x_{1}^{2}x_{2}^{2}\\ \hline\cr s_{2}&\emptyset&s_{1}&tx_{1}^{2}x_{3}^{2}\\ &\{2\}&id&(1-t)x_{1}^{2}x_{2}x_{3}\\ &\{4\}&id&(1-t)x_{1}^{2}x_{3}^{2}\\ \hline\cr\end{array}$           $\begin{array}[]{c|c|c|r}\hline\cr w=s_{1}s_{2}&\emptyset&s_{1}s_{2}&t^{2}x_{2}^{2}x_{3}^{2}\\ &\{1\}&s_{2}&t(1-t)x_{1}x_{2}x_{3}^{2}\\ &\{2\}&s_{1}&t(1-t)x_{1}x_{2}^{2}x_{3}\\ &\{3\}&s_{2}&t(1-t)x_{2}^{2}x_{3}^{2}\\ &\{4\}&s_{1}&t(1-t)x_{2}^{2}x_{3}^{2}\\ &\{1,2\}&id&(1-t)^{2}x_{1}x_{2}^{2}x_{3}\\ &\{1,4\}&id&(1-t)^{2}x_{1}x_{2}x_{3}^{2}\\ &\{3,4\}&id&(1-t)^{2}x_{2}^{2}x_{3}^{2}\\ \hline\cr\end{array}.$
We get $HL_{\lambda}({\mathbf{x}};t)=s_{22}-ts_{211}$, where
$s_{22}=x_{1}^{2}x_{2}^{2}+x_{1}^{2}x_{3}^{2}+x_{2}^{2}x_{3}^{2}+x_{1}^{2}x_{2}x_{3}+x_{1}x_{2}^{2}x_{3}+x_{1}x_{2}x_{3}^{2}\/,$
and
$s_{211}=x_{1}^{2}x_{2}x_{3}+x_{1}x_{2}^{2}x_{3}+x_{1}x_{2}x_{3}^{2}\/.$
The 8.9 sums over the following ten terms:
$\begin{array}[]{c|c|c|r}w&J&u&\\ \hline\cr\hline\cr id&\emptyset&id&t^{2}x_{1}^{2}x_{2}^{2}\\ \hline\cr s_{2}&\emptyset&s_{2}&tx_{1}^{2}x_{3}^{2}\\ &\{2\}&id&t(1-t)x_{1}^{2}x_{2}x_{3}\\ &\{4\}&id&t(1-t)x_{1}^{2}x_{2}^{2}\\ \hline\cr\end{array}$          $\begin{array}[]{c|c|c|r}\hline\cr s_{1}s_{2}&\emptyset&s_{1}s_{2}&x_{2}^{2}x_{3}^{2}\\ &\{1\}&s_{2}&(1-t)x_{1}x_{2}x_{3}^{2}\\ &\{2\}&s_{1}&(1-t)x_{1}x_{2}^{2}x_{3}\\ &\{3\}&s_{2}&(1-t)x_{1}^{2}x_{3}^{2}\\ &\{4\}&s_{1}&(1-t)x_{1}^{2}x_{2}^{2}\\ &\{2,3\}&id&(1-t)^{2}x_{1}^{2}x_{2}x_{3}\\ \hline\cr\end{array}.$
The summation also gives $HL_{\lambda}({\mathbf{x}};t)=(x_{1}^{2}x_{2}^{2}+x_{1}^{2}x_{3}^{2}+x_{2}^{2}x_{3}^{2})+(1-t)(x_{1}^{2}x_{2}x_{3}+x_{1}x_{2}^{2}x_{3}+x_{1}x_{2}x_{3}^{2})=s_{22}-ts_{211}$.
Appendix A Chevalley formulae for the Chern–Schwartz–MacPherson classes of Schubert cells
In this appendix, we give a short proof of the Chevalley formulae for the equivariant Chern–Schwartz–MacPherson (CSM) and Segre–MacPherson (SM) classes of the Schubert cells in the (partial) flag varieties, see [AMSS17, Thm  9.10] and Theorem A.4 below. Our proof again relies on the action of the appropriate Hecke algebra, this time on the equivariant cohomology of $G/P$. Besides the intrinsic interest in these Chevalley formulae, we mention that recursions based on it were recently utilized to obtain proofs of Nakada’s colored hook formula for finite Weyl groups [MNS22a] and more general Coxeter groups [MNS24].
A.1. Degenerate affine Hecke algebra
A.1.1. A change of bases formula
Recall that the degenerate affine Hecke algebra ${\mathcal{H}}$ is generated by $T_{w},w\in W$ and $x_{\lambda}$, $\lambda\in X^{*}(T)$, such that
  • $T_{w}T_{u}=T_{wu}$ for any $w,u\in W$;
  • $x_{\lambda}x_{\mu}=x_{\mu}x_{\lambda}$ for any $\lambda,\mu\in X^{*}(T)$;
  • $x_{\lambda+\mu}=x_{\lambda}+x_{\mu}$ for any $\lambda,\mu\in X^{*}(T)$;
  • for any simple root $\alpha_{i}$, $T_{s_{i}}x_{\lambda}-x_{s_{i}\lambda}T_{s_{i}}=-\langle\lambda,\alpha_{i}^{\vee}\rangle$.
We have the following commutation relation.
Lemma A.1.
For any $w\in W$ and $\lambda\in X^{*}(T)$, the following holds in ${\mathcal{H}}$
$T_{w}x_{\lambda}=x_{w\lambda}T_{w}-\sum_{\alpha>0,ws_{\alpha}<w}\langle\lambda,\alpha^{\vee}\rangle T_{ws_{\alpha}}.$
Proof.
We utilize induction on $\ell(w)$. The claim is clear when $\ell(w)=0$ or $1$. Now assume $\ell(w)>1$, and that the claim holds for any Weyl group element with length smaller than $\ell(w)$. Pick a simple root $\alpha_{i}$, such that $w>ws_{i}$. By induction,
$T_{ws_{i}}x_{s_{i}\lambda}=x_{w\lambda}T_{ws_{i}}-\sum_{\alpha>0,ws_{i}s_{\alpha}<ws_{i}}\langle s_{i}\lambda,\alpha^{\vee}\rangle T_{ws_{i}s_{\alpha}}.$
Using the commutation of $T_{s_{i}}$ and $x_{\lambda}$ and the induction hypothesis we obtain:
$\begin{split}T_{w}x_{\lambda}-x_{w\lambda}T_{w}=&\quad T_{ws_{i}}x_{s_{i}\lambda}T_{s_{i}}-\langle\lambda,\alpha_{i}^{\vee}\rangle T_{ws_{i}}-x_{w\lambda}T_{w}\\ =&-\sum_{\alpha>0,ws_{i}s_{\alpha}<ws_{i}}\langle\lambda,s_{i}\alpha^{\vee}\rangle T_{ws_{i}s_{\alpha}}T_{s_{i}}-\langle\lambda,\alpha_{i}^{\vee}\rangle T_{ws_{i}}\\ =&-\sum_{\alpha>0,ws_{\alpha}<w}\langle\lambda,\alpha^{\vee}\rangle T_{ws_{\alpha}}.\end{split}$
Here in the last equality, we have used the facts that if $ws_{i}<w$ then
$\{\alpha>0\mid ws_{\alpha}<w\}$ =$=\{\alpha>0\mid w\alpha<0\}$
=$=\{\alpha_{i}\}\sqcup\{s_{i}\alpha\mid\alpha>0,ws_{i}s_{\alpha}<ws_{i}\},$
and $T_{ws_{i}s_{\alpha}}T_{s_{i}}=T_{ws_{i}s_{\alpha}s_{i}}=T_{ws_{s_{i}(\alpha)}}$. ∎
A.1.2. The Hecke action on the equivariant cohomology
Since $G$ acts on $G/P$ by left multiplication, there is a natural left Weyl group action on $H_{T}^{*}(G/P)$ for any partial flag variety $G/P$ and which acts on the base ring $H_{T}^{*}(\operatorname{pt})$ by the usual Weyl group action; see e.g. [MNS22b]. For any $w\in W$, we use $w^{L}$ to denote this action.
For any simple root $\alpha_{i}$, define the left Demazure–Lusztig operator on $H_{T}^{*}(G/P)$ by the following formula (see [MNS22b, Section 3.2]):
${\mathcal{T}}_{i}^{L}:=\frac{\alpha_{i}+1}{\alpha_{i}}s_{i}^{L}-\frac{1}{\alpha_{i}}.$
Then it is proved in loc. cit. that these operators satisfy the braid relations and $({\mathcal{T}}_{i}^{L})^{2}=\operatorname{id}$. Moreover, it is immediate to check the following lemma.
Lemma A.2.
There is an action $\Psi$ of the degenerate affine Hecke algebra ${\mathcal{H}}$ on $H_{T}^{*}(G/P)$, sending $T_{i}$ to ${\mathcal{T}}_{i}^{L}$ and $x_{\lambda}$ to $\lambda\in H_{T}^{*}(\operatorname{pt})$.
A.2. Definition of the CSM/SM classes
Next we recall the basic definitions and properties of CSM classes for the Schubert cells in $G/P$, we will be brief. We refer the reader e.g. to [Ohm06, AMSS17] for details, including a construction of these classes in the equivariant setting and for general varieties.
[…]
There is an equivariant version of MacPherson’s transformation defined by Ohmoto [Ohm06]. In this case one starts with a variety $X$ with a $T$-action, and the equivariant version ${\mathcal{F}}^{T}(X)$ of the group of constructible functions ${\mathcal{F}}(X)$ contains the characteristic functions $1\hskip-3.5pt1_{Z}$ for $Z$ stable under the $T$-action. If $f:X\to Y$ is a proper $T$-equivariant morphism of algebraic varieties the induced homomorphism and $Z\subset X$ is constructible and $T$-stable then one defines $f_{*}^{T}:{\mathcal{F}}^{T}(X)\to{\mathcal{F}}^{T}(Y)$ with the property that $f_{*}^{T}(1\hskip-3.5pt1_{Z})=f_{*}(1\hskip-3.5pt1_{Z})$. Ohmoto proves [Ohm06, Theorem 1.1] that there is an equivariant version of MacPherson transformation $c_{*}^{T}:{\mathcal{F}}^{T}(X)\to H_{*}^{T}(X)$ that satisfies $c_{*}^{T}(1\hskip-3.5pt1_{X})=c^{T}(T_{X})\cap[X]_{T}$ if $X$ is a non-singular $T$-variety, which and is functorial with respect to proper push-forwards. The last statement means that for all proper $T$-equivariant morphisms $Y\to X$ the following diagram commutes:
[…]
Definition A.3.
Let $Z$ be a $T$-invariant constructible subvariety of $X$.
  1. (1)
    We denote by $c_{\mathrm{SM}}(Z):=c_{*}^{T}(1\hskip-3.5pt1_{Z}){~{}\in H_{*}^{T}(X)}$ the equivariant Chern–Schwartz–MacPherson (CSM) class of $Z$.
  2. (2)
    If $X$ is smooth, we denote by $s_{\mathrm{M}}(Z\subset X):=\frac{c_{*}^{T}(1\hskip-3.5pt1_{Z})}{c(T_{X})}{~{}\in H_{*}^{T}(X)_{\operatorname{loc}}}$ the equivariant Segre–MacPherson (SM) class of $Z$, where $H_{*}^{T}(X)_{\operatorname{loc}}:=H_{*}^{T}(X)\otimes_{H_{*}^{T}(pt)}\operatorname{Frac}H_{*}^{T}(pt)$ denotes the localization of $H_{*}^{T}(X)$, and $\operatorname{Frac}H_{*}^{T}(pt)$ is the fraction field of $H_{*}^{T}(pt)$.
A.3. The Chevalley formula in cohomology
We now specialize to $X=G/P$ with the usual $T$-action. For simplicity we will denote by $s_{\mathrm{M}}(Z\subset G/P)$ simply by $s_{\mathrm{M}}(Z)$. We will identify the equivariant (Borel-Moore) homology group $H_{*}^{T}(G/P)$ with the equivariant cohomology $H^{*}_{T}(G/P)$, using the Poincaré duality. The sets of CSM classes of Schubert cells $\{c_{\mathrm{SM}}(X(wW_{P})^{\circ})\mid w\in W^{P}\}$ and of the SM classes $\{s_{\mathrm{M}}(X(wW_{P})^{\circ})\mid w\in W^{P}\}$ form bases for $H_{T}^{*}(G/P)_{\operatorname{loc}}:=H_{T}^{*}(G/P)\otimes_{H_{T}^{*}(pt)}\operatorname{Frac}H_{T}^{*}(pt)$. Moreover, if one takes the opposite Schubert cells in any of these sets, then the two bases are dual under the usual intersection pairing, see [AMSS17, Theorem 9.4]:
(49) $\langle c_{\mathrm{SM}}(X(wW_{P})^{\circ}),s_{\mathrm{M}}(Y(uW_{P})^{\circ})\rangle_{G/P}=\delta_{w,u}\textit{ for any }w,u\in W^{P}.$
The left Demazure–Lusztig operator acts on the CSM classes by the following formula (see [MNS22b, Theorem 4.3])
${\mathcal{T}}_{i}^{L}(c_{\mathrm{SM}}(X(wW_{P})^{\circ}))=c_{\mathrm{SM}}(X(s_{i}wW_{P})^{\circ}).$
Hence, for any $w\in W$,
(50) $c_{\mathrm{SM}}(X(wW_{P})^{\circ})={\mathcal{T}}_{w}^{L}([X(\operatorname{id})]).$
Recall for any $\lambda\in X^{*}(T)_{P}$, ${\mathcal{L}}_{\lambda}:=G\times^{P}{\mathbb{C}}_{\lambda}\in\operatorname{Pic}_{T}(G/P)$. The following is our main result in this Appendix, and it has also been proved in [AMSS17, Thm. 9.10] using the Chevalley formula for the cohomological stable envelopes from [Su16, Thm. 3.7]. Here we give a direct proof based on the action of the degenerate affine Hecke algebra.
Theorem A.4.
For any $w\in W^{P}$ and $\lambda\in X^{*}(T)_{P}$, the following holds in $H_{T}^{*}(G/P)$:
$c_{1}^{T}({\mathcal{L}}_{\lambda})\cup c_{\mathrm{SM}}(X(wW_{P})^{\circ})=w(\lambda)c_{\mathrm{SM}}(X(wW_{P})^{\circ})-\sum_{\alpha>0,ws_{\alpha}<w}\langle\lambda,\alpha^{\vee}\rangle c_{\mathrm{SM}}(X(ws_{\alpha}W_{P})^{\circ}),$
and
$c_{1}^{T}({\mathcal{L}}_{\lambda})\cup s_{\mathrm{M}}(Y(wW_{P})^{\circ})=w(\lambda)s_{\mathrm{M}}(Y(wW_{P})^{\circ})-\sum_{\alpha>0,ws_{\alpha}>w}\langle\lambda,\alpha^{\vee}\rangle s_{\mathrm{M}}(Y(ws_{\alpha}W_{P})^{\circ})\/.$
Proof.
Applying the Hecke action $\Psi$ in Lemma A.2 to the equation in Lemma A.1, and acting on the point class $[X(\operatorname{id})]$, we get
$c_{1}^{T}({\mathcal{L}}_{\lambda})\cup c_{\mathrm{SM}}(X(wW_{P})^{\circ})=$ $c_{1}^{T}({\mathcal{L}}_{\lambda})\cup{\mathcal{T}}^{L}_{w}([X(\operatorname{id})])$
$=$ ${\mathcal{T}}^{L}_{w}(c_{1}^{T}({\mathcal{L}}_{\lambda})\cup[X(\operatorname{id})])$
$=$ ${\mathcal{T}}^{L}_{w}(\lambda\cdot[X(\operatorname{id})])$
$=$ $\Psi(T_{w}x_{\lambda})([X(\operatorname{id})])$
$=$ $\Psi(x_{w\lambda}T_{w}-\sum_{\alpha>0,ws_{\alpha}<w}\langle\lambda,\alpha^{\vee}\rangle T_{ws_{\alpha}})([X(\operatorname{id})])$
$=$ $w(\lambda)c_{\mathrm{SM}}(X(wW_{P})^{\circ})-\sum_{\alpha>0,ws_{\alpha}<w}\langle\lambda,\alpha^{\vee}\rangle c_{\mathrm{SM}}(X(ws_{\alpha}W_{P})^{\circ}).$
The second equality follows from the fact that the left operator ${\mathcal{T}}_{w}^{L}$ commutes with $c_{1}^{T}({\mathcal{L}}_{\lambda})$ because the latter is Weyl-group invariant, as ${\mathcal{L}}_{\lambda}$ is a $G$-equivariant line bundle; see [MNS22b]. Finally, the Chevalley formula for the SM classes follows from the one on CSM via the duality in Equation 49, similar to the proof of Lemma 5.1 above. ∎
Appendix B An example of the $\lambda$-chain formula
We consider Lie type $A_{2}$, with the Weyl group $W=S_{3}$, with $\lambda=2\varpi_{1}+\varpi_{2}$, and $w=s_{2}s_{1}$. We can find an alcove walk $p_{-\lambda}$ from $\operatorname{A_{\circ}}$ to $\operatorname{A_{\circ}}-\lambda$ as indicated by the red path in diagram (b). This gives the corresponding reduced expression $v_{-\lambda}=s_{2}s_{1}s_{2}s_{0}s_{1}s_{2}$ and the corresponding alcove path
$\operatorname{A_{\circ}}=A_{0}\stackrel{{\scriptstyle-\beta_{1}}}{{\longrightarrow}}A_{1}\stackrel{{\scriptstyle-\beta_{2}}}{{\longrightarrow}}A_{2}\stackrel{{\scriptstyle-\beta_{3}}}{{\longrightarrow}}A_{3}\stackrel{{\scriptstyle-\beta_{4}}}{{\longrightarrow}}A_{4}\stackrel{{\scriptstyle-\beta_{5}}}{{\longrightarrow}}A_{5}\stackrel{{\scriptstyle-\beta_{6}}}{{\longrightarrow}}A_{6}=\operatorname{A_{\circ}}-\lambda\;(A_{i}=r_{i}A_{i-1},1\leq i\leq 6),$
with $\lambda$-chain $\beta_{1}=\alpha_{2},\;\beta_{2}=\alpha_{1}+\alpha_{2},\;\beta_{3}=\alpha_{1},\;\beta_{4}=\alpha_{1}+\alpha_{2},\;\beta_{5}=\alpha_{1},\;\beta_{6}=\alpha_{1}+\alpha_{2}$.
The associated sequence of hyperplanes is
$h_{1}=H_{\alpha_{2},0},h_{2}=H_{\alpha_{1}+\alpha_{2},0},h_{3}=H_{\alpha_{1},0},h_{4}=H_{\alpha_{1}+\alpha_{2},-1},h_{5}=H_{\alpha_{1},-1},h_{6}=H_{\alpha_{1}+\alpha_{2},-2}.$
(a) Alcove walk $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}p_{\lambda}$ from $\operatorname{A_{\circ}}$ to $\operatorname{A_{\circ}}+\lambda$$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}p_{w}=c^{-}_{2}c^{-}_{1}$ , $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}p_{\lambda}=c^{+}_{0}c^{+}_{2}c^{+}_{1}c^{+}_{0}c^{+}_{2}c^{+}_{0}$
[…]
(a) Alcove walk $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}p_{\lambda}$ from $\operatorname{A_{\circ}}$ to $\operatorname{A_{\circ}}+\lambda$$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}p_{w}=c^{-}_{2}c^{-}_{1}$ , $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}p_{\lambda}=c^{+}_{0}c^{+}_{2}c^{+}_{1}c^{+}_{0}c^{+}_{2}c^{+}_{0}$
(b) Alcove walk $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}p_{-\lambda}$ from $\operatorname{A_{\circ}}$ to $\operatorname{A_{\circ}}-\lambda$$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}p_{w}=c^{-}_{2}c^{-}_{1}$, $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}p_{-\lambda}=c^{-}_{2}c^{-}_{1}c^{-}_{2}c^{-}_{0}c^{-}_{1}c^{-}_{2}$
We first calculate $c_{u,\mu}^{w,\lambda}$ according to the formula Theorem 3.10. In the proof we introduced $\lambda$-shifted (reversed) hyperplane sequence which can be seen in diagram (a).
$h^{\prime}_{1}=H_{\alpha_{1}+\alpha_{2},1},h^{\prime}_{2}=H_{\alpha_{1},1},h^{\prime}_{3}=H_{\alpha_{1}+\alpha_{2},2},h^{\prime}_{4}=H_{\alpha_{1},2},,h^{\prime}_{5}=H_{\alpha_{1}+\alpha_{2},3},h^{\prime}_{6}=H_{\alpha_{2},1}.$
According to the formula, we need to choose $J\subset\{1,2,\ldots,l\}$ such that $u\stackrel{{\scriptstyle J_{<}}}{{\longrightarrow}}w$.
If $u=s_{1}$, $J=\{6\},\{4\},\{2\}$ as $s_{\alpha_{1}+\alpha_{2}}=s_{1}s_{2}s_{1}$. For the weight $\mu$ , when $J=\{6\}$, we need to calculate $\mu=w\tilde{r}_{h_{6}}(\lambda)$. But as is explained in the proof, $\tilde{r}_{h_{6}}=\hat{r}_{h^{\prime}_{1}}$, so, $\mu=w\hat{r}_{h^{\prime}_{1}}(\lambda)=w(-\varpi_{2})=-\varpi_{1}+\varpi_{2}$. Likewise when $J=\{4\}$, $\mu=w\tilde{r}_{h_{4}}(\lambda)=w\hat{r}_{h^{\prime}_{3}}(\lambda)=-\varpi_{2}$, and when $J=\{2\}$, $\mu=w\tilde{r}_{h_{2}}(\lambda)=w\hat{r}_{h^{\prime}_{5}}(\lambda)=\varpi_{1}-3\varpi_{2}$. If $u=id$, there are two possible Bruhat chains $u=id<s_{1}<s_{2}s_{1}=w$ and $u=id<s_{2}<s_{2}s_{1}=w$. For the first case, $J=\{5,6\},\{3,6\},\{3,4\}$, and for the second case, $J=\{1,5\},\{1,3\}$. When $J=\{5,6\}$, $\mu=w\tilde{r}_{h_{6}}\tilde{r}_{h_{5}}(\lambda)=w\hat{r}_{h^{\prime}_{1}}\hat{r}_{h^{\prime}_{2}}(\lambda)=w\hat{r}_{h^{\prime}_{1}}(2\varpi_{2})=w(-\varpi_{1}+\varpi_{2})=\varpi_{1}$. All the possibilities and the corresponding (folded) alcove walks $\tilde{p}_{\lambda}$ are listed in the table below. (We can also see the bijection of Lemma 2.7, cf. diagram (a).)
$\begin{array}[]{|c|c|c|c|c|c|}\hline\cr p_{w}\tilde{p}_{\lambda}&\tilde{p}_{\lambda}&\mathcal{M}&J&\varphi(p)=u&{\rm wt}(p)=w\tilde{r}^{\lambda}_{J_{>}}(\lambda)\\ \hline\cr\hline\cr(0)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{0}c^{-}_{2}c^{-}_{1}c^{-}_{0}c^{-}_{2}c^{+}_{0}}&\{\}&\{\}&s_{2}s_{1}&\varpi_{1}-3\varpi_{2}\\ \hline\cr(1)&f^{+}_{0}c^{-}_{2}c^{+}_{1}c^{-}_{0}c^{+}_{2}c^{+}_{0}&\{h_{1}\}&\{6\}&s_{1}&-\varpi_{1}+\varpi_{2}\\ (2)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{0}c^{-}_{2}}f^{+}_{1}c^{-}_{0}c^{+}_{2}c^{+}_{0}&\{h_{3}\}&\{4\}&s_{1}&-\varpi_{2}\\ (3)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{0}c^{-}_{2}c^{-}_{1}c^{-}_{0}}f^{+}_{2}c^{+}_{0}&\{h_{5}\}&\{2\}&s_{1}&\varpi_{1}-3\varpi_{2}\\ \hline\cr(4)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{0}}f^{+}_{2}c^{+}_{1}c^{+}_{0}c^{+}_{2}c^{-}_{0}&\{h_{2}\}&\{5\}&s_{2}&2\varpi_{1}-2\varpi_{2}\\ (5)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{0}c^{-}_{2}c^{-}_{1}}f^{+}_{0}c^{+}_{2}c^{-}_{0}&\{h_{4}\}&\{3\}&s_{2}&\varpi_{1}-3\varpi_{2}\\ \hline\cr(6)&f^{+}_{0}f^{+}_{2}c^{+}_{1}c^{+}_{0}c^{+}_{2}c^{+}_{0}&\{h_{1},h_{2}\}&\{5,6\}&id&\varpi_{1}\\ (7)&f^{+}_{0}c^{-}_{2}c^{+}_{1}f^{+}_{0}c^{+}_{2}c^{+}_{0}&\{h_{1},h_{4}\}&\{3,6\}&id&-\varpi_{1}+\varpi_{2}\\ (8)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{0}}f^{+}_{2}c^{+}_{1}c^{+}_{0}c^{+}_{2}f^{+}_{0}&\{h_{2},h_{6}\}&\{1,5\}&id&2\varpi_{1}-2\varpi_{2}\\ (9)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{0}c^{-}_{2}}f^{+}_{1}f^{+}_{0}c^{+}_{2}c^{+}_{0}&\{h_{3},h_{4}\}&\{3,4\}&id&-\varpi_{2}\\ (10)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{0}c^{-}_{2}c^{-}_{1}}f^{+}_{0}c^{+}_{2}f^{+}_{0}&\{h_{4},h_{6}\}&\{1,3\}&id&\varpi_{1}-3\varpi_{2}\\ \hline\cr\end{array}$
By this table, we get all the coefficients $c^{w,\lambda}_{u,\mu}$ as follows.
$c^{w,\lambda}_{s_{2}s_{1},\mu}=1\text{ for }\mu=\varpi_{1}-3\varpi_{2}$, $c^{w,\lambda}_{s_{1},\mu}=(q-1)q^{-1}\text{ for }\mu=-\varpi_{1}+\varpi_{2},-\varpi_{2},\varpi_{1}-3\varpi_{2}$, $c^{w,\lambda}_{s_{2},\mu}=(q-1)q^{-1}\text{ for }\mu=2\varpi_{1}-2\varpi_{2},\varpi_{1}-3\varpi_{2}$, $c^{w,\lambda}_{id,\mu}=(q-1)^{2}q^{-2}\text{ for }\mu=\varpi_{1},-\varpi_{1}+\varpi_{2},2\varpi_{1}-2\varpi_{2},-\varpi_{2},\varpi_{1}-3\varpi_{2}$.
Next we calculate $c^{w,-\lambda}_{u,\mu}$ according to Theorem 3.9. For this case we need to chose $J\subset\{1,2,\ldots,l\}$ such that $u\stackrel{{\scriptstyle J_{>}}}{{\longrightarrow}}w$.
For example $u=id$, then there are three possible $J$ corresponding to the Bruhat chains $u<us_{\beta_{3}}<us_{\beta_{3}}s_{\beta_{2}}=w$, $u<us_{\beta_{5}}<us_{\beta_{5}}s_{\beta_{2}}=w$, $u<us_{\beta_{5}}<us_{\beta_{5}}s_{\beta_{4}}=w$. For the first case the weight $\mu$ can be calculated as (using diagram (b)),
$\mu=w\hat{r}_{J_{<}}(-\lambda)=s_{2}s_{1}\hat{r}_{h_{2}}\hat{r}_{h_{3}}(-\lambda)=s_{2}s_{1}(3\varpi_{1}-2\varpi_{2})=-2\varpi_{1}-\varpi_{2}.$
All the possible $J$ and corresponding alcove walks $\tilde{p}_{-\lambda}$ are listed in the table below.
$\begin{array}[]{|c|c|c|c|c|c|}\hline\cr p_{w}\tilde{p}_{-\lambda}&\tilde{p}_{-\lambda}&\mathcal{M}&J&\varphi(p)=u&wt(p)=w\hat{r}_{J_{<}}(-\lambda)\\ \hline\cr\hline\cr(0)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{2}c^{+}_{1}c^{+}_{2}c^{+}_{0}c^{+}_{1}c^{+}_{2}}&\{\}&\{\}&s_{2}s_{1}&-\varpi_{1}+3\varpi_{2}\\ \hline\cr(1)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{2}c^{+}_{1}c^{+}_{2}c^{+}_{0}c^{+}_{1}}f^{-}_{2}&\{h_{6}\}&\{6\}&s_{1}&\varpi_{2}\\ (2)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{2}c^{+}_{1}c^{+}_{2}}f^{-}_{0}c^{+}_{1}c^{-}_{2}&\{h_{4}\}&\{4\}&s_{1}&\varpi_{1}-\varpi_{2}\\ (3)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{2}}f^{-}_{1}c^{+}_{2}c^{-}_{0}c^{+}_{1}c^{-}_{2}&\{h_{2}\}&\{2\}&s_{1}&2\varpi_{1}-3\varpi_{2}\\ \hline\cr(4)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{2}c^{+}_{1}c^{+}_{2}c^{+}_{0}}f^{-}_{1}c^{-}_{2}&\{h_{5}\}&\{5\}&s_{2}&-2\varpi_{1}+2\varpi_{2}\\ (5)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{2}c^{+}_{1}}f^{-}_{2}c^{-}_{0}c^{-}_{1}c^{-}_{2}&\{h_{3}\}&\{3\}&s_{2}&-3\varpi_{1}+\varpi_{2}\\ \hline\cr(6)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{2}}f^{-}_{1}f^{-}_{2}c^{-}_{0}c^{-}_{1}c^{-}_{2}&\{h_{2},h_{3}\}&\{2,3\}&id&-2\varpi_{1}-\varpi_{2}\\ (7)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{2}}f^{-}_{1}c^{+}_{2}c^{-}_{0}f^{-}_{1}c^{-}_{2}&\{h_{2},h_{5}\}&\{2,5\}&id&-2\varpi_{2}\\ (8)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{2}c^{+}_{1}c^{+}_{2}}f^{-}_{0}f^{-}_{1}c^{-}_{2}&\{h_{4},h_{5}\}&\{4,5\}&id&-\varpi_{1}\\ \hline\cr\end{array}$
By this table, we get all the coefficients $c^{w,-\lambda}_{u,\mu}$ as follows.
$c^{w,-\lambda}_{s_{2}s_{1},\mu}=1\text{ for }\mu=-\varpi_{1}+3\varpi_{2}$, $c^{w,-\lambda}_{s_{1},\mu}=(1-q)q^{-1}\text{ for }\mu=\varpi_{2},\varpi_{1}-\varpi_{2},2\varpi_{1}-3\varpi_{2}$, $c^{w,-\lambda}_{s_{2},\mu}=(1-q)q^{-1}\text{ for }\mu=-2\varpi_{1}+2\varpi_{2},-3\varpi_{1}+\varpi_{2}$, $c^{w,-\lambda}_{id,\mu}=(1-q)^{2}q^{-2}\text{ for }\mu=-2\varpi_{1}-\varpi_{2},-2\varpi_{2},-\varpi_{1}$.

### v4

@@
 Primary 14M15, 14C17; Secondary 14N15, 17B10, 33D80.
-L. C. Mihalcea was supported in part by the NSF grant DMS-2152294 and a Simons Collaboration Grant. H. Naruse was supported in part by JSPS KAKENHI Grant Number 16H03921.
+L.
+C.
+Mihalcea was supported in part by the NSF grant DMS-2152294 and a Simons Collaboration Grant.
+H.
+Naruse was supported in part by JSPS KAKENHI Grant Number 16H03921.
+C.
+Su is supported by the National Key R&D Program of China (No.
+2024YFA1014700).
 1. Introduction
@@
 Let $W$ be the Weyl group determined by $(G,T)$.
-In the study of cohomology and K-theory rings of (generalized) flag manifolds $G/P$, the Chevalley formula expresses the multiplication of a Schubert class by the class of line bundle, or, equivalently, a Schubert divisor in $G/P$.
+In the study of cohomology and K-theory rings of (generalized) flag manifolds $G/P$, the Chevalley formula expresses the multiplication of a Schubert class by the class of a line bundle, or a Schubert divisor in $G/P$.
 If one works equivariantly, this formula determines completely the multiplication in the equivariant K ring.
@@
 If $y=0$, the motivic class $MC_{y}(X(w)^{\circ})$ is equal to the class of the ideal sheaf $[{\mathcal{O}}_{X(w)}(-\partial X(w))]$ of the boundary $\partial X(w)=X(w)\setminus X(w)^{\circ}$, where $X(w)=\overline{X(w)^{\circ}}$ is the Schubert variety.
-If $y=-1$, then $MC_{y}(X(w)^{\circ})$ is equal to the class of the unique $T$-fixed point in $X(w)^{\circ}$; see [AMSS22].
-The Poincaré duals of the classes $MC_{y}(X(w)^{\circ})$, the Segre motivic classes [AMSS19, MSA22], specialize when $y=0$ to the Grothendieck classes of the structure sheaves of the opposite Schubert varieties.
+If $y=-1$, then $MC_{y}(X(w)^{\circ})$ is equal to the class of the unique $T$-fixed point in $X(w)^{\circ}$; see [AMSS24b].
+The Poincaré duals of the classes $MC_{y}(X(w)^{\circ})$, the Segre motivic classes [AMSS24a, MS22], specialize when $y=0$ to the Grothendieck classes of the structure sheaves of the opposite Schubert varieties.
 Our Chevalley formula (1), and its analogous formula for the Segre motivic classes, specialize to known Chevalley formulae for K-theoretic Schubert classes and ideal sheaves from [GR04, LP07].
@@
 In fact, the power $y^{\ell(w)-\ell(u)}$ from the formula (5) is absorbed into $c_{u,\mu}^{w,\lambda}$ so it becomes polynomial in $y$.
+As mentioned above, our Chevalley formula generalizes to the motivic situation the classical Chevalley multiplication in $K_{T}(G/B)$.
+It also generalizes the Chevalley multiplication by (equivariant) Chern–Schwartz–MacPherson classes of Schubert cells from [AMSS23]; a short, self-contained, proof of this is added in an Appendix.
 […]
 We now give a rough idea on the proof of Theorem 1.1.
-The key connection between the Chevalley formula in the Hecke algebra to motivic Chern classes, proved in [MNS22b], and ultimately based on results from [AMSS19], is that the motivic Chern classes are recursively obtained by certain left Demazure–Lusztig operators $\mathcal{T}_{w}^{L}$ acting on $K_{T}(G/B)[y]$:
+The key connection between the Chevalley formula in the Hecke algebra to motivic Chern classes, proved in [MNS22b], and ultimately based on results from [AMSS24a], is that the motivic Chern classes are recursively obtained by certain left Demazure–Lusztig operators $\mathcal{T}_{w}^{L}$ acting on $K_{T}(G/B)[y]$:
 $MC_{y}(X(w)^{\circ})=\mathcal{T}_{w}^{L}[\mathcal{O}_{1.B}]\/.$
@@
 Combining these dualities shows that the polynomials $C_{v,\lambda}^{w}(y)$ are palindromic.
-A remarkable property of the motivic Chern classes of Schubert cells, proved in [AMSS19, FRW21], is that they are equivalent to the K-theoretic version of Maulik and Okounkov’s stable envelopes, see [MO19, AO21].
+A remarkable property of the motivic Chern classes of Schubert cells, proved in [AMSS24a, FRW21], is that they are equivalent to the K-theoretic version of Maulik and Okounkov’s stable envelopes, see [MO19, AO21].
 The stable envelopes are elements in the $T\times{\mathbb{C}}^{*}$-equivariant K-theory of the cotangent bundle, $K_{T\times{\mathbb{C}}^{*}}(T^{*}(G/B))$.
@@
 The stable envelopes depend on three parameters: a chamber, a polarization, and a slope, and the precise normalizations are essential for this paper.
-A variation in the chamber results in conjugating by the Borel subgroup [AMSS19], and it is encoded in the left Weyl group action [MNS22b] and certain $R$-matrix operators [RTV15, RTV19].
+A variation in the chamber results in conjugating by the Borel subgroup [AMSS24a], and it is encoded in the left Weyl group action [MNS22b] and certain $R$-matrix operators [RTV15, RTV19].
 Varying the polarization, or the slope, results in the multiplication of $\operatorname{stab}(w)$ by a line bundle $\mathcal{L}_{\lambda}$ pulled back from $G/B$; cf.
-[AMSS19, Oko17], see also §7 below.
+[AMSS24a, Oko17], see also §7 below.
 In particular, the coefficients $C_{v,\lambda}^{w}(y)$ from (1) give ‘wall-crossing’ formulae, recording the change of stable envelopes when its defining parameters are varied.
-While these wall crossing formulae have been worked out in [Oko17, SZZ20, SZZ21] (see also [KW23]), in §7 we revisit some of these from the point of view of Theorem 1.1.
+While these wall crossing formulae have been worked out in [Oko17, SZZ20, SZZ21] (see also [KW25]), in §7 we revisit some of these from the point of view of Theorem 1.1.
 In particular, we utilize the Chevalley formula to give an explicit combinatorial rule relating the stable envelope for the fundamental alcove $\operatorname{A_{\circ}}$ to the one corresponding to any translation $\operatorname{A_{\circ}}+\lambda$; see Theorem 7.8.
@@
 Finally, in an Appendix we obtain an analogue of the Chevalley formula (1) for the homological analogue of the motivic Chern classes, the Chern–Schwartz–MacPherson classes.
-While this formula may be obtained by a specialization argument as in [AMSS22], the degenerate affine Hecke algebra is much simpler in this case, and a direct proof of the Chevalley formula may be obtained rather quickly.
+While this formula may be obtained by a specialization argument as in [AMSS24b], the degenerate affine Hecke algebra is much simpler in this case, and a direct proof of the Chevalley formula may be obtained rather quickly.
 Acknowledgments.
@@
 Let $X^{*}(T)_{P}:=\{\lambda\in X^{*}(T)\mid\langle\lambda,\gamma^{\vee}\rangle=0\text{ for all }\gamma\in R^{+}_{P}\}$ be the set of integral weights which vanish on $(R^{+}_{P})^{\vee}$.
-For any $\lambda\in X^{*}(T)_{P}$, we still use ${\mathcal{L}}_{\lambda}$ to denote the line bundle $G\times^{P}{\mathbb{C}}_{\lambda}\in\operatorname{Pic}(G/P)$, which has fiber over $1.P$ the $T$-module of weight $\lambda$.
+For any $\lambda\in X^{*}(T)_{P}$, we still use ${\mathcal{L}}_{\lambda}$ to denote the line bundle $G\times^{P}{\mathbb{C}}_{\lambda}\in\operatorname{Pic}(G/P)$, which has fiber over $1.P$ the one dimensional $T$-module of weight $\lambda$.
 2. Affine Hecke algebra via alcove walk algebra
@@
 Example 2.1.
+We consider the example of the root system of type $A_{1}$.
+Let $\alpha$ be the positive root, and $\omega=\alpha/2$ be the fundamental weight.
+The weight lattice $X^{*}(T)={\mathbb{Z}}\omega$, the root lattice $Q={\mathbb{Z}}\alpha$, and the finite Weyl group $W=\{id,s_{1}=s_{\alpha}\}$.
+The affine Weyl group $W_{\operatorname{aff}}=Q\rtimes W$ has Coxeter generators $s_{1}$ and $s_{0}=t_{\alpha}s_{1}$.
+The subgroup of length zero elements in $W^{\textit{ext}}_{\operatorname{aff}}$ is $\Omega=\{id,g=t_{\omega}s_{1}\}\simeq{\mathbb{Z}}/2{\mathbb{Z}}$.
+In the following picture for $\Omega\times\mathfrak{t}_{\mathbb{R}}^{*}$, the lower sheet is the identity sheet, while the upper sheet is the sheet $g\times\mathfrak{t}_{\mathbb{R}}^{*}$.
+Each alcove $w\operatorname{A_{\circ}}$ is labeled by the corresponding $w\in W^{\textit{ext}}_{\operatorname{aff}}$, both in the Coxeter presentation and the translation presentation.
+In the lower sheet, the walls $H_{\alpha,n}$ are labeled by $1$ if $n$ is even, and 00 if $n$ is odd.
+On the upper sheet, the labelings are in the opposite way.
 […]
@@
 In terms of pictures, these generators can be drawn as follows:
---$+$$i$$c_{i}^{+}$-$+$$i$$c_{i}^{-}$-$+$$i$$f_{i}^{+}$-$+$$i$$f_{i}^{-}$
+--$+$$i$$c_{i}^{+}$
+--$+$$i$$c_{i}^{-}$
+--$+$$i$$f_{i}^{+}$
+--$+$$i$$f_{i}^{-}$
 $(0\leq i\leq r)$
@@
 We check the above theorem for $w=s_{1}$ and $\lambda=\omega$.
-From the alcove picture in Example 2.1, $T_{s_{1}}^{-1}$ is represented by the minimal length walk $q^{-1/2}c_{1}^{-1}$, while $X^{\omega}$ is represented by the walk $gc_{1}^{+}=c_{0}^{+}g$.
+From the alcove picture in Example 2.1, $T_{s_{1}}^{-1}$ is represented by the minimal length walk $q^{-1/2}c_{1}^{-}$, while $X^{\omega}$ is represented by the walk $gc_{1}^{+}=c_{0}^{+}g$.
 Thus, the sum in the right hand side of the Theorem is over the alcove walks $c_{1}^{-}c_{0}^{-}g$ and $c_{1}^{-}f_{0}^{+}g$, which end at the alcoves $t_{-\omega}s_{1}\operatorname{A_{\circ}}$ and $t_{-\omega}\operatorname{A_{\circ}}$, respectively.
@@
 There is also an equivariant version of the motivic Chern class transformation, which uses equivariant varieties and morphisms, and has values in the suitable equivariant K-theory group.
-Its definition was given in [AMSS19, FRW21], following closely the approach of [BSY10].
+Its definition was given in [AMSS24a, FRW21], following closely the approach of [BSY10].
 Assume $X$ is smooth and there is a torus $T$ acting on $X$.
-Let $K_{T}(X)$ denote the equivariant K-theory of $X$, see [CG09].
+Let $K_{T}(X)$ denote the equivariant K-theory of $X$, see [CG10].
 If $X$ is a point, $K_{T}(pt)=K^{0}(\operatorname{Rep}(T))={\mathbb{Z}}[T]$.
@@
 Theorem 4.1.
-[AMSS19, FRW21] Let $X$ be a quasi-projective, non-singular, complex algebraic variety with an action of the torus $T$.
+[AMSS24a, FRW21] Let $X$ be a quasi-projective, non-singular, complex algebraic variety with an action of the torus $T$.
 There exists a unique natural transformation $MC_{y}:G_{0}^{T}(var/X)\to K_{T}(X)[y]$ satisfying the following properties:
@@
   1. (1)
-    By [AMSS22, Thm. 5.1 and Cor. 5.3],
+    By [AMSS24b, Thm. 5.1 and Cor. 5.3],
     $MC_{0}(X(w)^{\circ})=[{\mathcal{O}}_{X(w)}(-\partial X(w))],\textit{\quad and \quad}SMC_{0}(Y(w)^{\circ})=[{\mathcal{O}}_{Y(w)}],$
@@
 4.3. Partial flag variety case
-For any $w\in W$, let $\ell(wW_{P})$ denote the length of the minimal length representative in $wW_{P}$. Let $\pi:G/B\to G/P$ be the natural projection. It is proved in [AMSS19, Remeark 5.7] that
+For any $w\in W$, let $\ell(wW_{P})$ denote the length of the minimal length representative in $wW_{P}$. Let $\pi:G/B\to G/P$ be the natural projection. It is proved in [AMSS24a, Remeark 5.7] that
 (25) $\pi_{*}MC_{y}(X(w)^{\circ})=(-y)^{\ell(w)-\ell(wW_{P})}MC_{y}(X(wW_{P})^{\circ})\/.$
@@
 In particular, the following equation holds for the Chevalley coefficients in (27):
-(29) $C_{u,-\lambda}^{w}=\sum_{\mu\in X^{*}(T)}y^{\ell(w)-\ell(u)}e^{-\mu}c_{u,\mu}^{w,\lambda}|_{q=-y}.$
+$C_{u,-\lambda}^{w}=\sum_{\mu\in X^{*}(T)}y^{\ell(w)-\ell(u)}e^{-\mu}c_{u,\mu}^{w,\lambda}|_{q=-y}.$
 From Equation 27, the Chevalley coefficients $C_{u,\lambda}^{v}$ belong to a localization $K_{T}(pt)[y]$ which allows division by $1+e^{\alpha}y$ for any root $\alpha$.
@@
 Applying the map $\Psi$ in Lemma 4.5 to Equation 7, we get
-(30) ${\mathcal{T}}^{L}_{w}e^{-\lambda}=\sum_{\mu\in X^{*}(T),u\in W}y^{\ell(w)-\ell(u)}e^{-\mu}c_{u,\mu}^{w,\lambda}|_{q=-y}{\mathcal{T}}^{L}_{u}\in\operatorname{End}_{{\mathbb{C}}}K_{T}(G/B)[y].$
+${\mathcal{T}}^{L}_{w}e^{-\lambda}=\sum_{\mu\in X^{*}(T),u\in W}y^{\ell(w)-\ell(u)}e^{-\mu}c_{u,\mu}^{w,\lambda}|_{q=-y}{\mathcal{T}}^{L}_{u}\in\operatorname{End}_{{\mathbb{C}}}K_{T}(G/B)[y].$
 Then the theorem follows by applying both sides to $[{\mathcal{O}}_{X(id)}]$, and utilizing that
@@
 In a paper in preparation we use this theorem to give a combinatorial Chevalley formula for minuscule flag varieties and a $K$-theoretic generalization of Nakada’s colored hook formula.
-We also recently learned about work by Neil Fan, Peter Guo, Changjian Su, and Rui Xiong who proved Pieri formulae for the motivic Chern classes on Grassmannians.
+See also [FGSX24] for a Pieri formula for the motivic Chern classes of Schubert cells in the equivariant K-theory of Grassmannians.
 6. Dualities of Chevalley coefficients
@@
 This finishes the proof of the theorem by comparing the above equation with Equation 27. ∎
-The star duality *** acts on $K_{T}(G/B)[y^{\pm 1}]$ and $K_{T}(pt)[y^{\pm 1}]$ by sending a vector bundle to its dual, and leaving $y^{i}$ unchanged. Consider the composition
+The star duality * acts on $K_{T}(G/B)[y^{\pm 1}]$ and $K_{T}(pt)[y^{\pm 1}]$ by sending a vector bundle to its dual, and leaving $y^{i}$ unchanged. Consider the composition
 $\iota:=w_{0}*:K_{T}(\operatorname{pt})[y^{\pm 1}]\to K_{T}(\operatorname{pt})[y^{\pm 1}]\/.$
-Then we have the following result by combining Theorem 9.1(a) and Remark 4.7 of [AMSS22].
+Then we have the following result by combining Theorem 9.1(a) and Remark 4.7 of [AMSS24b].
 (32) ${\mathbb{C}}_{-\rho}\otimes{\mathcal{L}}_{-\rho}\otimes MC_{y}(X(w)^{\circ})=(-1)^{\dim G/B-\ell(w)}\prod_{\alpha>0}(1+ye^{-\alpha})*\bigg{(}SMC_{y}(X(w)^{\circ})\bigg{)}.$
@@
 Proof.
-Applying the star duality functor *** to Equation 27, then use Equation 34 to get
+Applying the star duality functor * to Equation 27, then use Equation 32 to get
 ${\mathcal{L}}_{-\lambda}\otimes SMC_{y}(X(w)^{\circ})=\sum_{u}(-1)^{\ell(w)-\ell(u)}*(C_{u,\lambda}^{w})SMC_{y}(X(u)^{\circ}).$
@@
     (Serre duality + star duality)
-    $(C^{w}_{u,\lambda})|_{y\to y^{-1}}\times y^{\ell(w)-\ell(u)}=C^{w}_{u,\lambda}.$
+    $(C^{w}_{u,\lambda})|_{y\to y^{-1}}\times y^{\ell(w)-\ell(u)}=C^{w}_{u,\lambda}\/.$
   2. (2)
     (Star duality + Dynkin duality)
-    $C^{w}_{u,\lambda}=(-1)^{\ell(w)-\ell(u)}C^{uw_{0}}_{ww_{0},w_{0}\lambda}.$
+    $C^{w}_{u,\lambda}=(-1)^{\ell(w)-\ell(u)}C^{uw_{0}}_{ww_{0},w_{0}\lambda}\/.$
 Proof.
@@
 In this section we apply the Chevalley formula for motivic classes to calculate the transformation of stable envelopes in $T^{*}(G/B)$ under the change of arbitrary alcoves.
-For alcoves adjacent to the fundamental alcove, a formula for this transformation was obtained in [SZZ20, Theorem 5.4], see also [KW23].
+For alcoves adjacent to the fundamental alcove, a formula for this transformation was obtained in [SZZ20, Theorem 5.4], see also [KW25].
 7.1. Definition of the stable envelopes
@@
 From the definition, it is immediate to see that $\{\operatorname{stab}_{\mathfrak{C},T^{\frac{1}{2}},A}(w)\mid w\in W\}$ forms a basis for the localized equivariant K-theory $K_{T\times{\mathbb{C}}^{*}}(T^{*}(G/B))_{\operatorname{loc}}$, called the stable basis.
-Explicit combinatorial formulae and recursions for the localizations of stable envelopes in $T^{*}(G/B)$ may be found in [AMSS19, §8.3].
+Explicit combinatorial formulae and recursions for the localizations of stable envelopes in $T^{*}(G/B)$ may be found in [AMSS24a, §8.3].
 7.2. Changing the polarizations
@@
 More precisely, the group $G$ acts on $T^{*}(G/B)$ by left multiplication, which induces a left Weyl group action on $K_{T\times{\mathbb{C}}^{*}}(T^{*}(G/B))$, see [MNS22b].
-If we change the chamber $\mathfrak{C}$ to another chamber $w(\mathfrak{C})$ ($w\in W$), we have the following formula (see [AMSS19, Lemma 8.2(a)])
+If we change the chamber $\mathfrak{C}$ to another chamber $w(\mathfrak{C})$ ($w\in W$), we have the following formula (see [AMSS24a, Lemma 8.2(a)])
 $w\cdot(\operatorname{stab}_{{\mathcal{C}},T^{\frac{1}{2}},A}(u))=\operatorname{stab}_{w({\mathcal{C}}),w(T^{\frac{1}{2}}),A}(wu).$
@@
 The alcoves in $\mathfrak{t}^{*}_{\mathbb{R}}$ are of the form $x(\operatorname{A_{\circ}})+\lambda$ for some $x\in W$ and some $\lambda$ in the root lattice.
-It was proved in [AMSS19, Lemma 8.2],[SZZ20, Rmk.
+It was proved in [AMSS24a, Lemma 8.2],[SZZ20, Rmk.
 2.3] that
@@
 Let $Z=T^{*}(G/B)\times_{\mathcal{N}}T^{*}(G/B)$ be the Steinberg variety where ${\mathcal{N}}\subset\mathfrak{g}$ denotes the nilpotent cone.
-There is an algebra isomorphism due to Kazhdan–Lusztig and Ginzburg [KL87, CG09]
+There is an algebra isomorphism due to Kazhdan–Lusztig and Ginzburg [KL87, CG10]
 ${\mathbb{H}}\simeq K_{G\times{\mathbb{C}}^{*}}(Z),$
@@
 Lemma 7.7.
-[AMSS19, Theorem 8.6] Let $\iota:G/B\hookrightarrow T^{*}(G/B)$ denote the inclusion of the zero section. For any $w\in W$,
+[AMSS24a, Theorem 8.6] Let $\iota:G/B\hookrightarrow T^{*}(G/B)$ denote the inclusion of the zero section. For any $w\in W$,
 $\iota^{*}(\operatorname{stab}_{\operatorname{A_{\circ}}}(w))=(-1)^{\dim G/B}q^{\dim G/B-\frac{\ell(w)}{2}}MC_{-q^{-1}}(Y(w)^{\circ})\otimes{\mathcal{L}}_{-2\rho}.$
@@
 In this section we study Whittaker functions.
-These appear in $p$-adic representation theory, and in this note we utilize a cohomological construction of these functions from [AMSS19], see also [MSA22].
+These appear in $p$-adic representation theory, and in this note we utilize a cohomological construction of these functions from [AMSS24a], see also [MS22].
 Recall the definition of the Demazure–Lusztig operators on $K_{T}(pt)[y]$:
@@
 The operators satisfy the usual quadratic and braid relations in the Hecke algebra, therefore for any $w\in W$ there are operators $\widetilde{T_{w}}$ and $\widetilde{T^{\vee}_{w}}$ acting on $K_{T}(pt)[y]$, and defined using any reduced decomposition of $w$.
-The following has been proved in [MSA22, Thm.
+The following has been proved in [MS22, Thm.
 1.1]:
@@
 (35) $\chi_{T}\left(G/B,{\mathcal{L}}_{\lambda}\otimes MC_{y}^{\prime}(X(w)^{\circ})\right)=\widetilde{T_{w}}(e^{\lambda})={\mathcal{W}}_{\lambda,w},$
-where ${\mathcal{W}}_{\lambda,w}$ is the Iwahori–Whittaker function for the Langlands dual group over a non-archimedean local field; we refer to [MSA22] for more details, including the number theoretic definition of ${\mathcal{W}}_{\lambda,w}$.
+where ${\mathcal{W}}_{\lambda,w}$ is the Iwahori–Whittaker function for the Langlands dual group over a non-archimedean local field; we refer to [MS22] for more details, including the number theoretic definition of ${\mathcal{W}}_{\lambda,w}$.
 Here we identified $y$ with $-q^{-1}$, where $q$ is the number of elements in the residue field.
-As explained in [MSA22], from the fact that
+As explained in [MS22], from the fact that
 $\sum_{w\in W}MC_{y}^{\prime}(X(w)^{\circ})=1$
@@
 Here $\chi_{w_{0}\lambda}$ denotes the character for the irreducible representation of $G$ of highest weight $w_{0}(\lambda)$.
-We also note that, in type A, an interpretation of the Iwahori–Whittaker function in terms of the partition function of the Iwahori lattice model has been obtained in [BBBG19].
+We also note that, in type A, an interpretation of the Iwahori–Whittaker function in terms of the partition function of the Iwahori lattice model has been obtained in [BBBG24].
 Using the Chevalley coefficients in Equation 27, we obtain the following formula for the Iwahori–Whittaker function ${\mathcal{W}}_{\lambda,w}$. Let $\rho$ denote the half sum of the positive roots.
@@
 (37) $\chi_{T}(G/B,MC_{y}(X(u)^{\circ}))=MC_{y}[X(u)^{\circ}\rightarrow\operatorname{pt}]=MC_{y}[{\mathbb{A}}^{1}\rightarrow\operatorname{pt}]^{\ell(u)}=(-y)^{\ell(u)},$
-where the second equality follows from [AMSS19, Theorem 4.2(3)]. Therefore, taking the equivariant Euler characteristics of both sides of Equation 27, we get
+where the second equality follows from [AMSS24a, Theorem 4.2(3)]. Therefore, taking the equivariant Euler characteristics of both sides of Equation 27, we get
 $\widetilde{T^{\vee}_{w}}(e^{\lambda})=\chi_{T}\left(G/B,{\mathcal{L}}_{\lambda}\otimes MC_{y}(X(w)^{\circ})\right)=\sum_{u}C_{\lambda,u}^{w}(-y)^{\ell(u)}.$
@@
 $\chi_{T}\left(G/B,{\mathcal{L}}_{\rho}\otimes MC_{y}^{\prime}(X(w)^{\circ})\right)=(-1)^{\ell(w)}e^{\rho}.$
-As a corollary we prove a variant of the Casselman–Shalika formula (38), obtained by Li [Li92], see also [BBBG19, Proposition 9.4]. First define
+As a corollary we prove a variant of the Casselman–Shalika formula (36), obtained by Li [Li92], see also [BBBG24, Proposition 9.4]. First define
 (39) $R_{\lambda}(y):=\chi_{T}(G/B,\lambda_{y}(T^{*}_{G/B})\otimes{\mathcal{L}}_{\lambda})\in K_{T}(pt)[y].$
@@
 Equation 42 follows from the localization formula (cf.
-[Nie74], [MSA22, Theorem 2.1 (c)].) To be more specific, the torus fixed points in $G/P_{\lambda}$ are $\{wP_{\lambda}\mid w\in W^{P}\}$, and the torus weights of the tangent space at the fixed point $wP_{\lambda}$ are $\{-w\alpha\mid\alpha\in R^{+}\setminus R^{+}_{\lambda}\}$.
+[Nie74], [MS22, Theorem 2.1 (c)].) To be more specific, the torus fixed points in $G/P_{\lambda}$ are $\{wP_{\lambda}\mid w\in W^{P}\}$, and the torus weights of the tangent space at the fixed point $wP_{\lambda}$ are $\{-w\alpha\mid\alpha\in R^{+}\setminus R^{+}_{\lambda}\}$.
 ∎
@@
 Appendix A Chevalley formulae for the Chern–Schwartz–MacPherson classes of Schubert cells
-In this appendix, we give a short proof of the Chevalley formulae for the equivariant Chern–Schwartz–MacPherson (CSM) and Segre–MacPherson (SM) classes of the Schubert cells in the (partial) flag varieties, see [AMSS17, Thm  9.10] and Theorem A.4 below.
+In this appendix, we give a short proof of the Chevalley formulae for the equivariant Chern–Schwartz–MacPherson (CSM) and Segre–MacPherson (SM) classes of the Schubert cells in the (partial) flag varieties, see [AMSS23, Thm  9.10] and Theorem A.4 below.
 Our proof again relies on the action of the appropriate Hecke algebra, this time on the equivariant cohomology of $G/P$.
@@
 We refer the reader e.g.
-to [Ohm06, AMSS17] for details, including a construction of these classes in the equivariant setting and for general varieties.
-[…]
+to [Ohm06, AMSS23] for details, including a construction of these classes in the equivariant setting and for general varieties.
+The (additive) group of constructible functions ${\mathcal{F}}(X)$ consists of functions $\varphi=\sum_{Z}c_{Z}1\hskip-3.5pt1_{Z}$, where the sum is over a finite set of constructible subsets $W\subset X$, $c_{Z}\in{\mathbb{Z}}$ are integers, and $1\hskip-3.5pt1_{Z}$ is the characteristic function of $Z$.
+For a proper morphism $f:Y\rightarrow X$, there is a linear map $f_{*}:{\mathcal{F}}(Y)\rightarrow{\mathcal{F}}(X)$, such that for any constructible subset $Z\subset Y$, $f_{*}(1\hskip-3.5pt1_{Z})(x)=\chi_{\textit{top}}(f^{-1}(x)\cap Z)$, where $x\in X$ and $\chi_{\textit{top}}$ denotes the topological Euler characteristic.
+A conjecture attributed to Deligne and Grothendieck states that there is a unique natural transformation $c_{*}:{\mathcal{F}}\to H_{*}$ from the functor of constructible functions on a complex algebraic variety $X$ to the homology functor, where all morphisms are proper, such that if $X$ is smooth then $c_{*}(1\hskip-3.5pt1_{X})=c(T_{X})\cap[X]$.
+This conjecture was proved by MacPherson [Mac74]; the class $c_{*}(1\hskip-3.5pt1_{X})$ for possibly singular $X$ was shown to coincide with a class defined earlier by M.-H.
+Schwartz [Sch65a, Sch65b, BS81].
 There is an equivariant version of MacPherson’s transformation defined by Ohmoto [Ohm06].
@@
 The sets of CSM classes of Schubert cells $\{c_{\mathrm{SM}}(X(wW_{P})^{\circ})\mid w\in W^{P}\}$ and of the SM classes $\{s_{\mathrm{M}}(X(wW_{P})^{\circ})\mid w\in W^{P}\}$ form bases for $H_{T}^{*}(G/P)_{\operatorname{loc}}:=H_{T}^{*}(G/P)\otimes_{H_{T}^{*}(pt)}\operatorname{Frac}H_{T}^{*}(pt)$.
-Moreover, if one takes the opposite Schubert cells in any of these sets, then the two bases are dual under the usual intersection pairing, see [AMSS17, Theorem 9.4]:
+Moreover, if one takes the opposite Schubert cells in any of these sets, then the two bases are dual under the usual intersection pairing, see [AMSS23, Theorem 9.4]:
 (47) $\langle c_{\mathrm{SM}}(X(wW_{P})^{\circ}),s_{\mathrm{M}}(Y(uW_{P})^{\circ})\rangle_{G/P}=\delta_{w,u}\textit{ for any }w,u\in W^{P}.$
@@
 Recall for any $\lambda\in X^{*}(T)_{P}$, ${\mathcal{L}}_{\lambda}:=G\times^{P}{\mathbb{C}}_{\lambda}\in\operatorname{Pic}_{T}(G/P)$.
-The following is our main result in this Appendix, and it has also been proved in [AMSS17, Thm.
+The following is our main result in this Appendix, and it has also been proved in [AMSS23, Thm.
 9.10] using the Chevalley formula for the cohomological stable envelopes from [Su16, Thm.
@@
 $h_{1}=H_{\alpha_{2},0},h_{2}=H_{\alpha_{1}+\alpha_{2},0},h_{3}=H_{\alpha_{1},0},h_{4}=H_{\alpha_{1}+\alpha_{2},-1},h_{5}=H_{\alpha_{1},-1},h_{6}=H_{\alpha_{1}+\alpha_{2},-2}.$
-(a) Alcove walk $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}p_{\lambda}$ from $\operatorname{A_{\circ}}$ to $\operatorname{A_{\circ}}+\lambda$$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}p_{w}=c^{-}_{2}c^{-}_{1}$ , $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}p_{\lambda}=c^{+}_{0}c^{+}_{2}c^{+}_{1}c^{+}_{0}c^{+}_{2}c^{+}_{0}$
 […]
+(a) Alcove walk $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}p_{\lambda}$ from $\operatorname{A_{\circ}}$ to $\operatorname{A_{\circ}}+\lambda$
+$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}p_{w}=c^{-}_{2}c^{-}_{1}$ , $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}p_{\lambda}=c^{+}_{0}c^{+}_{2}c^{+}_{1}c^{+}_{0}c^{+}_{2}c^{+}_{0}$
 […]
-(a) Alcove walk $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}p_{\lambda}$ from $\operatorname{A_{\circ}}$ to $\operatorname{A_{\circ}}+\lambda$$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}p_{w}=c^{-}_{2}c^{-}_{1}$ , $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}p_{\lambda}=c^{+}_{0}c^{+}_{2}c^{+}_{1}c^{+}_{0}c^{+}_{2}c^{+}_{0}$
-(b) Alcove walk $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}p_{-\lambda}$ from $\operatorname{A_{\circ}}$ to $\operatorname{A_{\circ}}-\lambda$$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}p_{w}=c^{-}_{2}c^{-}_{1}$, $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}p_{-\lambda}=c^{-}_{2}c^{-}_{1}c^{-}_{2}c^{-}_{0}c^{-}_{1}c^{-}_{2}$
+(b) Alcove walk $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}p_{-\lambda}$ from $\operatorname{A_{\circ}}$ to $\operatorname{A_{\circ}}-\lambda$
+$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}p_{w}=c^{-}_{2}c^{-}_{1}$, $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}p_{-\lambda}=c^{-}_{2}c^{-}_{1}c^{-}_{2}c^{-}_{0}c^{-}_{1}c^{-}_{2}$
 We first calculate $c_{u,\mu}^{w,\lambda}$ according to the formula Theorem 3.10. In the proof we introduced $\lambda$-shifted (reversed) hyperplane sequence which can be seen in diagram (a).
@@
 diagram (a).)
-$\begin{array}[]{|c|c|c|c|c|c|}\hline\cr p_{w}\tilde{p}_{\lambda}&\tilde{p}_{\lambda}&\mathcal{M}&J&\varphi(p)=u&{\rm wt}(p)=w\tilde{r}^{\lambda}_{J_{>}}(\lambda)\\ \hline\cr\hline\cr(0)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{0}c^{-}_{2}c^{-}_{1}c^{-}_{0}c^{-}_{2}c^{+}_{0}}&\{\}&\{\}&s_{2}s_{1}&\varpi_{1}-3\varpi_{2}\\ \hline\cr(1)&f^{+}_{0}c^{-}_{2}c^{+}_{1}c^{-}_{0}c^{+}_{2}c^{+}_{0}&\{h_{1}\}&\{6\}&s_{1}&-\varpi_{1}+\varpi_{2}\\ (2)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{0}c^{-}_{2}}f^{+}_{1}c^{-}_{0}c^{+}_{2}c^{+}_{0}&\{h_{3}\}&\{4\}&s_{1}&-\varpi_{2}\\ (3)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{0}c^{-}_{2}c^{-}_{1}c^{-}_{0}}f^{+}_{2}c^{+}_{0}&\{h_{5}\}&\{2\}&s_{1}&\varpi_{1}-3\varpi_{2}\\ \hline\cr(4)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{0}}f^{+}_{2}c^{+}_{1}c^{+}_{0}c^{+}_{2}c^{-}_{0}&\{h_{2}\}&\{5\}&s_{2}&2\varpi_{1}-2\varpi_{2}\\ (5)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{0}c^{-}_{2}c^{-}_{1}}f^{+}_{0}c^{+}_{2}c^{-}_{0}&\{h_{4}\}&\{3\}&s_{2}&\varpi_{1}-3\varpi_{2}\\ \hline\cr(6)&f^{+}_{0}f^{+}_{2}c^{+}_{1}c^{+}_{0}c^{+}_{2}c^{+}_{0}&\{h_{1},h_{2}\}&\{5,6\}&id&\varpi_{1}\\ (7)&f^{+}_{0}c^{-}_{2}c^{+}_{1}f^{+}_{0}c^{+}_{2}c^{+}_{0}&\{h_{1},h_{4}\}&\{3,6\}&id&-\varpi_{1}+\varpi_{2}\\ (8)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{0}}f^{+}_{2}c^{+}_{1}c^{+}_{0}c^{+}_{2}f^{+}_{0}&\{h_{2},h_{6}\}&\{1,5\}&id&2\varpi_{1}-2\varpi_{2}\\ (9)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{0}c^{-}_{2}}f^{+}_{1}f^{+}_{0}c^{+}_{2}c^{+}_{0}&\{h_{3},h_{4}\}&\{3,4\}&id&-\varpi_{2}\\ (10)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{0}c^{-}_{2}c^{-}_{1}}f^{+}_{0}c^{+}_{2}f^{+}_{0}&\{h_{4},h_{6}\}&\{1,3\}&id&\varpi_{1}-3\varpi_{2}\\ \hline\cr\end{array}$
+$\begin{array}[]{|c|c|c|c|c|c|}\hline\cr p_{w}\tilde{p}_{\lambda}&\tilde{p}_{\lambda}&\mathcal{M}&J&\varphi(p)=u&{\rm wt}(p)=w\tilde{r}^{\lambda}_{J_{>}}(\lambda)\\ \hline\cr\hline\cr(0)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}c^{-}_{0}c^{-}_{2}c^{-}_{1}c^{-}_{0}c^{-}_{2}c^{+}_{0}}&\{\}&\{\}&s_{2}s_{1}&\varpi_{1}-3\varpi_{2}\\ \hline\cr(1)&f^{+}_{0}c^{-}_{2}c^{+}_{1}c^{-}_{0}c^{+}_{2}c^{+}_{0}&\{h_{1}\}&\{6\}&s_{1}&-\varpi_{1}+\varpi_{2}\\ (2)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}c^{-}_{0}c^{-}_{2}}f^{+}_{1}c^{-}_{0}c^{+}_{2}c^{+}_{0}&\{h_{3}\}&\{4\}&s_{1}&-\varpi_{2}\\ (3)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}c^{-}_{0}c^{-}_{2}c^{-}_{1}c^{-}_{0}}f^{+}_{2}c^{+}_{0}&\{h_{5}\}&\{2\}&s_{1}&\varpi_{1}-3\varpi_{2}\\ \hline\cr(4)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}c^{-}_{0}}f^{+}_{2}c^{+}_{1}c^{+}_{0}c^{+}_{2}c^{-}_{0}&\{h_{2}\}&\{5\}&s_{2}&2\varpi_{1}-2\varpi_{2}\\ (5)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}c^{-}_{0}c^{-}_{2}c^{-}_{1}}f^{+}_{0}c^{+}_{2}c^{-}_{0}&\{h_{4}\}&\{3\}&s_{2}&\varpi_{1}-3\varpi_{2}\\ \hline\cr(6)&f^{+}_{0}f^{+}_{2}c^{+}_{1}c^{+}_{0}c^{+}_{2}c^{+}_{0}&\{h_{1},h_{2}\}&\{5,6\}&id&\varpi_{1}\\ (7)&f^{+}_{0}c^{-}_{2}c^{+}_{1}f^{+}_{0}c^{+}_{2}c^{+}_{0}&\{h_{1},h_{4}\}&\{3,6\}&id&-\varpi_{1}+\varpi_{2}\\ (8)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}c^{-}_{0}}f^{+}_{2}c^{+}_{1}c^{+}_{0}c^{+}_{2}f^{+}_{0}&\{h_{2},h_{6}\}&\{1,5\}&id&2\varpi_{1}-2\varpi_{2}\\ (9)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}c^{-}_{0}c^{-}_{2}}f^{+}_{1}f^{+}_{0}c^{+}_{2}c^{+}_{0}&\{h_{3},h_{4}\}&\{3,4\}&id&-\varpi_{2}\\ (10)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}c^{-}_{0}c^{-}_{2}c^{-}_{1}}f^{+}_{0}c^{+}_{2}f^{+}_{0}&\{h_{4},h_{6}\}&\{1,3\}&id&\varpi_{1}-3\varpi_{2}\\ \hline\cr\end{array}$
 By this table, we get all the coefficients $c^{w,\lambda}_{u,\mu}$ as follows.
@@
 All the possible $J$ and corresponding alcove walks $\tilde{p}_{-\lambda}$ are listed in the table below.
-$\begin{array}[]{|c|c|c|c|c|c|}\hline\cr p_{w}\tilde{p}_{-\lambda}&\tilde{p}_{-\lambda}&\mathcal{M}&J&\varphi(p)=u&wt(p)=w\hat{r}_{J_{<}}(-\lambda)\\ \hline\cr\hline\cr(0)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{2}c^{+}_{1}c^{+}_{2}c^{+}_{0}c^{+}_{1}c^{+}_{2}}&\{\}&\{\}&s_{2}s_{1}&-\varpi_{1}+3\varpi_{2}\\ \hline\cr(1)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{2}c^{+}_{1}c^{+}_{2}c^{+}_{0}c^{+}_{1}}f^{-}_{2}&\{h_{6}\}&\{6\}&s_{1}&\varpi_{2}\\ (2)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{2}c^{+}_{1}c^{+}_{2}}f^{-}_{0}c^{+}_{1}c^{-}_{2}&\{h_{4}\}&\{4\}&s_{1}&\varpi_{1}-\varpi_{2}\\ (3)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{2}}f^{-}_{1}c^{+}_{2}c^{-}_{0}c^{+}_{1}c^{-}_{2}&\{h_{2}\}&\{2\}&s_{1}&2\varpi_{1}-3\varpi_{2}\\ \hline\cr(4)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{2}c^{+}_{1}c^{+}_{2}c^{+}_{0}}f^{-}_{1}c^{-}_{2}&\{h_{5}\}&\{5\}&s_{2}&-2\varpi_{1}+2\varpi_{2}\\ (5)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{2}c^{+}_{1}}f^{-}_{2}c^{-}_{0}c^{-}_{1}c^{-}_{2}&\{h_{3}\}&\{3\}&s_{2}&-3\varpi_{1}+\varpi_{2}\\ \hline\cr(6)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{2}}f^{-}_{1}f^{-}_{2}c^{-}_{0}c^{-}_{1}c^{-}_{2}&\{h_{2},h_{3}\}&\{2,3\}&id&-2\varpi_{1}-\varpi_{2}\\ (7)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{2}}f^{-}_{1}c^{+}_{2}c^{-}_{0}f^{-}_{1}c^{-}_{2}&\{h_{2},h_{5}\}&\{2,5\}&id&-2\varpi_{2}\\ (8)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{2}c^{+}_{1}c^{+}_{2}}f^{-}_{0}f^{-}_{1}c^{-}_{2}&\{h_{4},h_{5}\}&\{4,5\}&id&-\varpi_{1}\\ \hline\cr\end{array}$
+$\begin{array}[]{|c|c|c|c|c|c|}\hline\cr p_{w}\tilde{p}_{-\lambda}&\tilde{p}_{-\lambda}&\mathcal{M}&J&\varphi(p)=u&wt(p)=w\hat{r}_{J_{<}}(-\lambda)\\ \hline\cr\hline\cr(0)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}c^{-}_{2}c^{+}_{1}c^{+}_{2}c^{+}_{0}c^{+}_{1}c^{+}_{2}}&\{\}&\{\}&s_{2}s_{1}&-\varpi_{1}+3\varpi_{2}\\ \hline\cr(1)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}c^{-}_{2}c^{+}_{1}c^{+}_{2}c^{+}_{0}c^{+}_{1}}f^{-}_{2}&\{h_{6}\}&\{6\}&s_{1}&\varpi_{2}\\ (2)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}c^{-}_{2}c^{+}_{1}c^{+}_{2}}f^{-}_{0}c^{+}_{1}c^{-}_{2}&\{h_{4}\}&\{4\}&s_{1}&\varpi_{1}-\varpi_{2}\\ (3)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}c^{-}_{2}}f^{-}_{1}c^{+}_{2}c^{-}_{0}c^{+}_{1}c^{-}_{2}&\{h_{2}\}&\{2\}&s_{1}&2\varpi_{1}-3\varpi_{2}\\ \hline\cr(4)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}c^{-}_{2}c^{+}_{1}c^{+}_{2}c^{+}_{0}}f^{-}_{1}c^{-}_{2}&\{h_{5}\}&\{5\}&s_{2}&-2\varpi_{1}+2\varpi_{2}\\ (5)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}c^{-}_{2}c^{+}_{1}}f^{-}_{2}c^{-}_{0}c^{-}_{1}c^{-}_{2}&\{h_{3}\}&\{3\}&s_{2}&-3\varpi_{1}+\varpi_{2}\\ \hline\cr(6)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}c^{-}_{2}}f^{-}_{1}f^{-}_{2}c^{-}_{0}c^{-}_{1}c^{-}_{2}&\{h_{2},h_{3}\}&\{2,3\}&id&-2\varpi_{1}-\varpi_{2}\\ (7)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}c^{-}_{2}}f^{-}_{1}c^{+}_{2}c^{-}_{0}f^{-}_{1}c^{-}_{2}&\{h_{2},h_{5}\}&\{2,5\}&id&-2\varpi_{2}\\ (8)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}c^{-}_{2}c^{+}_{1}c^{+}_{2}}f^{-}_{0}f^{-}_{1}c^{-}_{2}&\{h_{4},h_{5}\}&\{4,5\}&id&-\varpi_{1}\\ \hline\cr\end{array}$
 By this table, we get all the coefficients $c^{w,-\lambda}_{u,\mu}$ as follows.